### v1

# DiverseNet: Decision Diversified Semi-supervised Semantic Segmentation Networks for Remote Sensing Imagery

Abstract
Semi-supervised learning is designed to help reduce the cost of the manual labelling process by exploiting the utilisation of useful features from a large quantity of unlabelled data during training. Since pixel-level manual labelling in large-scale remote sensing imagery is expensive, semi-supervised learning becomes an appropriate solution to this. However, most of the existing semi-supervised learning methods still lack efficient perturbation methods to promote the precision and diversity of features (or pseudo labels) during training. In order to fill this gap, we propose DiverseNet architectures which explore multi-head and multi-model semi-supervised learning algorithms by simultaneously promoting precision and diversity during training. The two proposed methods of DiverseNet, namely the DiverseHead and DiverseModel, significantly achieve the highest semantic segmentation performance in four widely utilised remote sensing imagery data sets compared to state-of-the-art semi-supervised learning methods. Meanwhile, the proposed DiverseHead architecture is relatively lightweight in terms of parameter space compared to the state-of-the-art methods whilst reaching high-performance results for all utilised data sets.
Index Terms:
Semi-supervised Learning, Consistency Learning, Land Cover Classification
I Introduction
Supervised deep learning has gradually become the dominant technique in computer vision, especially during the last decade. Specifically, with its great success in semantic segmentation, automated land cover classification approaches have been significantly improved because of the high similarity between the two tasks [vali2020deep]. Nevertheless, supervised learning necessitates a substantial, meticulously labelled dataset. In the case of extensive remote sensing data, such as satellite imagery and drone-captured images in complex terrains, acquiring pixel-wise expert annotations proves to be a time-consuming, labour-intensive, and costly endeavour. While the field of computer vision boasts numerous well-annotated datasets, transferring deep learning models trained on these datasets to the remote sensing domain is a formidable challenge. This is primarily because of the significant disparities between natural images and remote sensing data, including hyperspectral and synthetic aperture radar (SAR) imagery, which often possesses unconventional image geometries, making the process non-intuitive. Fortunately, thanks to open-access remote sensing sources, a large amount of unlabelled imagery can be accessed easily and freely. Thus, only using a limited amount of labelled data and exploiting the benefits of a large amount of unlabelled data could be a feasible solution for solving the problem of the lack of labelled remote sensing datasets [wang2022semi].
In computer vision, semi-supervised learning has become a trending research direction to eliminate the labour-intensive and expensive annotation stages mentioned above. It leads to competitive performance compared to supervised learning in applications such as image classification [zhang2021flexmatch, huang2021embedding, wang2021semi2] and segmentation [hu2021semi, wang2022semi_u2pl, wang2021semi], which just use a part of labelled data and a large amount of unlabelled data. Specifically, mainstream semi-supervised learning methods exploit “pseudo” labels that are fake and/or auto-generated label information obtained as a result of an unsupervised step. Since pseudo labels are involved in the model training stage, the accuracy of pseudo labels is regarded as a key factor affecting the performance of semi-supervised learning approaches. However, the accuracy of pseudo labels is not always satisfactory especially when labelled training data is not complete and sufficient enough. In this case, augmenting the pseudo-labels to increase diversity has emerged as another key research direction within the semi-supervised learning domain. A greater diversity aids in constructing more comprehensive pseudo-labels, which, in turn, indirectly enhances the quality of these labels and the training effectiveness of the model.
(a) DiverseHead
(b) DiverseModel
Figure 1: Two kinds of Pseudo label generation methods based on multiple heads (a) DiverseHead and multiple models (b) DiverseModel.
In order to empower networks to harness the unexplored potentials within unlabeled data, a notable technique known as consistency regularization has emerged as a widely adopted method for semi-supervised learning. [french2019semi, filipiak2021n]. Specifically, consistency regularization methods rely on the assumption of smoothing, which posits that if two points reside in a high-density region and are close to each other, their corresponding labels should be the same or consistent [chen2010semi]. In practice, these consistency regularization methods achieve this assumption by forcing networks to produce consistent predictions for unlabeled inputs that undergo diverse perturbations.
The strategies for the above-mentioned perturbations can be categorised into three groups, namely input, feature, and network perturbations. Input perturbation is a method that involves making adjustments or alterations to the input data as part of the training process. In this way, the semi-supervised learning approach promotes the consistency of predictions made in response to these modified inputs. Widely used input perturbation methods employed in semi-supervised learning include Cutout[devries2017improved], MixUp [zhang2017mixup], and CutMix [yun2019cutmix]. Among them, CutMix stands out by combining aspects of MixUp and CutOut techniques, demonstrating superior performance compared to them [french2019semi]. However, input perturbation introduces artificial noise to the data, which might lead to incorrect or noisy labels when generating pseudo-labels for unlabeled examples. This might negatively impact the training process by providing incorrect guidance to the model when using pseudo labels. Cross-consistency training (CCT) [ouali2020semi] presents a feature perturbation method that introduces noise to both low- and high-level features. These perturbed features are then fed into multi-decoders to generate multiple outputs, followed by the enforcement of consistency among the outputs obtained from different decoders. However, feature perturbation alters the intrinsic characteristics and relationships between features, which also might lead to the creation of artificial feature representations that do not accurately reflect the underlying patterns in the original data. Consequently, the perturbed features could result in incorrect or noisy labels when generating pseudo-labels for unlabeled examples, which negatively impacts the learning process. Lastly, the network perturbation uses different networks or the same network with different initialization weights to promote diversity of predictions, such as cross pseudo supervision (CPS)[chen2021semi], confidence-guided semi-supervised learning (CGSSL)[ma2023confidence], mean teacher (MT), [tarvainen2017mean], CCT [ouali2020semi], guided collaborative training (GCT) [ke2020guided] and cross teaching between CNN and transformer [luo2022semi]. Compared to the previous two perturbation methods, although network perturbation methods introduce perturbations created by the model itself instead of artificial noise, they generally require rapidly increased computational resources due to the increased number of networks (whole or their inner stages) utilized.
I-A Contributions
The preceding conversation underscores that, although perturbation techniques have their advantages, they also bring inherent drawbacks. Therefore, the creation of methods based on perturbations is a complex task, necessitating the exploration of more sophisticated approaches to augment the potential of semi-supervised learning frameworks by incorporating advanced perturbation methods. This paper proposes perturbation-based semi-supervised network architectures, coined as DiverseNet, which consist of two novel feature and decision diversifying network architectures for the purposes of semi-supervised semantic segmentation of remote sensing imagery. First, we develop a lightweight semi-supervised semantic segmentation network called DiverseHead based on multiple decision heads to promote precision and diversity of pseudo labels whilst using two novel perturbation methods of dropout (DT) and dynamic freezing (DF). In addition, we also propose a high-performance multi-model structure called DiverseModel for scenarios equipped with high-memory computational resources. Specifically, the contributions of this work are as follows:
  1. 1.
    We promote utilisation of multiple decision heads to provide a semi-supervised learning structure, called DiverseHead. This structure is inspired by Bagging (also called bootstrap aggregating) and provides diversity for parameters and features. The main source of perturbation is from the dynamic choosing heads to be updated for each iteration. This architecture does not distinguish the main or auxiliary decoder. Also, all decoders are trained by both supervised and unsupervised losses, which comes with the advantages of high-quality perturbation and diversity of semi-supervised segmentation tasks for remote sensing imagery, whilst providing a relatively lightweight architecture.
  2. 2.
    For the purposes of creating perturbation in diversifying decisions, we ustilised two important techniques, namely ’dynamic freezing’ and ’dropout’. We make these techniques fit with DiverseHead semi-supervised learning architecture for the purpose of diversifying the parameters of the network. The two perturbation methods differ from input/feature perturbation methods, in which the source of feature diversity is from parameter diversity instead of introducing artificial noise to data or features.
  3. 3.
    A previously proposed architecture DiverseModel [ma2023confidence] has been extended for detailed evaluation and in a detailed comparison study for various semantic segmentation data sets. This architecture is based on consistency learning among multiple neural networks with distinct weights.
The rest of the paper is organised as follows: subsection I-B discusses the related work on semi-supervised semantic segmentat ion in remote sensing and some basic knowledge on ensemble machine learning whilst in Section II and III, the proposed algorithms DiverseHead and DiverseModel are presented. Section IV describes the segmentation dataset of remote sensing imagery. The experimental setting and the qualitative and quantitative analyses of the experimental results are performed in Section V. Section VI concludes the paper with a summary.
I-B Related Work
Semantic segmentation is one of the well-studied and wide-spread research areas in remote sensing society, for applications such as land cover classification/mapping [dong2019land], building change detection [zheng2022hfa], road extraction [ghandorh2022semantic], and marine debris detection [booth2023high]. Also, the number of research on semantic segmentation in remote sensing in the literature is rapidly increasing with the success of deep learning in computer vision, due to the strong task-similarity between two areas. Fully convolutional networks (FCNs) [long2015fully] is one group of the most widespread networks for segmentation tasks. They have made a considerable contribution to various segmentation tasks either in remote sensing or computer vision. Following the FCNs’ success, SegNet [badrinarayanan2017segnet] and UNet [ronneberger2015u] adopt a symmetrical encoder-decoder structure with skip connections, leveraging multi-stage features within the encoder. Alternatively, PSPNet [zhao2017pyramid] introduced a pyramid pooling structure that helps provide a global contextual understanding for pixel-level scene parsing. DeepLab architecture [chen2017deeplab] proposed a new type of convolution called atrous convolution and an atrous spatial pyramid pooling (ASPP) operation to enable the network to have the ability to control the spatial perception range of convolution kernel via setting different dilation rates. DeepLab has then been extended to DeepLabv3+ [chen2018encoder] as an improved hybrid approach that combines an encoder-decoder structure and the ASPP. Within DeepLabv3+, the ASPP module is incorporated following a ResNet backbone, enabling the network to delve into deeper layers and extract high-level features, which is helpful for accurately segmenting pixels around boundaries[xia2021cloud, zhang2019multi]. DeepLabv3+ is one of the most used networks in the literature for semi-supervised learning segmentation in the computer vision area.
Semi-supervised learning aims to alleviate the need for expensive annotation work and improve the models’ performances by making use of unlabelled data. Self-training [lee2013pseudo] (also known as pseudo labelling) represents one of the primitive semi-supervised learning strategies for both classification [kim2020distribution] and segmentation [zhu2021improving] purposes that involves the generation of pseudo labels through model predictions. These pseudo-labels are subsequently utilized to retrain the model. Various variations of self-training have been proposed focusing on determining the appropriate pseudo labels. Another widely developed semi-supervised learning approach called consistency regularization [french2019semi] is to force networks to give consistent predictions for unlabeled inputs that undergo diverse perturbations. CCT [ouali2020semi] employs an encoder-decoder architecture with multiple auxiliary decoders. These decoders introduce diversity in the output by feature perturbations specific to each auxiliary decoder. They calculate MSE loss between predictions of the main decoder and each auxiliary decoder without creating pseudo labels. It is worth noting that the unsupervised loss is not used to supervise the main decoder. Following CCT, subsequent methods like GCT [ke2020guided] and CPS [chen2021semi] are proposed to introduce network perturbation for consistency regularization. Specifically, they use two same network structures but with different weight initialization. CPS differs from GCT by employing pseudo labels to enforce consistency, whereas GCT achieves consistency regularization by utilizing predictions of networks cooperating with a flow detector. Although the network perturbation achieved through different weight initialization in the previous two works shows a good ability to generate diversity, we believe that employing different networks could yield even greater diversity.
Ensemble machine learning is a concept that employs multiple learners and combines their predictions [sewell2008ensemble]. Bagging, a form of ensemble learning, is a technique aimed at reducing prediction variance by creating multiple iterations of a predictor and then utilizing them to form an aggregated predictor [breiman1996bagging]. Specifically, Bagging creates sample subsets by randomly selecting from the training dataset and subsequently utilizes these acquired subsets to train the foundational models for integration. When forecasting a numerical result, the aggregation involves taking an average of the versions, while predicting a class, it relies on a majority vote. Based on the concept, bagging is a commonly employed approach for enhancing the robustness and precision of machine learning algorithms in terms of classification and regression applications [ren2016ensemble].
II Cross-head Supervision (DiverseHead)
Figure 2: DiverseHead: an online semi-supervised learning approach. The slide windows of the freezer randomly select a certain number of heads to freeze the parameter update of heads. A mainstream head is randomly selected to be supervised by unsupervised loss.
This section introduces the details of the proposed DiverseHead Semi-Supervised Learning method. Inspired by Bagging (also called bootstrap aggregating), DiverseHead is proposed to create various predictions by multiple heads with even the same structure. The source of perturbation is from the dynamic choosing heads to be updated for each iteration, called dynamic freezing, which is equivalent to using different training data for different heads in each iteration. The proposed DiverseHead differs from CCT, which does not distinguish the main or auxiliary decoder. Also, all decoders are trained by both supervised and unsupervised losses. In addition, the proposed method combines self-training and consistency regularization, which calculates unsupervised loss by using optimized pseudo labels. Specially, given are both a labelled dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ containing M images and an unlabelled data set $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$ with N images. The network $Q$ is constructed with multiple heads denoted as $\left\{head^{i}\right\}_{i=1}^{len(heads)}$. Each of these heads is initialized differently. The expectation is to gain the trained network $Q$ leveraging these labelled and unlabelled data based on the proposed semi-supervised learning methods.
The output of each head refers to a probability-based prediction for all classifications corresponding to its input. When working with labelled data, the supervised loss $\mathcal{L}_{sup}^{j}$ between the $k\leavevmode\nobreak\ th$ ground truth $y_{k}$ and its corresponding prediction $p_{k}^{j}$ for $j\leavevmode\nobreak\ th$ head is defined by using the standard cross-entropy loss function $\ell_{ce}$:
$\mathcal{L}_{sup}^{j}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k}^{j},y_{k}\right)\right),$ (1)
where $W$ and $H$ refer to the weight and height of input images. The ultimate supervised loss is determined by the mean of the losses across all heads
$\mathcal{L}_{sup}=\frac{1}{len(heads)}\sum_{j=1}^{len(heads)}\mathcal{L}_{sup}^{j}.$ (2)
In each iteration, we also get prediction $r_{k}^{j}$ corresponding to unlabelled data $u_{k}^{j}$ from $j\leavevmode\nobreak\ th$ head. In order to get higher precision pseudo labels, inspired by ensemble learning, the predictions from all heads are aggregated to create a combined prediction $\hat{r}_{k}^{sum}$. Subsequently, this combined prediction is utilized to calculate pseudo labels $\hat{y}_{k}^{sum}$ through an $argmax$ operation, which returns the indices of the maximum values of the prediction along the class dimension. Next, there is a random selector to pick only one head as the main branch to be supervised by minimizing the unsupervised loss
$\mathcal{L}_{unsup}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(r_{k}^{main},\hat{y}_{k}^{sum}\right)\right),$ (3)
where $r_{k}^{main}$ is the prediction from the selected head.
Finally, the whole loss is written as
$\mathcal{L}=\mathcal{L}_{sup}+\lambda\mathcal{L}_{unsup},$ (4)
where $\lambda$ is the trade-off weight between the supervised and unsupervised losses.
Two perturbation methods. Based on the framework of DiverseHead, we propose a parameter perturbation method called dynamic freezing and also apply dropout as another perturbation to increase the diversity of features during training. Specifically, the first step of dynamic freezing refers to a process where, during each iteration, half of the heads are randomly selected to be frozen. This implies that the parameters within these selected heads will remain unchanged and not undergo updates. Every head has an equal probability of being chosen. Before going to the next iteration, those frozen heads are unfrozen. The algorithm of the DiverseHead semi-supervised learning approach with the perturbation of dynamic freezing is expressed in the Algorithm 1.
Another form of perturbation involves the utilization of dropout layers. We introduce a dropout layer before the last convolutional layer for each head. The pseudocode of this method is expressed in Algorithm 2. In the training phase, specific components of the feature tensor are randomly assigned a value of zero with a dropout rate (determining the probability), using samples derived from a Bernoulli distribution. This dropout operation is employed to enhance the variability of the output predictions. In the conducted experiments, the dropout rate is set as a hyperparameter with a value of 0.3. Similar to the training procedure of the semi-supervised learning approach based on dynamic freezing, the dropout-based approach adheres to the same training pipeline outlined above, with the distinction that all heads remain unfrozen. To evaluate the efficacy of individual perturbation methods in conjunction with the proposed DiverseHead techniques, each method is applied independently within the proposed framework. The performance of each combination is assessed in Section V.
Algorithm 1 DiverseHead Semi-Supervised Learning with Dynamic Freezing Pseudocode
  Initialization:
              Initialise backbone using ResNet-50
              Randomly initialise parameters for each head
  INPUT: Labelled Training Dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$
               Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
[…]
        $\left\{head^{i}\right\}_{i=random}^{len(heads)/2}\leftarrow sample\left(\left\{head^{i}\right\}_{i=1}^{len(heads)}\right)$
        $Freeze(\left\{head^{i}\right\}_{i=random}^{len(heads)/2})$
        for $j\in\left\{1,...,len(heads)\right\}$ do
            $p_{k}^{j}=Q^{head_{j}}(x_{k}^{j})$
            $\mathcal{L}_{sup}^{j}=loss(p_{k}^{j},y_{k})$ based on (1)
            $r_{k}^{j}=Q^{head_{j}}(u_{k}^{j})$
        $r_{k}^{sum}=sum\left(\left\{r_{k}\right\}_{k=1}^{len(heads)}\right)$
        $\hat{y}_{k}^{sum}\leftarrow argmax(r_{k}^{sum})$
        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{len(heads)}\right\}\right)$
        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{sum})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{len(heads)}\sum_{j=1}^{len(heads)}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
        Update model to minimize $\mathcal{L}$
        $Unfreeze(\left\{head^{i}\right\}_{i=random}^{len(heads)/2})$
  OUTPUT: Trained Model Q
Algorithm 2 DiverseHead Semi-Supervised Learning with Dropout Pseudocode
  Initialization:
          Initialise backbone using ResNet-50
          Randomly initialise parameters for each head
          Add dropout layer before the last convolutional layer
  INPUT: Labelled Training Dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$
               Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
[…]
        for $j\in\left\{1,...,len(heads)\right\}$ do
            $p_{k}^{j}=Q^{head_{j}}(x_{k}^{j})$
            $\mathcal{L}_{sup}^{j}=loss(p_{k}^{j},y_{k})$ based on (1)
            $r_{k}^{j}=Q^{head_{j}}(u_{k}^{j})$
        $r_{k}^{sum}=sum\left(\left\{r_{k}\right\}_{k=1}^{len(heads)}\right)$
        $\hat{y}_{k}^{sum}\leftarrow argmax(r_{k}^{sum})$
        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{len(heads)}\right\}\right)$
        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{sum})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{len(heads)}\sum_{j=1}^{len(heads)}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
        Update model to minimize $\mathcal{L}$
  OUTPUT: Trained Model Q
III Cross-model Supervision (DiverseModel)
Figure 3: DiverseModel: an online semi-supervised learning approach.
The proposed DiverseModel differs from CPS, exploring different networks in parallel to generate comprehensive pseudo labels. Since different networks pay different and complementary attention to the same input, this offers the basis that they are able to benefit from each other. DiverseModel structure includes three distinct networks which are PSPNet [zhao2017pyramid], UNet [ronneberger2015u], SegNet [badrinarayanan2017segnet], respectively. Algorithm 3 presents the pseudocode of the DiverseModel method.
The labelled data is used in a regular supervised learning manner to train these models by using the standard cross-entropy loss function $\ell_{ce}$. The supervised loss is expressed as:
$\mathcal{L}_{sup}=\sum_{n=1}^{3}\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k}^{n},y_{k}\right)\right),$ (5)
where $p_{k}^{n}$ represents the prediction of $k\leavevmode\nobreak\ th$ pixel from $n\leavevmode\nobreak\ th$ network.
In addition, unlabelled data is used to generate pseudo labels, which are then exploited for cross-supervision to inform each network. Different from the version presented in [ma2023confidence], all loss calculations in this work solely focus on cross-entropy loss to avoid the effects resulting from different types of loss functions on performance. The predictions obtained by each network are denoted as $\{p^{1},p^{2},p^{3}\}$, which are used for generating pseudo labels $\{\hat{r}^{1},\hat{r}^{2},\hat{r}^{3}\}$ through $argmax$ operation. For instance, the cross pseudo supervision loss $\mathcal{L}_{unsup}^{12}$ between the prediction $p^{1}$ from first network and the pseudo label $r^{2}$ generated by the second network is defined as:
$\mathcal{L}_{unsup}^{12}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k}^{1},\hat{r}_{k}^{2}\right)\right).$ (6)
Algorithm 3 DiverseModel Semi-supervised Learning Pseudocode
  Initialization:
               Randomly initialise 3 models, PSPNet $Q^{1}$, UNet $Q^{2}$, SegNet $Q^{3}$
  INPUT: Labelled Training Dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$
               Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
[…]
        $\mathcal{L}_{sup}=loss(Q^{1}(x_{k}),y_{k})$ + $loss(Q^{2}(x_{k}),y_{k})$
                  + $loss(Q^{3}(x_{k}),y_{k})$ based on (1)
        $\hat{r}_{k}^{1}\leftarrow argmax(Q^{3}(u_{k}))$
        $\hat{r}_{k}^{2}\leftarrow argmax(Q^{3}(u_{k}))$
        $\hat{r}_{k}^{3}\leftarrow argmax(Q^{3}(u_{k}))$
        $\mathcal{L}_{unsup}=loss(Q^{1}(u_{k}),\hat{r}_{k}^{2})$ + $loss(Q^{1}(u_{k}),\hat{r}_{k}^{3})$
                  + $loss(Q^{2}(u_{k}),\hat{r}_{k}^{1})$ + $loss(Q^{2}(u_{k}),\hat{r}_{k}^{3})$
                  + $loss(Q^{3}(u_{k}),\hat{r}_{k}^{1})$ + $loss(Q^{3}(u_{k}),\hat{r}_{k}^{2})$
        $\mathcal{L}\leftarrow\mathcal{L}_{sup}+\lambda\mathcal{L}_{unsup}$
        Update model to minimize $\mathcal{L}$
  OUTPUT: Trained Model $Q^{1}$, $Q^{2}$, $Q^{3}$
The cross-pseudo supervision among three networks creates six losses by using the same way. The unsupervised loss $\mathcal{L}_{unsup}$ is the result of linearly adding the six individual losses, as shown below
$\begin{array}[]{r}\mathcal{L}_{unsup}=\mathcal{L}_{unsup}^{12}+\mathcal{L}_{unsup}^{13}+\mathcal{L}_{unsup}^{21}\\ +\mathcal{L}_{unsup}^{23}+\mathcal{L}_{unsup}^{31}+\mathcal{L}_{unsup}^{32}.\end{array}$ (7)
The total loss $\mathcal{L}$ is the linear addition of $\mathcal{L}_{sup}$ and $\mathcal{L}_{unsup}$, which is previously given in equation (4)
IV Dataset Description
We employed diverse datasets to assess both the proposed techniques and state-of-the-art methods, specifically including (1) the ISPRS Potsdam dataset [rottensteiner2012isprs], (2) the DFC2020 dataset [robinson2021global], (3) the RoadNet dataset [liu2018roadnet], and (4) the Massachusetts Buildings dataset [MnihThesis]. In the sequel, we share the details of each datasets utilised in this paper.
1) ISPRS Potsdam Semantic Labeling dataset is an open-access benchmark dataset provided by the International Society for Photogrammetry and Remote Sensing (ISPRS). The ground sampling distance of the two modalities of the true orthophoto (TOP) and the DSM is 5 cm. This dataset was annotated manually into six land cover classes, which are impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background. This dataset provides 38 multi-source patches (all with the size of $6000\times 6000$ pixels), which contain infrared (IR), red, green and blue bands orthorectified optical images with corresponding digital surface models (DSM). For computation purposes, we partitioned all these data tiles into $512\times 512$ patches, resulting in 3456, 201, and 1815 samples for training, validation, and test sets, respectively. We randomly select 1/4 of training data as labelled data and use the remaining 3/4 as unlabelled data for semi-supervised learning.
2) DFC2020 is the 2022 IEEE GRSS Data Fusion Contest dataset which is based on the SEN12MS dataset [schmitt2019sen12ms]. It provides Sentinel-1 SAR imagery, Sentinel-2 multispectral imagery, and corresponding land cover maps across seven areas namely Mexico City, Kippa-Ring, Khabarovsk, Black Forest, Mumbai, Cape Town, Bandar Anzali between 2016 and 2017. The size of all patches is $256\times 256$ pixels. There are 6112, 986 and 5127 images for training, validation and test sets, respectively. In this paper, we utilize only one-fifth of the labelled data for training, while the remaining four-fifths are employed as unlabeled data. The fine-grained International Geosphere-Biosphere Program (IGBP) classification scheme in SEN12MS was aggregated to ten coarser-grained classes, which are forest (11.3%), shrubland (6.9%), savanna (23.6%), grassland (16.8%), wetlands (1.1%), croplands (17.9%), urban/built-up (10.6%), snow/ice (0.0%), barren (5.2%), and water (6.5%).
3) RoadNet is a benchmark dataset for road detection with 0.21-m spatial resolution per pixel covering 21 typical urban regions about 8 $km^{2}$ in Ottawa, Canada. The annotation data includes road surfaces, road edges, and road centerlines for each image. we only used road surfaces for the segmentation task. the number of samples for training, validation and testing is 410, 45 and 387, respectively. In the semi-supervised learning approaches, only 1/4 of the annotated training data is used, while the remaining 3/4 of the training data are employed as unlabeled data for training purposes.
4) Massachusetts Buildings Dataset predominantly encompasses urban and suburban regions, encompassing structures of varying scales. This includes diverse buildings such as individual houses and garages, all of which are included in the labels. The dataset consists of 151 aerial images capturing the Boston area covering roughly 340 square kilometres, with a resolution of $1\leavevmode\nobreak\ m^{2}$ per pixel. The dataset was divided into subsets as follows: 137 images for training, 10 images for testing, and 4 images for validation. Each subset consists of images sized $1500\times 1500$ pixels. One-quarter of the annotated training images and the remaining three-quarters of the unlabelled training data are utilized in semi-supervised learning.
V experiment and results analysis
V-A Implementation Details
We implemented our methods by using the PyTorch framework. Following [chen2018encoder], we used a mini-batch SGD optimizer and adopted a poly learning rate policy, where the current learning rate equals the initial learning rate multiplied by $\left(1-\frac{\text{ iter }}{\text{ max-iter }}\right)^{\text{power }}$. The initial learning rate and power are set to 0.01 and 0.9, respectively. All experiments were conducted on the GW4 Isambard with an NVIDIA A100-sxm GPU and an AMD EPYC 7543P CPU  [gw4].
We implemented various classic semi-supervised learning methods (MT, CCT, GCT and CPS) on remote sensing datasets. All these methods use DeepLabv3+ with ResNet-50 pre-trained on ImageNet as the semantic segmentation network. To be fair the proposed DiverseHead is also based on DeepLabv3+ with the pretrained ResNet-50.
We comprehensively analyzed all methods by quantifying performance via class-related measures, including overall accuracy (OA), user’s accuracy (UA), producer’s accuracy (PA), mean intersection over union (mIoU), and $F_{1}$-score. Expressions of all five performance metrics are given as follows: $\mathrm{OA}=\frac{TP+TN}{TP+TN+FP+FN}$, $\mathrm{UA}=\frac{TP}{TP+FP}$, $\mathrm{PA}=\frac{TP}{TP+FN}$, $\mathrm{mIoU}=\frac{|\mathrm{TP}|}{|\mathrm{TP}+\mathrm{FN}+\mathrm{FP}|}$, $F_{1}=\frac{2\cdot PA\cdot UA}{PA+UA}$, where TP, TN, FP, and FN refer to the numbers of pixels that are true positives, true negatives, false positives, and false negatives for each class, respectively.
V-B Quantitative Results and Analysis
We evaluated the performance of the proposed approach and reference networks using 4 aforementioned datasets, and the average results are presented in Table I. The reason we first share average results across datasets is to provide a global applicability display for all utilised techniques, and each data set’s performance will be discussed in the sequel. Evaluating the performance results in Table I, we conclude that the proposed DiverseModel achieves the best performance among all metrics. In particular, for metrics like UA and mIoU, the proposed DiverseModel exhibits an improvement of over 5% compared to another cross-model-based approach, CPS. Although the performance of DiverseHead is lower than that of DiverseModel, it outperforms other reference methods and reaches the second best. Upon convergence of contemporary methodologies with remote sensing datasets, MT demonstrates notably inferior average performance across various metrics, with the exception of UA. Both CCT and GCT exhibit comparable performance, yet the performance superiority of CCT is notably evident in its PA metric. CPS consistently yields superior outcomes across the majority of metrics when juxtaposed against other aforementioned state-of-the-art methodologies.
Meanwhile, it is worth noting that the DiverseHead consists of a notably lighter architecture compared to all other models where Table II presents the required number of parameters during training for each semi-supervised learning architecture. It is evident that the architectures using multiple segmentation networks during training typically need higher memory demands. In contrast, the proposed DiverseHead is remarkably lightweight as it only employs a single model with multiple heads which consists of only 2 convolution layers, eliminating the need for multiple models during training. Thus, the parameter size of the DiverseHead (DT&DF) is only 16% bigger than that of the single network (Base in Table II), whereas the parameter size of other reference semi-supervised architectures is at least twice that of the single network. Despite this need for a lower number of parameters, the DiverseHead proposed method outperforms all state-of-the-art with at least 1% accuracy whilst obtaining similar performance to the companion method of DiverseModel.
Table I: Average performance comparison with the state of the arts on 4 datasets. DT and DF present Dropout and dynamic freezing, respectively.
OA UA PA mIoU $F_{1}$
MT[tarvainen2017mean] 85.63% 73.14% 80.10% 63.93% 76.39%
CCT[ouali2020semi] 86.60% 72.17% 83.29% 64.07% 77.22%
GCT[ke2020guided] 86.53% 73.33% 80.50% 64.59% 76.72%
CPS[chen2021semi] 87.17% 72.89% 84.62% 65.02% 78.18%
DiverseHead (DT) 88.42% 77.80% 84.95% 69.70% 81.05%
DiverseHead (DF) 88.29% 76.98% 84.87% 68.85% 80.58%
DiverseModel 88.86% 78.24% 86.06% 70.53% 81.79%
Table II: The required parameter size in each semi-supervised learning approach. Base means supervised segmentation network DeepLabv3+. DT and DF present Dropout and dynamic freezing, respectively.
CPS [chen2021semi] MT [tarvainen2017mean] CCT [ouali2020semi] GCT [ke2020guided]
Size (MB) 303.378 303.378 337.655 335.049
Base DiverseHead(DT) DiverseHead(DF) DiverseModel
Size (MB) 151.689 175.806 175.806 936.262
Specifically, we present the results of each semi-supervised learning method for the 4 utilised datasets in Table III. DiverseNet architectures reach a remarkable performance for each of the datasets. Specifically, the proposed DiverseModel achieves the best performance on both RoadNet and Massachusetts building datasets for all metrics, and the DiverseHead (DT) shows the second best for the most of metrics on these two datasets. For the other two datasets, Potsdam and DFC2020, The DiverseHead (DF) method shows competitive and better performance compared to the companion DiverseModel for most of the metrics.
Table III: Performance comparison with the state of the arts on 4 datasets. DT and DF present Dropout and dynamic freezing, respectively.
OA UA PA mIoU $F_{1}$
Potsdam MT[tarvainen2017mean] 81.98% 73.66% 78.39% 63.07% 75.95%
CCT[ouali2020semi] 82.66% 74.64% 77.62% 64.16% 76.10%
GCT[ke2020guided] 83.99% 75.65% 80.81% 65.80% 78.14%
CPS[chen2021semi] 85.00% 75.76% 82.94% 66.69% 79.19%
DiverseHead (DT) 84.60% 77.42% 80.49% 67.18% 78.92%
DiverseHead (DF) 85.37% 78.43% 81.86% 68.49% 80.11%
DiverseModel 85.76% 76.75% 83.45% 67.85% 79.96%
DFC2020 MT[tarvainen2017mean] 78.64% 59.59% 73.57% 50.44% 65.85%
CCT[ouali2020semi] 79.71% 59.87% 76.40% 51.04% 67.13%
GCT[ke2020guided] 80.84% 61.47% 71.43% 52.17% 66.07%
CPS[chen2021semi] 81.50% 61.75% 79.47% 53.21% 69.50%
DiverseHead (DT) 82.11% 62.31% 80.38% 53.93% 70.20%
DiverseHead (DF) 82.19% 62.29% 80.73% 54.02% 70.32%
DiverseModel 81.87% 62.20% 80.69% 53.71% 70.25%
RoadNet MT[tarvainen2017mean] 93.00% 80.74% 85.54% 72.96% 83.07%
CCT[ouali2020semi] 93.90% 83.85% 87.26% 76.23% 85.52%
GCT[ke2020guided] 93.57% 79.26% 89.25% 73.51% 83.96%
CPS[chen2021semi] 94.36% 84.21% 88.85% 77.44% 86.47%
DiverseHead (DT) 96.28% 90.89% 91.95% 84.84% 91.42%
DiverseHead (DF) 95.36% 87.99% 90.31% 81.35% 89.13%
DiverseModel 96.79% 92.82% 92.57% 86.87% 92.70%
Massachusetts MT[tarvainen2017mean] 88.92% 78.56% 82.90% 69.25% 80.67%
CCT[ouali2020semi] 89.65% 79.91% 84.23% 70.95% 82.01%
GCT[ke2020guided] 87.73% 76.95% 80.51% 66.88% 78.69%
CPS[chen2021semi] 87.85% 69.85% 87.24% 62.76% 77.58%
DiverseHead (DT) 90.69% 80.56% 86.97% 72.86% 83.64%
DiverseHead (DF) 90.25% 79.22% 86.59% 71.53% 82.74%
DiverseModel 91.01% 81.19% 87.53% 73.67% 84.24%
We also considered a downgraded version of the current DiverseHead approach, called single-head supervision (SHS). The segmentation network consists of only a single head. During each iteration, the current model’s predictions for unlabeled data are utilized to generate pseudo labels through an $\arg\max$ operation. These generated pseudo labels are then used to calculate the unsupervised loss. The baseline is the segmentation network only trained by labelled data. The performance of the baseline and SHS, along with the proposed DiverseHead, are presented in Table IV in order to provide experimental evidence of the novelty of utilisation of multiple heads in DiverseHead architecture. While the single-head supervision benefits from the pseudo labels and exhibits better performance compared to the baseline model especially improving the metric of PA, both versions of DiverseHead approaches demonstrate a further significant improvement by taking advantage of multiple heads. Specifically, the mIoU of DiverseHead (DF) is 5.7% higher than that of Baseline and approximately 4% higher than that of SHS for Potsdam. Also, the mIoU value of DiverseHead (DT) is about 10% higher than that of SHS.
Table IV: performance comparison with single-head supervision (SHS) and baseline model which is only supervised by labelled data. DT and DF present Dropout and dynamic freezing, respectively.
OA UA PA mIoU $F_{1}$
Potsdam Baseline 81.64% 73.86% 76.05% 62.72% 74.94%
SHS 83.43% 74.57% 79.93% 64.51% 77.16%
DiverseHead (DT) 84.60% 77.42% 80.49% 67.18% 78.92%
DiverseHead (DF) 85.37% 78.43% 81.86% 68.49% 80.11%
RoadNet Baseline 93.56% 82.85% 86.53% 75.06% 84.65%
SHS 93.79% 81.19% 88.62% 74.87% 84.74%
DiverseHead (DT) 96.28% 90.89% 91.95% 84.84% 91.42%
DiverseHead (DF) 95.36% 87.99% 90.31% 81.35% 89.13%
Table V: Ablation study about the number of heads for DiverseHead with dynamic freezing
# of heads OA UA PA mIoU $F_{1}$
Potsdam 8 85.70% 78.47% 82.09% 68.82% 80.24%
10 85.37% 78.43% 81.86% 68.49% 80.11%
12 85.51% 77.23% 83.03% 68.14% 80.02%
RoadNet 8 94.28% 80.20% 92.23% 75.59% 85.79%
10 95.36% 87.99% 90.31% 81.35% 89.13%
12 95.35% 85.57% 92.22% 80.52% 88.77%
Average 8 89.99% 79.33% 87.16% 72.21% 83.02%
10 90.36% 83.21% 86.08% 74.92% 84.62%
12 90.43% 81.40% 87.62% 74.33% 84.40%
An ablation study was conducted on the Potsdam and RoadNet datasets to explore the impact of the number of heads for the proposed DiverseHead method with dynamic freezing perturbation. The results in Table V indicate that using 10 heads yields the optimal choice across most metrics in terms of average performance. The performance difference between the versions with different numbers of heads is very small on Potsdam, but the version of 8 heads shows a performance drop. Considering the stability, We chose the 10-head version of DiverseHead (DF) for all experiments in this paper.
Table VI: Performance comparison of the DiverseModel with its constituent networks
OA UA PA mIoU $F_{1}$
UNet 83.28% 74.18% 78.88% 64.51% 76.46%
SegNet 82.37% 73.53% 77.46% 63.24% 75.45%
PSPNet 81.96% 74.23% 76.99% 63.08% 75.58%
DiverseModel 85.76% 76.75% 83.45% 67.85% 79.96%
Since the DiverseModel technique uses three different segmentation networks, we also evaluated the performance of each individual member network within the DiverseModel in Table VI. The results indicate that the cross-supervision of all networks shows a significant improvement in the segmentation task performance compared to each separate network.
Figure 4: Visual results of each method on Potsdam Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
Figure 5: Visual results of each method on RoadNet Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
Figure 6: Visual results of each method on DFC2020 Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
Figure 7: Visual results of each method on Massachusetts Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
V-C Qualitative Results and Analysis
The visible results are presented in Figures from 4 to 7. For each dataset, we chose one case and depicted the RGB image, ground truth, and predictions of all the methods. Compared to the ground truth, the proposed three methods (right-top 3 sub-figures) achieve better similarity to the ground truth compared to the reference models. We also calculated the IoU scores for all predictions with the ground truth for this case, which are shown in the sub-caption of each prediction. The IoU scores also demonstrate that the performance of the proposed DiverseNet methods is more competitive than that of the other approaches.
VI Conclusion
In this paper, we proposed a lightweight semi-supervised learning based on a multi-head structure called DiverseHead, which equipped a dynamic freezing and dropout perturbation approaches. The multi-head idea benefits from bagging and shows competitive performance for the segmentation of remote sensing imagery. Also, another high-efficiency multiple network based semi-supervised learning method called DiverseModel is investigated for further analysis. DiverModel shows the strongest performance on various datasets compared to the state of the arts but requires more computational resources, whilst the performance of DiverseHead is slightly lower than DiverseModel but dramatically lighter than all reference models.

### v2

@@
 Abstract
-Semi-supervised learning is designed to help reduce the cost of the manual labelling process by exploiting the utilisation of useful features from a large quantity of unlabelled data during training.
+Semi-supervised learning aims to help reduce the cost of the manual labelling process by leveraging valuable features extracted from a substantial pool of unlabeled data alongside a limited set of labelled data during the training phase.
 Since pixel-level manual labelling in large-scale remote sensing imagery is expensive, semi-supervised learning becomes an appropriate solution to this.
-However, most of the existing semi-supervised learning methods still lack efficient perturbation methods to promote the precision and diversity of features (or pseudo labels) during training.
-In order to fill this gap, we propose DiverseNet architectures which explore multi-head and multi-model semi-supervised learning algorithms by simultaneously promoting precision and diversity during training.
-The two proposed methods of DiverseNet, namely the DiverseHead and DiverseModel, significantly achieve the highest semantic segmentation performance in four widely utilised remote sensing imagery data sets compared to state-of-the-art semi-supervised learning methods.
-Meanwhile, the proposed DiverseHead architecture is relatively lightweight in terms of parameter space compared to the state-of-the-art methods whilst reaching high-performance results for all utilised data sets.
-Index Terms:
-Semi-supervised Learning, Consistency Learning, Land Cover Classification
-I Introduction
+However, most of the existing consistency learning frameworks based on network perturbation are very bulky.
+There is still a lack of lightweight and efficient perturbation methods to promote the diversity of features and the precision of pseudo labels during training.
+In order to fill this gap, we propose DiverseNet which explores multi-head and multi-model semi-supervised learning algorithms by simultaneously enhancing precision and diversity during training.
+The two proposed methods in the DiverseNet family, namely DiverseHead and DiverseModel, both achieve the better semantic segmentation performance in four widely utilised remote sensing imagery data sets compared to state-of-the-art semi-supervised learning methods.
+Meanwhile, the proposed DiverseHead architecture is simple and relatively lightweight in terms of parameter space compared to the state-of-the-art methods whilst reaching high-performance results for all the tested data sets.
+keywords:
+Semi-supervised Learning, Consistency Learning, Semantic Segmentation
+[…]
+organization=School of Computer Science and Informatics, Cardiff University,city=Cardiff, postcode=CF24 4HJ, country=United Kingdom
+1 Introduction
 Supervised deep learning has gradually become the dominant technique in computer vision, especially during the last decade.
-Specifically, with its great success in semantic segmentation, automated land cover classification approaches have been significantly improved because of the high similarity between the two tasks [vali2020deep].
-Nevertheless, supervised learning necessitates a substantial, meticulously labelled dataset.
-In the case of extensive remote sensing data, such as satellite imagery and drone-captured images in complex terrains, acquiring pixel-wise expert annotations proves to be a time-consuming, labour-intensive, and costly endeavour.
-While the field of computer vision boasts numerous well-annotated datasets, transferring deep learning models trained on these datasets to the remote sensing domain is a formidable challenge.
-This is primarily because of the significant disparities between natural images and remote sensing data, including hyperspectral and synthetic aperture radar (SAR) imagery, which often possesses unconventional image geometries, making the process non-intuitive.
-Fortunately, thanks to open-access remote sensing sources, a large amount of unlabelled imagery can be accessed easily and freely.
-Thus, only using a limited amount of labelled data and exploiting the benefits of a large amount of unlabelled data could be a feasible solution for solving the problem of the lack of labelled remote sensing datasets [wang2022semi].
-In computer vision, semi-supervised learning has become a trending research direction to eliminate the labour-intensive and expensive annotation stages mentioned above.
-It leads to competitive performance compared to supervised learning in applications such as image classification [zhang2021flexmatch, huang2021embedding, wang2021semi2] and segmentation [hu2021semi, wang2022semi_u2pl, wang2021semi], which just use a part of labelled data and a large amount of unlabelled data.
+Specifically, with its great success in semantic segmentation in computer vision, automated land cover classification approaches have been significantly improved  (Vali et al., 2020).
+Nevertheless, supervised learning necessitates a substantial and meticulously labelled dataset.
+In the case of extensive remote sensing data, such as satellite imagery and drone-captured images in complex terrains, acquiring pixel-wise expert annotations is a time-consuming, labour-intensive, and costly endeavour.
+While the field of computer vision provides numerous well-annotated datasets, transferring deep learning models trained on these datasets to the remote sensing domain is a formidable challenge.
+This is primarily due to the significant differences between images generally used in computer vision and remote sensing data, including hyperspectral and synthetic aperture radar (SAR) imagery, in which images often possess unconventional and non-intuitive characteristics.
+Fortunately, thanks to open-access remote sensing data sources, a large amount of unlabelled imagery can be accessed easily and freely.
+Thus, only using a limited amount of labelled data and exploiting the benefits of a large amount of unlabelled data could be a feasible solution for solving the problem of the lack of labelled remote sensing datasets (Wang et al., 2022a; Yang et al., 2023; Wang et al., 2023).
+In computer vision, semi-supervised learning has become a popular research direction to eliminate the labour-intensive and expensive annotation stages mentioned above.
+It leads to competitive performance compared to supervised learning in applications such as image classification (Zhang et al., 2021; Huang et al., 2021; Wang et al., 2021b) and segmentation (Hu et al., 2021; Wang et al., 2022b, 2021a), which exploit just a part of labelled data and a large amount of unlabelled data.
 Specifically, mainstream semi-supervised learning methods exploit “pseudo” labels that are fake and/or auto-generated label information obtained as a result of an unsupervised step.
@@
 However, the accuracy of pseudo labels is not always satisfactory especially when labelled training data is not complete and sufficient enough.
-In this case, augmenting the pseudo-labels to increase diversity has emerged as another key research direction within the semi-supervised learning domain.
-A greater diversity aids in constructing more comprehensive pseudo-labels, which, in turn, indirectly enhances the quality of these labels and the training effectiveness of the model.
+Meanwhile, increasing diversity of the pseudo-labels has become another key research direction within the semi-supervised learning domain, especially for consistency regularization (French et al., 2019).
+This is because a greater diversity of pseudo-labels enhances the effectiveness of training deep learning networks.
+In order to empower networks to harness the potentials within unlabeled data, a notable technique known as consistency regularization has emerged as a widely adopted method for semi-supervised learning.
+(French et al., 2019; Filipiak et al., 2021).
+Specifically, consistency regularization methods are built up on the theory of assumption of smoothness, which posits that if two points reside in a high-density region of feature space and are close to each other, their corresponding labels should be the same or consistent (Chen and Wang, 2010; Luo et al., 2018; French et al., 2019).
+In practice, these methods for consistency regularization ensure adherence to this assumption by compelling networks to generate predictions that remain consistent across modified iterations of unlabeled inputs.
+The strategies for the aforementioned perturbations can be categorised into three groups, namely input, feature, and network perturbations.
+Input perturbation is a method that involves making adjustments or alterations to the input data as part of the training process.
+In this way, the semi-supervised learning approach promotes the consistency of predictions made in response to these modified inputs.
+Widely used input perturbation methods employed in semi-supervised learning include Cutout (DeVries and Taylor, 2017), MixUp (Zhang et al., 2017), and CutMix (Yun et al., 2019).
+Among them, CutMix stands out by combining aspects of MixUp and CutOut techniques, demonstrating superior performance compared to them (French et al., 2019).
+However, input perturbation introduces artificial noise to the data, which might lead to incorrect or noisy labels when generating pseudo-labels for unlabeled examples.
+This might negatively impact the training process by providing incorrect guidance to the model when using pseudo labels.
+Cross-consistency training (CCT) (Ouali et al., 2020) presents a feature perturbation method that introduces noise to both low- and high-level features.
+These perturbed features are then fed into multiple decoders to generate multiple outputs, followed by the enforcement of consistency among the outputs obtained from different decoders.
+However, feature perturbation alters the intrinsic characteristics and relationships between features, which also might lead to the creation of artificial feature representations that do not accurately reflect the underlying patterns in the original data.
+Consequently, the perturbed features could result in incorrect or noisy labels when generating pseudo-labels for unlabeled examples, which negatively impacts the learning process.
+Lastly, network perturbation uses different networks or the same network with different initialization weights to promote diversity of predictions, such as cross pseudo supervision (CPS) (Chen et al., 2021), confidence-guided semi-supervised learning (CGSSL) (Ma et al., 2023), mean teacher (MT), (Tarvainen and Valpola, 2017), CCT (Ouali et al., 2020), guided collaborative training (GCT) (Ke et al., 2020) and cross teaching between CNN and transformer (Luo et al., 2022).
+Compared to the previous two perturbation methods, although network perturbation methods introduce perturbations created by the model itself instead of artificial noise, they generally require rapidly increased computational resources due to the increased number of networks (whole or their inner stages) that are used.
 (a) DiverseHead
 (b) DiverseModel
-Figure 1: Two kinds of Pseudo label generation methods based on multiple heads (a) DiverseHead and multiple models (b) DiverseModel.
-In order to empower networks to harness the unexplored potentials within unlabeled data, a notable technique known as consistency regularization has emerged as a widely adopted method for semi-supervised learning.
-[french2019semi, filipiak2021n].
-Specifically, consistency regularization methods rely on the assumption of smoothing, which posits that if two points reside in a high-density region and are close to each other, their corresponding labels should be the same or consistent [chen2010semi].
-In practice, these consistency regularization methods achieve this assumption by forcing networks to produce consistent predictions for unlabeled inputs that undergo diverse perturbations.
-The strategies for the above-mentioned perturbations can be categorised into three groups, namely input, feature, and network perturbations.
-Input perturbation is a method that involves making adjustments or alterations to the input data as part of the training process.
-In this way, the semi-supervised learning approach promotes the consistency of predictions made in response to these modified inputs.
-Widely used input perturbation methods employed in semi-supervised learning include Cutout[devries2017improved], MixUp [zhang2017mixup], and CutMix [yun2019cutmix].
-Among them, CutMix stands out by combining aspects of MixUp and CutOut techniques, demonstrating superior performance compared to them [french2019semi].
-However, input perturbation introduces artificial noise to the data, which might lead to incorrect or noisy labels when generating pseudo-labels for unlabeled examples.
-This might negatively impact the training process by providing incorrect guidance to the model when using pseudo labels.
-Cross-consistency training (CCT) [ouali2020semi] presents a feature perturbation method that introduces noise to both low- and high-level features.
-These perturbed features are then fed into multi-decoders to generate multiple outputs, followed by the enforcement of consistency among the outputs obtained from different decoders.
-However, feature perturbation alters the intrinsic characteristics and relationships between features, which also might lead to the creation of artificial feature representations that do not accurately reflect the underlying patterns in the original data.
-Consequently, the perturbed features could result in incorrect or noisy labels when generating pseudo-labels for unlabeled examples, which negatively impacts the learning process.
-Lastly, the network perturbation uses different networks or the same network with different initialization weights to promote diversity of predictions, such as cross pseudo supervision (CPS)[chen2021semi], confidence-guided semi-supervised learning (CGSSL)[ma2023confidence], mean teacher (MT), [tarvainen2017mean], CCT [ouali2020semi], guided collaborative training (GCT) [ke2020guided] and cross teaching between CNN and transformer [luo2022semi].
-Compared to the previous two perturbation methods, although network perturbation methods introduce perturbations created by the model itself instead of artificial noise, they generally require rapidly increased computational resources due to the increased number of networks (whole or their inner stages) utilized.
-I-A Contributions
-The preceding conversation underscores that, although perturbation techniques have their advantages, they also bring inherent drawbacks.
-Therefore, the creation of methods based on perturbations is a complex task, necessitating the exploration of more sophisticated approaches to augment the potential of semi-supervised learning frameworks by incorporating advanced perturbation methods.
-This paper proposes perturbation-based semi-supervised network architectures, coined as DiverseNet, which consist of two novel feature and decision diversifying network architectures for the purposes of semi-supervised semantic segmentation of remote sensing imagery.
-First, we develop a lightweight semi-supervised semantic segmentation network called DiverseHead based on multiple decision heads to promote precision and diversity of pseudo labels whilst using two novel perturbation methods of dropout (DT) and dynamic freezing (DF).
-In addition, we also propose a high-performance multi-model structure called DiverseModel for scenarios equipped with high-memory computational resources.
+[…]
+1.1 Contributions
+The preceding overview underscores that, although previous perturbation techniques in semi-supervised learning have their advantages, they also bring inherent drawbacks which could be a key consideration affecting the performance in semi-supervised learning.
+In addressing the challenges posed by perturbation-based semi-supervised learning and its complexity, it becomes imperative to explore more sophisticated and efficient approaches.
+These approaches aim to enhance the capabilities of semi-supervised learning frameworks by integrating advanced perturbation methods.
+This paper proposes two perturbation-based semi-supervised network architectures, coined as DiverseNet, which consist of decision (DiverseHead) and novel feature (DiverseModel) diversifying network architectures for the purposes of semi-supervised semantic segmentation of remote sensing imagery.
+The proposed lightweight semi-supervised semantic segmentation network called DiverseHead is shown in Figure 1-(a) based on multiple decision heads to promote precision and diversity of pseudo labels whilst using two novel perturbation methods of dropout (DT) and dynamic freezing (DF).
+In addition, we also analysed a previously proposed high-performance multi-model structure called DiverseModel (Ma et al., 2023) as shown in Figure 1-(b) for scenarios equipped with high-memory computational resources.
 Specifically, the contributions of this work are as follows:
   1. 1.
-    We promote utilisation of multiple decision heads to provide a semi-supervised learning structure, called DiverseHead.
+    We proposed a simple but efficient semi-supervised learning architecture, promoting the utilisation of multiple decision heads called DiverseHead.
 This structure is inspired by bagging (also called bootstrap aggregating) and provides diversity for parameters and features.
-The main source of perturbation is from the dynamic choosing heads to be updated for each iteration.
-This architecture does not distinguish the main or auxiliary decoder.
-Also, all decoders are trained by both supervised and unsupervised losses, which comes with the advantages of high-quality perturbation and diversity of semi-supervised segmentation tasks for remote sensing imagery, whilst providing a relatively lightweight architecture.
+Since each head only consists of a couple of convolutional layers instead of a whole network, the architecture is relatively lightweight and adaptable to various segmentation networks.
   2. 2.
-    For the purposes of creating perturbation in diversifying decisions, we ustilised two important techniques, namely ’dynamic freezing’ and ’dropout’.
-We make these techniques fit with DiverseHead semi-supervised learning architecture for the purpose of diversifying the parameters of the network.
-The two perturbation methods differ from input/feature perturbation methods, in which the source of feature diversity is from parameter diversity instead of introducing artificial noise to data or features.
+    In order to provide efficient pseudo labels in the training stage, we provide a comprehensive voting mechanism based on the proposed DiverseHead structure.
+The voting mechanism not only uses the joint output from multiple heads but also individually considers pseudo labels generated from these outputs, which are called Mean Voting and Max Voting in this paper, respectively.
   3. 3.
-    A previously proposed architecture DiverseModel [ma2023confidence] has been extended for detailed evaluation and in a detailed comparison study for various semantic segmentation data sets.
-This architecture is based on consistency learning among multiple neural networks with distinct weights.
-The rest of the paper is organised as follows: subsection I-B discusses the related work on semi-supervised semantic segmentat ion in remote sensing and some basic knowledge on ensemble machine learning whilst in Section II and III, the proposed algorithms DiverseHead and DiverseModel are presented.
-Section IV describes the segmentation dataset of remote sensing imagery.
-The experimental setting and the qualitative and quantitative analyses of the experimental results are performed in Section V.
-Section VI concludes the paper with a summary.
-I-B Related Work
-Semantic segmentation is one of the well-studied and wide-spread research areas in remote sensing society, for applications such as land cover classification/mapping [dong2019land], building change detection [zheng2022hfa], road extraction [ghandorh2022semantic], and marine debris detection [booth2023high].
-Also, the number of research on semantic segmentation in remote sensing in the literature is rapidly increasing with the success of deep learning in computer vision, due to the strong task-similarity between two areas.
-Fully convolutional networks (FCNs) [long2015fully] is one group of the most widespread networks for segmentation tasks.
+    For the purposes of creating perturbation in diversifying decisions, we used two important techniques, namely ‘dynamic freezing’ and ‘dropout’.
+We make these techniques fit with the DiverseHead semi-supervised learning architecture to diversify the parameters of the network.
+The two perturbation methods differ from input/feature perturbation methods since the source of feature diversity is from parameter diversity instead of introducing artificial noise to data or features.
+  4. 4.
+    Further analysis and more detailed evaluation has been carried out on a previously proposed architecture DiverseModel (Ma et al., 2023) which is based on consistency learning among multiple neural networks with distinct weights.
+We provide a detailed comparison study for various semantic segmentation datasets in this paper.
+Also, we use Grad-CAM (Selvaraju et al., 2020) to verify the observation that different networks exhibit varied attention to the same input, culminating in the combined pseudo label derived from the multiple model predictions showcasing superior quality.
+The rest of the paper is organised as follows: Section 2 discusses the related work on semi-supervised semantic segmentation in remote sensing and some basic knowledge on ensemble machine learning whilst in Section 3 and 4, the proposed algorithms DiverseHead and DiverseModel are presented.
+Section 5 describes the utilised segmentation dataset of remote sensing imagery.
+The experimental setting and the qualitative and quantitative analyses of the experimental results are performed in Section 6.
+Section 7 concludes the paper with a summary.
+2 Related Work
+Semantic segmentation is a well-studied and wide-spread research topic in remote sensing, incorporated within applications such as land cover classification/mapping (Dong et al., 2019), building change detection (Zheng et al., 2022), road extraction (Ghandorh et al., 2022), and marine debris detection (Kikaki et al., 2022; Booth et al., 2023).
+The volume of research on semantic segmentation in remote sensing in the literature is rapidly increasing with the success of deep learning in computer vision, due to the strong task-similarity between two areas.
+Fully convolutional networks (FCNs) (Long et al., 2015) is one group of the most widespread networks for segmentation tasks.
 They have made a considerable contribution to various segmentation tasks either in remote sensing or computer vision.
-Following the FCNs’ success, SegNet [badrinarayanan2017segnet] and UNet [ronneberger2015u] adopt a symmetrical encoder-decoder structure with skip connections, leveraging multi-stage features within the encoder.
-Alternatively, PSPNet [zhao2017pyramid] introduced a pyramid pooling structure that helps provide a global contextual understanding for pixel-level scene parsing.
-DeepLab architecture [chen2017deeplab] proposed a new type of convolution called atrous convolution and an atrous spatial pyramid pooling (ASPP) operation to enable the network to have the ability to control the spatial perception range of convolution kernel via setting different dilation rates.
-DeepLab has then been extended to DeepLabv3+ [chen2018encoder] as an improved hybrid approach that combines an encoder-decoder structure and the ASPP.
-Within DeepLabv3+, the ASPP module is incorporated following a ResNet backbone, enabling the network to delve into deeper layers and extract high-level features, which is helpful for accurately segmenting pixels around boundaries[xia2021cloud, zhang2019multi].
+Following the FCNs’ success, SegNet (Badrinarayanan et al., 2017) and UNet (Ronneberger et al., 2015) adopt a symmetrical encoder-decoder structure with skip connections, leveraging multi-stage features within the encoder.
+Alternatively, PSPNet (Zhao et al., 2017) introduces a pyramid pooling structure that helps provide a global contextual understanding for pixel-level scene parsing.
+DeepLab architecture (Chen et al., 2017) proposes a new type of convolution called atrous convolution and an atrous spatial pyramid pooling (ASPP) operation to enable the network to have the ability to control the spatial perception range of convolution kernels by setting different dilation rates.
+DeepLab has been extended to DeepLabv3+ (Chen et al., 2018) with an improved hybrid approach that combines an encoder-decoder structure and the ASPP.
+Within DeepLabv3+, the ASPP module is incorporated following a ResNet backbone, enabling the network to delve into deeper layers and extract high-level features, which is helpful for accurately segmenting pixels around boundaries (Xia et al., 2021; Zhang et al., 2019).
 DeepLabv3+ is one of the most used networks in the literature for semi-supervised learning segmentation in the computer vision area.
 Semi-supervised learning aims to alleviate the need for expensive annotation work and improve the models’ performances by making use of unlabelled data.
-Self-training [lee2013pseudo] (also known as pseudo labelling) represents one of the primitive semi-supervised learning strategies for both classification [kim2020distribution] and segmentation [zhu2021improving] purposes that involves the generation of pseudo labels through model predictions.
-These pseudo-labels are subsequently utilized to retrain the model.
+Self-training (Lee et al., 2013) (also known as pseudo labelling) represents one of the primitive semi-supervised learning strategies for both classifications (Kim et al., 2020) and segmentation (Zhu et al., 2021), and involves the generation of pseudo labels through model predictions.
+These pseudo-labels are subsequently used to retrain the model.
 Various variations of self-training have been proposed focusing on determining the appropriate pseudo labels.
-Another widely developed semi-supervised learning approach called consistency regularization [french2019semi] is to force networks to give consistent predictions for unlabeled inputs that undergo diverse perturbations.
-CCT [ouali2020semi] employs an encoder-decoder architecture with multiple auxiliary decoders.
+Another widely developed semi-supervised learning approach called consistency regularization (French et al., 2019) is to force networks to give consistent predictions for unlabeled inputs that undergo diverse perturbations.
+CCT (Ouali et al., 2020) employs an encoder-decoder architecture with multiple auxiliary decoders.
 These decoders introduce diversity in the output by feature perturbations specific to each auxiliary decoder.
@@
 It is worth noting that the unsupervised loss is not used to supervise the main decoder.
-Following CCT, subsequent methods like GCT [ke2020guided] and CPS [chen2021semi] are proposed to introduce network perturbation for consistency regularization.
+Following CCT, subsequent methods like GCT (Ke et al., 2020) and CPS (Chen et al., 2021) have been proposed to introduce network perturbation for consistency regularization.
 Specifically, they use two same network structures but with different weight initialization.
 CPS differs from GCT by employing pseudo labels to enforce consistency, whereas GCT achieves consistency regularization by utilizing predictions of networks cooperating with a flow detector.
-Although the network perturbation achieved through different weight initialization in the previous two works shows a good ability to generate diversity, we believe that employing different networks could yield even greater diversity.
-Ensemble machine learning is a concept that employs multiple learners and combines their predictions [sewell2008ensemble].
-Bagging, a form of ensemble learning, is a technique aimed at reducing prediction variance by creating multiple iterations of a predictor and then utilizing them to form an aggregated predictor [breiman1996bagging].
-Specifically, Bagging creates sample subsets by randomly selecting from the training dataset and subsequently utilizes these acquired subsets to train the foundational models for integration.
-When forecasting a numerical result, the aggregation involves taking an average of the versions, while predicting a class, it relies on a majority vote.
-Based on the concept, bagging is a commonly employed approach for enhancing the robustness and precision of machine learning algorithms in terms of classification and regression applications [ren2016ensemble].
-II Cross-head Supervision (DiverseHead)
+Although the network perturbation was achieved through different weight initialization in GCT and CPS, the ability to generate diversity is still limited.
+ICNet (Wang et al., 2022a) was proposed to use teacher networks for increasing the model difference to enhance the diversity of pseudo labels based on an iterative contrast network.
+Nonetheless, these methods consistently rely on the same network architecture.
+To introduce more diversity, we advocate the use of different model architectures.
+By leveraging various networks, we can obtain distinct and complementary features from these models even with the same input data.
+Thus, the proposed DiverseModel differs from the aforementioned methods, exploring different networks in parallel to generate more diversity of pseudo labels to improve training effectiveness.
+Ensemble machine learning is a concept that employs multiple learners and combines their predictions (Sewell, 2008).
+Bagging, a form of ensemble learning, is a technique aimed at reducing prediction variance by creating multiple iterations of a predictor and then utilizing them to form an aggregated predictor (Breiman, 1996).
+Specifically, bagging creates sample subsets by randomly selecting from the training data set and subsequently utilises these acquired subsets to train the foundational models for integration.
+When forecasting a numerical result, the aggregation involves taking an average of the versions, while for predicting a class it relies on a majority vote.
+Bagging is a commonly employed approach for enhancing the robustness and precision of machine learning algorithms for classification and regression (Ren et al., 2016).
+3 DiverseHead (Cross-head Supervision)
 Figure 2: DiverseHead: an online semi-supervised learning approach.
-The slide windows of the freezer randomly select a certain number of heads to freeze the parameter update of heads.
-A mainstream head is randomly selected to be supervised by unsupervised loss.
-This section introduces the details of the proposed DiverseHead Semi-Supervised Learning method.
-Inspired by Bagging (also called bootstrap aggregating), DiverseHead is proposed to create various predictions by multiple heads with even the same structure.
-The source of perturbation is from the dynamic choosing heads to be updated for each iteration, called dynamic freezing, which is equivalent to using different training data for different heads in each iteration.
-The proposed DiverseHead differs from CCT, which does not distinguish the main or auxiliary decoder.
-Also, all decoders are trained by both supervised and unsupervised losses.
-In addition, the proposed method combines self-training and consistency regularization, which calculates unsupervised loss by using optimized pseudo labels.
-Specially, given are both a labelled dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ containing M images and an unlabelled data set $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$ with N images.
-The network $Q$ is constructed with multiple heads denoted as $\left\{head^{i}\right\}_{i=1}^{len(heads)}$.
+This figure applies the dynamic freezing strategy: slide windows of the freezer randomly select a certain number of heads to freeze the parameter of heads (not updated by backpropagation).
+Additionally, during every iteration, all heads undergo supervision through a supervised loss, yet only a single head is randomly chosen to be updated by unsupervised loss.
+Figure 3: The Proposed Voting Module: a voting mechanism for the pseudo label creation.
+In the unsupervised part, the voting module combines the mean output of multiple heads (mean voting) and individual pseudo labels (max voting) to generate more efficient pseudo labels.
+Argmax returns the indices of the maximum values of the prediction along the class dimension.
+The dashed arrow serves as an illustration of a pixel voting for its classification in a segmentation map.
+This section introduces the details of the proposed DiverseHead semi-supervised learning method which is proposed to create various predictions by multiple heads illustrated in Figure 2, specifically the number of convolutional layers for each head is 2.
+The source of perturbation is from two strategies, called dynamic freezing and dropout which are shown in Figure 4 and explained in detail in Section 3.1.
+The proposed DiverseHead does not distinguish between the main or auxiliary decoder, which differs from CCT.
+During every training iteration, all heads undergo supervision through supervised losses, yet only a single head is randomly chosen to be updated by an unsupervised loss.
+In addition, the proposed method combines self-training and consistency regularization, which uses their predictions to supervise itself and forces all perturbed head to give a consistent output.
+It makes an efficient voting module shown in Figure 3 to help calculate unsupervised loss by using optimized pseudo labels.
+To specifically explain the proposed semi-supervised framework, given both a labelled dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ containing $M$ images and an unlabelled data set $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$ with $N$ images, the network $Q$ is constructed with multiple heads denoted as $\left\{head^{i}\right\}_{i=1}^{L}$, where L is the number of heads.
 Each of these heads is initialized differently.
-The expectation is to gain the trained network $Q$ leveraging these labelled and unlabelled data based on the proposed semi-supervised learning methods.
+The expectation of the proposed semi-supervised learning architecture is to gain the trained network $Q$ leveraging these labelled and unlabelled data.
 The output of each head refers to a probability-based prediction for all classifications corresponding to its input.
-When working with labelled data, the supervised loss $\mathcal{L}_{sup}^{j}$ between the $k\leavevmode\nobreak\ th$ ground truth $y_{k}$ and its corresponding prediction $p_{k}^{j}$ for $j\leavevmode\nobreak\ th$ head is defined by using the standard cross-entropy loss function $\ell_{ce}$:
-$\mathcal{L}_{sup}^{j}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k}^{j},y_{k}\right)\right),$ (1)
-where $W$ and $H$ refer to the weight and height of input images. The ultimate supervised loss is determined by the mean of the losses across all heads
-$\mathcal{L}_{sup}=\frac{1}{len(heads)}\sum_{j=1}^{len(heads)}\mathcal{L}_{sup}^{j}.$ (2)
-In each iteration, we also get prediction $r_{k}^{j}$ corresponding to unlabelled data $u_{k}^{j}$ from $j\leavevmode\nobreak\ th$ head.
-In order to get higher precision pseudo labels, inspired by ensemble learning, the predictions from all heads are aggregated to create a combined prediction $\hat{r}_{k}^{sum}$.
-Subsequently, this combined prediction is utilized to calculate pseudo labels $\hat{y}_{k}^{sum}$ through an $argmax$ operation, which returns the indices of the maximum values of the prediction along the class dimension.
-Next, there is a random selector to pick only one head as the main branch to be supervised by minimizing the unsupervised loss
-$\mathcal{L}_{unsup}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(r_{k}^{main},\hat{y}_{k}^{sum}\right)\right),$ (3)
-where $r_{k}^{main}$ is the prediction from the selected head.
+When working with labelled data, the supervised loss $\mathcal{L}_{sup}^{j}$ between the $k^{\prime}th$ ground truth $y_{k}$ and its corresponding prediction $p_{k}^{j}$ for $j^{\prime}th$ head is defined by using the standard cross-entropy loss function $\ell_{ce}$:
+$\mathcal{L}_{sup}^{j}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{j},y_{k}\right),$ (1)
+where $W$ and $H$ refer to the weight and height of input images. The final supervised loss is determined by the mean of the losses across all heads
+$\mathcal{L}_{sup}=\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}$ (2)
+where L is the number of heads.
+In each iteration, we also get prediction $r_{k}^{j}$ corresponding to unlabelled data $u_{k}^{j}$ from the $j^{\prime}th$ head.
+Inspired by ensemble learning, a voting module is proposed to get high-precision pseudo labels.
+There are two voting mechanisms in the proposed voting module, called mean voting and max voting, respectively.
+The former aggregates the predictions from all heads to create a combined prediction $\hat{r}_{k}^{mean}$.
+Subsequently, this combined prediction is used to calculate the mean pseudo labels $\hat{y}_{k}^{mean}$ through an $argmax$ operation, which returns the indices of the maximum values of the prediction along the class dimension.
+Apart from the mean voting module, the individual pseudo label is generated from the output of each head by the $argmax$ function.
+The latter regards all pseudo labels as voters, in which the mean pseudo label contributes $\varphi$ weight and each individual pseudo label contributes unit weight.
+$\varphi$ is a learnable parameter and its value changes depending on the dataset and training processing.
+The max voting module returns the class number that gets the most votes for each pixel.
+After max voting, an optimal pseudo label $\hat{y}_{k}^{final}$ is created to calculate unsupervised loss by using the cross-entropy loss function:
+$\mathcal{L}_{unsup}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(r_{k}^{main},\hat{y}_{k}^{final}\right),$ (3)
+where $r_{k}^{main}$ is the prediction from the selected head. There is a random selector to pick only one head as the main branch to be supervised by minimizing the unsupervised loss.
 Finally, the whole loss is written as
@@
 where $\lambda$ is the trade-off weight between the supervised and unsupervised losses.
-Two perturbation methods.
-Based on the framework of DiverseHead, we propose a parameter perturbation method called dynamic freezing and also apply dropout as another perturbation to increase the diversity of features during training.
-Specifically, the first step of dynamic freezing refers to a process where, during each iteration, half of the heads are randomly selected to be frozen.
-This implies that the parameters within these selected heads will remain unchanged and not undergo updates.
-Every head has an equal probability of being chosen.
-Before going to the next iteration, those frozen heads are unfrozen.
-The algorithm of the DiverseHead semi-supervised learning approach with the perturbation of dynamic freezing is expressed in the Algorithm 1.
-Another form of perturbation involves the utilization of dropout layers.
-We introduce a dropout layer before the last convolutional layer for each head.
-The pseudocode of this method is expressed in Algorithm 2.
-In the training phase, specific components of the feature tensor are randomly assigned a value of zero with a dropout rate (determining the probability), using samples derived from a Bernoulli distribution.
-This dropout operation is employed to enhance the variability of the output predictions.
-In the conducted experiments, the dropout rate is set as a hyperparameter with a value of 0.3.
-Similar to the training procedure of the semi-supervised learning approach based on dynamic freezing, the dropout-based approach adheres to the same training pipeline outlined above, with the distinction that all heads remain unfrozen.
-To evaluate the efficacy of individual perturbation methods in conjunction with the proposed DiverseHead techniques, each method is applied independently within the proposed framework.
-The performance of each combination is assessed in Section V.
+(a) Dynamic Freezing
+(b) Dropout
+Figure 4: The proposed two perturbation methods: dropout and dynamic freezing.
+In (a), the dynamic selector is designed to amplify the parameter diversity of multiple heads.
+It randomly chooses a set number of heads not to be updated by backpropagation during each training iteration, thereby enhancing the variability of parameters.
+3.1 Perturbation Methods
 Algorithm 1 DiverseHead Semi-Supervised Learning with Dynamic Freezing Pseudocode
@@
               Initialise backbone using ResNet-50
+              Initialise model $Q$
               Randomly initialise parameters for each head
@@
                Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
-[…]
-        $\left\{head^{i}\right\}_{i=random}^{len(heads)/2}\leftarrow sample\left(\left\{head^{i}\right\}_{i=1}^{len(heads)}\right)$
-        $Freeze(\left\{head^{i}\right\}_{i=random}^{len(heads)/2})$
-        for $j\in\left\{1,...,len(heads)\right\}$ do
+  $L=length(heads)$
+  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
+        $\mathcal{R}=Randint(0,L,\frac{1}{2}L)$
+        $Freeze(\left\{head^{i}\right\}_{i\in\mathcal{R}})$
+        for $j\in\left\{1,...,L\right\}$ do
             $p_{k}^{j}=Q^{head_{j}}(x_{k}^{j})$
@@
             $r_{k}^{j}=Q^{head_{j}}(u_{k}^{j})$
-        $r_{k}^{sum}=sum\left(\left\{r_{k}\right\}_{k=1}^{len(heads)}\right)$
-        $\hat{y}_{k}^{sum}\leftarrow argmax(r_{k}^{sum})$
-        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{len(heads)}\right\}\right)$
-        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{sum})$ based on (3)
-        $\mathcal{L}\leftarrow\frac{1}{len(heads)}\sum_{j=1}^{len(heads)}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
+            $\hat{y}_{k}^{j}\leftarrow argmax(r_{k}^{j})$
+        $r_{k}^{mean}=sum\left(\left\{r_{k}^{j}\right\}_{j=1}^{L}\right)$
+        $\hat{y}_{k}^{mean}\leftarrow argmax(r_{k}^{mean})$
+        $\hat{y}_{k}^{final}\leftarrow voting(\left\{\hat{y}_{k}^{j}\right\}_{j=1}^{L},\hat{y}_{k}^{mean})$
+        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{L}\right\}\right)$
+        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
+        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
         Update model to minimize $\mathcal{L}$
-        $Unfreeze(\left\{head^{i}\right\}_{i=random}^{len(heads)/2})$
+        $\mbox{\it Unfreeze}(\left\{head^{i}\right\}_{i=R})$
   OUTPUT: Trained model $Q$
+Based on the framework of DiverseHead, we propose a parameter perturbation method called dynamic freezing as shown in Figure 4 (a).
+The pseudocode of the algorithm for DiverseHead with the perturbation of dynamic freezing is given in Algorithm 1.
+Specifically, we use DeepLabv3+ with a ResNet-50 backbone as the segmentation network in our framework, ensuring consistency with the methods compared in this paper.
+The backbone is pretrained on ImageNet, and other parameters are initialised by Kaiming Initialization.
+Both labelled and unlabelled data are input to the framework, then the first step of dynamic freezing refers to a process where half of the heads are randomly selected to be frozen during each iteration.
+This implies that the parameters within these selected heads will remain unchanged and not undergo updates for the current iteration.
+Every head has an equal probability of being chosen.
+Following the cross-head supervision above, the segmentation model is updated by the supervised and unsupervised losses.
+Before going to the next iteration, those frozen heads are unfrozen.
 Algorithm 2 DiverseHead Semi-Supervised Learning with Dropout Pseudocode
@@
                Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
-[…]
-        for $j\in\left\{1,...,len(heads)\right\}$ do
+  $L=length(heads)$
+  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
+        for $j\in\left\{1,...,L\right\}$ do
             $p_{k}^{j}=Q^{head_{j}}(x_{k}^{j})$
@@
             $r_{k}^{j}=Q^{head_{j}}(u_{k}^{j})$
-        $r_{k}^{sum}=sum\left(\left\{r_{k}\right\}_{k=1}^{len(heads)}\right)$
-        $\hat{y}_{k}^{sum}\leftarrow argmax(r_{k}^{sum})$
-        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{len(heads)}\right\}\right)$
-        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{sum})$ based on (3)
-        $\mathcal{L}\leftarrow\frac{1}{len(heads)}\sum_{j=1}^{len(heads)}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
+            $\hat{y}_{k}^{j}\leftarrow argmax(r_{k}^{j})$
+        $r_{k}^{sum}=sum\left(\left\{r_{k}^{j}\right\}_{j=1}^{L}\right)$
+        $\hat{y}_{k}^{mean}\leftarrow argmax(r_{k}^{mean})$
+        $\hat{y}_{k}^{final}\leftarrow voting(\left\{\hat{y}_{k}^{j}\right\}_{j=1}^{L},\hat{y}_{k}^{mean})$
+        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{L}\right\}\right)$
+        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
+        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
         Update model to minimize $\mathcal{L}$
   OUTPUT: Trained Model Q
-III Cross-model Supervision (DiverseModel)
+Another form of perturbation involves the use of dropout layers to the proposed DiverseHead structure to increase the diversity of features during training.
+We introduce a dropout layer after each convolutional block in each segmentation head of the network as shown in Figure 4 (b).
+The pseudocode of this method is given in Algorithm 2.
+As both dynamic freezing and dropout perturbation utilize the same semi-supervised learning framework, DiverseHead, the network initialization and supervised loss calculation procedures stay the same.
+Using dropout in DiverseHead, specific components of the weights in the heads of the network are randomly assigned a value of zero with a dropout rate (determining the probability), using samples derived from a Bernoulli distribution.
+This dropout operation is employed to enhance the variability of the output predictions.
+In the conducted experiments, the dropout rate is set as a hyperparameter with a value of 0.3.
+Although the dropout-based approach adheres to the same training pipeline with dynamic freezing outlined above, the distinction is that all heads remain unfrozen.
+To evaluate the efficacy of individual perturbation methods in conjunction with the proposed DiverseHead techniques, each method is applied independently within the proposed framework.
+The performance of each combination is assessed in Section 6.
+Considering various datasets may require differing levels of diversity in pseudo labels within the proposed DiverseHead framework, adjustments can be made by altering the number of frozen heads and dropout rates in the methods of dynamic freezing and dropout, respectively.
+4 Cross-model Supervision (DiverseModel)
 Figure 5: DiverseModel: an online semi-supervised learning approach.
 The proposed DiverseModel differs from CPS, exploring different networks in parallel to generate comprehensive pseudo labels.
+As shown in Figure 5, the DiverseModel structure includes three distinct networks, which can be various semantic segmentation networks.
+In this paper, we choose three widely used segmentation networks for experiments, which are PSPNet (Zhao et al., 2017), UNet (Ronneberger et al., 2015), SegNet (Badrinarayanan et al., 2017).
+Algorithm 3 presents the pseudocode of the DiverseModel method.
 Since different networks pay different and complementary attention to the same input, this offers the basis that they are able to benefit from each other.
-DiverseModel structure includes three distinct networks which are PSPNet [zhao2017pyramid], UNet [ronneberger2015u], SegNet [badrinarayanan2017segnet], respectively.
-Algorithm 3 presents the pseudocode of the DiverseModel method.
+In order to provide evidence for this claim, we executed the Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2020) technique for every network employed within the framework of the DiverseModel architecture.
+Grad-CAM visualizes the areas of an image that are important to the model predictions from each network.
+Figure 6 depicts an example grad-CAM analysis for the BUILDING class in the Potsdam data set.
+Examining Figure 6, we can see that different networks pay different and complementary attention to the same input, and the pseudo labels from the DiverseModel prediction show the highest quality.
+Figure 6: The upper section displays Grad-CAM outputs from individual networks within the DiverseModel architecture using the Potsdam dataset.
+The lower section showcases both the ground truth and pseudo labels generated through predictions from each network.
+In the lower right corner, the pseudo label from DiverseModel is presented.
 The labelled data is used in a regular supervised learning manner to train these models by using the standard cross-entropy loss function $\ell_{ce}$. The supervised loss is expressed as:
-$\mathcal{L}_{sup}=\sum_{n=1}^{3}\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k}^{n},y_{k}\right)\right),$ (5)
-where $p_{k}^{n}$ represents the prediction of $k\leavevmode\nobreak\ th$ pixel from $n\leavevmode\nobreak\ th$ network.
+$\mathcal{L}_{sup}=\frac{1}{3}\sum_{n=1}^{3}\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{n},y_{k}\right),$ (5)
+where $p_{k}^{n}$ represents the prediction of the $k\leavevmode\nobreak\ th$ pixel from the $n\leavevmode\nobreak\ th$ network.
 In addition, unlabelled data is used to generate pseudo labels, which are then exploited for cross-supervision to inform each network.
-Different from the version presented in [ma2023confidence], all loss calculations in this work solely focus on cross-entropy loss to avoid the effects resulting from different types of loss functions on performance.
-The predictions obtained by each network are denoted as $\{p^{1},p^{2},p^{3}\}$, which are used for generating pseudo labels $\{\hat{r}^{1},\hat{r}^{2},\hat{r}^{3}\}$ through $argmax$ operation.
-For instance, the cross pseudo supervision loss $\mathcal{L}_{unsup}^{12}$ between the prediction $p^{1}$ from first network and the pseudo label $r^{2}$ generated by the second network is defined as:
-$\mathcal{L}_{unsup}^{12}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k}^{1},\hat{r}_{k}^{2}\right)\right).$ (6)
+Different from the version presented in (Ma et al., 2023), all loss calculations in this work solely focus on cross-entropy loss to avoid the variations in performance resulting from different types of loss functions.
+The predictions obtained by each network are denoted as $\{p^{1},p^{2},p^{3}\}$, which are used for generating pseudo labels $\{\hat{r}^{1},\hat{r}^{2},\hat{r}^{3}\}$ through the $argmax$ operation.
+For instance, the cross pseudo supervision loss $\mathcal{L}_{unsup}^{12}$ between the prediction $p^{1}$ from the first network and the pseudo label $r^{2}$ generated by the second network is defined as:
+$\mathcal{L}_{unsup}^{12}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{1},\hat{r}_{k}^{2}\right).$ (6)
 Algorithm 3 DiverseModel Semi-supervised Learning Pseudocode
@@
                Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
-[…]
+  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
         $\mathcal{L}_{sup}=loss(Q^{1}(x_{k}),y_{k})$ + $loss(Q^{2}(x_{k}),y_{k})$
@@
   OUTPUT: Trained Model $Q^{1}$, $Q^{2}$, $Q^{3}$
-The cross-pseudo supervision among three networks creates six losses by using the same way.
-The unsupervised loss $\mathcal{L}_{unsup}$ is the result of linearly adding the six individual losses, as shown below
-$\begin{array}[]{r}\mathcal{L}_{unsup}=\mathcal{L}_{unsup}^{12}+\mathcal{L}_{unsup}^{13}+\mathcal{L}_{unsup}^{21}\\ +\mathcal{L}_{unsup}^{23}+\mathcal{L}_{unsup}^{31}+\mathcal{L}_{unsup}^{32}.\end{array}$ (7)
+The cross-pseudo supervision among three networks creates six losses in the same way. The unsupervised loss $\mathcal{L}_{unsup}$ is the average of the six individual losses, as shown below
+$\begin{array}[]{r}\mathcal{L}_{unsup}=\frac{1}{6}\left(\mathcal{L}_{unsup}^{12}+\mathcal{L}_{unsup}^{13}+\mathcal{L}_{unsup}^{21}+\mathcal{L}_{unsup}^{23}+\mathcal{L}_{unsup}^{31}+\mathcal{L}_{unsup}^{32}\right).\end{array}$ (7)
 The total loss $\mathcal{L}$ is the linear addition of $\mathcal{L}_{sup}$ and $\mathcal{L}_{unsup}$, which is previously given in equation (4)
-IV Dataset Description
-We employed diverse datasets to assess both the proposed techniques and state-of-the-art methods, specifically including (1) the ISPRS Potsdam dataset [rottensteiner2012isprs], (2) the DFC2020 dataset [robinson2021global], (3) the RoadNet dataset [liu2018roadnet], and (4) the Massachusetts Buildings dataset [MnihThesis].
-In the sequel, we share the details of each datasets utilised in this paper.
+5 Dataset Description
+We employed diverse remote sensing datasets to assess both the proposed techniques and state-of-the-art methods, specifically including (1) the ISPRS Potsdam dataset (Rottensteiner et al., 2012), (2) the DFC2020 dataset (Robinson et al., 2021), (3) the RoadNet dataset (Liu et al., 2018), and (4) the Massachusetts Buildings dataset (Mnih, 2013).
+In the sequel, we share the details of each dataset utilised in this paper.
 1) ISPRS Potsdam Semantic Labeling dataset is an open-access benchmark dataset provided by the International Society for Photogrammetry and Remote Sensing (ISPRS).
@@
 This dataset was annotated manually into six land cover classes, which are impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background.
-This dataset provides 38 multi-source patches (all with the size of $6000\times 6000$ pixels), which contain infrared (IR), red, green and blue bands orthorectified optical images with corresponding digital surface models (DSM).
+This dataset provides 38 patches (all of the size $6000\times 6000$ pixels), which contain infrared (IR), red, green and blue bands orthorectified optical images with corresponding digital surface models (DSM).
 For computation purposes, we partitioned all these data tiles into $512\times 512$ patches, resulting in 3456, 201, and 1815 samples for training, validation, and test sets, respectively.
-We randomly select 1/4 of training data as labelled data and use the remaining 3/4 as unlabelled data for semi-supervised learning.
-2) DFC2020 is the 2022 IEEE GRSS Data Fusion Contest dataset which is based on the SEN12MS dataset [schmitt2019sen12ms].
+We randomly select a quarter of the training data as labelled data and use the remaining three quarters as unlabelled data for semi-supervised learning.
+2) DFC2020 is the 2022 IEEE GRSS Data Fusion Contest dataset which is based on the SEN12MS dataset (Schmitt et al., 2019).
 It provides Sentinel-1 SAR imagery, Sentinel-2 multispectral imagery, and corresponding land cover maps across seven areas namely Mexico City, Kippa-Ring, Khabarovsk, Black Forest, Mumbai, Cape Town, Bandar Anzali between 2016 and 2017.
@@
 In this paper, we utilize only one-fifth of the labelled data for training, while the remaining four-fifths are employed as unlabeled data.
-The fine-grained International Geosphere-Biosphere Program (IGBP) classification scheme in SEN12MS was aggregated to ten coarser-grained classes, which are forest (11.3%), shrubland (6.9%), savanna (23.6%), grassland (16.8%), wetlands (1.1%), croplands (17.9%), urban/built-up (10.6%), snow/ice (0.0%), barren (5.2%), and water (6.5%).
-3) RoadNet is a benchmark dataset for road detection with 0.21-m spatial resolution per pixel covering 21 typical urban regions about 8 $km^{2}$ in Ottawa, Canada.
-The annotation data includes road surfaces, road edges, and road centerlines for each image.
+The fine-grained International Geosphere-Biosphere Program (IGBP) classification scheme in SEN12MS was aggregated into ten coarser-grained classes.
+3) RoadNet is a benchmark dataset for road detection with 0.21-m spatial resolution per pixel covering 21 typical urban regionsThe annotation data includes road surfaces, road edges, and road centerlines for each image.
 We only used road surfaces for the segmentation task.
 the number of samples for training, validation and testing is 410, 45 and 387, respectively.
-In the semi-supervised learning approaches, only 1/4 of the annotated training data is used, while the remaining 3/4 of the training data are employed as unlabeled data for training purposes.
+In the semi-supervised learning approaches, only a quarter of the annotated training data is used, while the remaining three-quarters of the training data are employed as unlabeled data for training purposes.
 4) Massachusetts Buildings Dataset predominantly encompasses urban and suburban regions, encompassing structures of varying scales.
@@
 Each subset consists of images sized $1500\times 1500$ pixels.
-One-quarter of the annotated training images and the remaining three-quarters of the unlabelled training data are utilized in semi-supervised learning.
-V experiment and results analysis
-V-A Implementation Details
+A quarter of the annotated training images and the remaining three-quarters of the unlabelled training data are used in semi-supervised learning.
+6 Experimental analysis
+6.1 Implementation Details
 We implemented our methods by using the PyTorch framework.
-Following [chen2018encoder], we used a mini-batch SGD optimizer and adopted a poly learning rate policy, where the current learning rate equals the initial learning rate multiplied by $\left(1-\frac{\text{ iter }}{\text{ max-iter }}\right)^{\text{power }}$.
+Following (Chen et al., 2018), we used a mini-batch SGD optimizer and adopted a poly learning rate policy, where the current learning rate equals the initial learning rate multiplied by $\left(1-\frac{\text{ iter }}{\text{ max-iter }}\right)^{\text{power }}$.
 The initial learning rate and power are set to 0.01 and 0.9, respectively.
-All experiments were conducted on the GW4 Isambard with an NVIDIA A100-sxm GPU and an AMD EPYC 7543P CPU  [gw4].
+All experiments were conducted on the GW4 Isambard with an NVIDIA A100-sxm GPU and an AMD EPYC 7543P CPU  (McIntosh-Smith, 2014).
 We implemented various classic semi-supervised learning methods (MT, CCT, GCT and CPS) on remote sensing datasets.
-All these methods use DeepLabv3+ with ResNet-50 pre-trained on ImageNet as the semantic segmentation network.
-To be fair the proposed DiverseHead is also based on DeepLabv3+ with the pretrained ResNet-50.
+All these methods use DeepLabv3+ with ResNet-50 backbone pre-trained on ImageNet as the semantic segmentation network.
+To be fair the proposed DiverseHead is also based on DeepLabv3+ with the pretrained ResNet-50 on ImageNet.
 We comprehensively analyzed all methods by quantifying performance via class-related measures, including overall accuracy (OA), user’s accuracy (UA), producer’s accuracy (PA), mean intersection over union (mIoU), and $F_{1}$-score.
 Expressions of all five performance metrics are given as follows: $\mathrm{OA}=\frac{TP+TN}{TP+TN+FP+FN}$, $\mathrm{UA}=\frac{TP}{TP+FP}$, $\mathrm{PA}=\frac{TP}{TP+FN}$, $\mathrm{mIoU}=\frac{|\mathrm{TP}|}{|\mathrm{TP}+\mathrm{FN}+\mathrm{FP}|}$, $F_{1}=\frac{2\cdot PA\cdot UA}{PA+UA}$, where TP, TN, FP, and FN refer to the numbers of pixels that are true positives, true negatives, false positives, and false negatives for each class, respectively.
-V-B Quantitative Results and Analysis
-We evaluated the performance of the proposed approach and reference networks using 4 aforementioned datasets, and the average results are presented in Table I.
-The reason we first share average results across datasets is to provide a global applicability display for all utilised techniques, and each data set’s performance will be discussed in the sequel.
-Evaluating the performance results in Table I, we conclude that the proposed DiverseModel achieves the best performance among all metrics.
-In particular, for metrics like UA and mIoU, the proposed DiverseModel exhibits an improvement of over 5% compared to another cross-model-based approach, CPS.
-Although the performance of DiverseHead is lower than that of DiverseModel, it outperforms other reference methods and reaches the second best.
-Upon convergence of contemporary methodologies with remote sensing datasets, MT demonstrates notably inferior average performance across various metrics, with the exception of UA.
-Both CCT and GCT exhibit comparable performance, yet the performance superiority of CCT is notably evident in its PA metric.
-CPS consistently yields superior outcomes across the majority of metrics when juxtaposed against other aforementioned state-of-the-art methodologies.
-Meanwhile, it is worth noting that the DiverseHead consists of a notably lighter architecture compared to all other models where Table II presents the required number of parameters during training for each semi-supervised learning architecture.
+6.2 Quantitative Results and Analysis
+We evaluated the performance of the proposed approaches and some existing semi-supervised frameworks using the five performance metrics above.
+Experiments are conducted on the four previously mentioned datasets, namely Potsdam, DFC2020, RoadNet, and Massachusetts Building.
+The average results are presented in Table 1.
+The reason we first share average results across datasets is to provide a global performance for all utilised techniques, and the detailed performance comparison for all methods on each dataset will be discussed later.
+From the average results in Table 1, the proposed DiverseHead (DF) achieves the best overall performance in which there are 4 of 5 metrics reach the best compared to other methods, which are marked as red.
+DiverseModel shows the second-best performance among all methods in Table 1.
+In particular, for metrics of UA, the proposed DiverseModel exhibits an improvement of over 3.6% compared to another cross-model-based approach, CPS.
+Although the performance of DiverseHead (DT) is marginally lower than that of DiverseHead (DF) and DiverseModel, the three methods exhibit similar performance and surpass other compared methods.
+With these used remote sensing datasets, MT demonstrates notably inferior average performance across various metrics.
+CCT, GCT, and CPS exhibit comparable performance, yet the performance superiority of CPS is notably evident in its PA metric.
+Table 1: Average performance comparison with the state-of-the-art methods on four datasets. DT and DF indicate dropout and dynamic freezing, respectively.
+Models OA UA PA mIoU $F_{1}$
+MT (Tarvainen and Valpola, 2017) 86.33% 74.90% 81.47% 65.97% 77.95%
+CCT (Ouali et al., 2020) 87.10% 76.61% 82.29% 67.71% 79.19%
+GCT (Ke et al., 2020) 87.43% 75.53% 82.04% 67.19% 78.63%
+CPS (Chen et al., 2021) 88.03% 75.90% 85.17% 68.10% 80.13%
+DiverseModel (Ma et al., 2023) 88.77% 79.54% 85.18% 70.92% 82.02%
+DiverseHead (DT) 88.51% 78.40% 84.84% 70.07% 81.29%
+DiverseHead (DF) 88.97% 78.99% 85.82% 71.14% 82.08%
+It is important to highlight that the DiverseHead is a significantly lightweight semi-supervised learning architecture aimed at segmentation tasks within remote sensing imagery.
+Table 2 presents the required number of parameters during training for each semi-supervised learning architecture.
+Apart from DiverseHead, all other semi-supervised learning approaches typically involve parameters exceeding three hundred megabytes during training.
+Aside from DiverseModel, all other methods are based on the same segmentation network namely DeepLabv3+.
+DiverseModel is the largest architecture although it provides competitive performance compared to other cross-model-based methods namely MT, CCT, GCT, and CPS as shown in Table 1.
 It is evident that the architectures using multiple segmentation networks during training typically need higher memory demands.
-In contrast, the proposed DiverseHead is remarkably lightweight as it only employs a single model with multiple heads which consists of only 2 convolution layers, eliminating the need for multiple models during training.
-Thus, the parameter size of the DiverseHead (DT&DF) is only 16% bigger than that of the single network (Base in Table II), whereas the parameter size of other reference semi-supervised architectures is at least twice that of the single network.
-Despite this need for a lower number of parameters, the DiverseHead proposed method outperforms all state-of-the-art with at least 1% accuracy whilst obtaining similar performance to the companion method of DiverseModel.
-Table I: Average performance comparison with the state of the arts on 4 datasets. DT and DF present Dropout and dynamic freezing, respectively.
-OA UA PA mIoU $F_{1}$
-MT[tarvainen2017mean] 85.63% 73.14% 80.10% 63.93% 76.39%
-CCT[ouali2020semi] 86.60% 72.17% 83.29% 64.07% 77.22%
-GCT[ke2020guided] 86.53% 73.33% 80.50% 64.59% 76.72%
-CPS[chen2021semi] 87.17% 72.89% 84.62% 65.02% 78.18%
-DiverseHead (DT) 88.42% 77.80% 84.95% 69.70% 81.05%
-DiverseHead (DF) 88.29% 76.98% 84.87% 68.85% 80.58%
-DiverseModel 88.86% 78.24% 86.06% 70.53% 81.79%
-Table II: The required parameter size in each semi-supervised learning approach. Base means supervised segmentation network DeepLabv3+. DT and DF present Dropout and dynamic freezing, respectively.
-CPS [chen2021semi] MT [tarvainen2017mean] CCT [ouali2020semi] GCT [ke2020guided]
+In contrast, the proposed DiverseHead is remarkably lightweight as it only employs a single model with multiple heads which consists of only a few small convolutional layers, eliminating the need for multiple models during training.
+Thus, the parameter size of the DiverseHead (DT&DF) is only 16% bigger than that of the single network (Base in Table 2), whereas the parameter size of other reference semi-supervised architectures is at least twice that of the single network.
+Although just a a few parameters are required, the DiverseHead proposed method outperforms this state-of-the-art with at least 1% accuracy and 3% mIoU whilst obtaining similar performance to the companion method of DiverseModel.
+Table 2: The required parameter size in each semi-supervised learning approach. Base means supervised segmentation network DeepLabv3+. DT and DF present dropout and dynamic freezing, respectively.
+CPS (Chen et al., 2021) MT (Tarvainen and Valpola, 2017) CCT (Ouali et al., 2020) GCT (Ke et al., 2020)
 Size (MB) 303.378 303.378 337.655 335.049
-Base DiverseHead(DT) DiverseHead(DF) DiverseModel
-Size (MB) 151.689 175.806 175.806 936.262
-Specifically, we present the results of each semi-supervised learning method for the 4 utilised datasets in Table III.
-DiverseNet architectures reach a remarkable performance for each of the datasets.
-Specifically, the proposed DiverseModel achieves the best performance on both RoadNet and Massachusetts building datasets for all metrics, and the DiverseHead (DT) shows the second best for the most of metrics on these two datasets.
-For the other two datasets, Potsdam and DFC2020, The DiverseHead (DF) method shows competitive and better performance compared to the companion DiverseModel for most of the metrics.
-Table III: Performance comparison with the state of the arts on 4 datasets. DT and DF present Dropout and dynamic freezing, respectively.
-OA UA PA mIoU $F_{1}$
-Potsdam MT[tarvainen2017mean] 81.98% 73.66% 78.39% 63.07% 75.95%
-CCT[ouali2020semi] 82.66% 74.64% 77.62% 64.16% 76.10%
-GCT[ke2020guided] 83.99% 75.65% 80.81% 65.80% 78.14%
-CPS[chen2021semi] 85.00% 75.76% 82.94% 66.69% 79.19%
-DiverseHead (DT) 84.60% 77.42% 80.49% 67.18% 78.92%
-DiverseHead (DF) 85.37% 78.43% 81.86% 68.49% 80.11%
-DiverseModel 85.76% 76.75% 83.45% 67.85% 79.96%
-DFC2020 MT[tarvainen2017mean] 78.64% 59.59% 73.57% 50.44% 65.85%
-CCT[ouali2020semi] 79.71% 59.87% 76.40% 51.04% 67.13%
-GCT[ke2020guided] 80.84% 61.47% 71.43% 52.17% 66.07%
-CPS[chen2021semi] 81.50% 61.75% 79.47% 53.21% 69.50%
-DiverseHead (DT) 82.11% 62.31% 80.38% 53.93% 70.20%
-DiverseHead (DF) 82.19% 62.29% 80.73% 54.02% 70.32%
-DiverseModel 81.87% 62.20% 80.69% 53.71% 70.25%
-RoadNet MT[tarvainen2017mean] 93.00% 80.74% 85.54% 72.96% 83.07%
-CCT[ouali2020semi] 93.90% 83.85% 87.26% 76.23% 85.52%
-GCT[ke2020guided] 93.57% 79.26% 89.25% 73.51% 83.96%
-CPS[chen2021semi] 94.36% 84.21% 88.85% 77.44% 86.47%
-DiverseHead (DT) 96.28% 90.89% 91.95% 84.84% 91.42%
-DiverseHead (DF) 95.36% 87.99% 90.31% 81.35% 89.13%
-DiverseModel 96.79% 92.82% 92.57% 86.87% 92.70%
-Massachusetts MT[tarvainen2017mean] 88.92% 78.56% 82.90% 69.25% 80.67%
-CCT[ouali2020semi] 89.65% 79.91% 84.23% 70.95% 82.01%
-GCT[ke2020guided] 87.73% 76.95% 80.51% 66.88% 78.69%
-CPS[chen2021semi] 87.85% 69.85% 87.24% 62.76% 77.58%
-DiverseHead (DT) 90.69% 80.56% 86.97% 72.86% 83.64%
-DiverseHead (DF) 90.25% 79.22% 86.59% 71.53% 82.74%
-DiverseModel 91.01% 81.19% 87.53% 73.67% 84.24%
-We also considered a downgraded version of the current DiverseHead approach, called single-head supervision (SHS).
-The segmentation network consists of only a single head.
-During each iteration, the current model’s predictions for unlabeled data are utilized to generate pseudo labels through an $\arg\max$ operation.
+DiverseModel (Ma et al., 2023) DiverseHead(DT) DiverseHead(DF) Base
+Size (MB) 936.262 175.806 175.806 151.689
+Specifically, we present the results of each semi-supervised learning method for the 4 utilised datasets in Table 3.
+The performance of the DiverseNet family, namely DiverseHead and DiverseModel, is noticeably superior to that of the other listed semi-supervised learning methods across all evaluated datasets, as indicated by the results.
+In particular, for the Potsdam and Massachusetts Building datasets, the proposed DiverseHead attains the highest performance across 4 out of 5 segmentation metrics whilst DiverseModel emerges as the second-best method based on its results.
+However, in the case of the RoadNet and DFC2020 datasets, DiverseModel reaches the best for most performance metrics, securing the top position whilst DiverseHead closely follows as the second-best method, despite their similar performance.
+The performance of DiverseHead (DT) consistently falls between that of DiverseModel and DiverseHead (DF) for most used datasets apart from DFC2020.
+This dataset provides very low-resolution labels, distinguishing it from other datasets.
+This could be one of the reasons why DiverseHead (DT) outperforms DiverseHead (DF) in DFC2020.
+Table 3: Performance comparison with the state of the arts on four datasets. DT and DF present Dropout and dynamic freezing, respectively.
+Model OA UA PA mIoU $F_{1}$
+Potsdam MT(Tarvainen and Valpola, 2017) 81.98% 73.66% 78.39% 63.07% 75.95%
+CCT(Ouali et al., 2020) 82.66% 74.64% 77.62% 64.16% 76.10%
+GCT(Ke et al., 2020) 83.99% 75.65% 80.81% 65.80% 78.14%
+CPS(Chen et al., 2021) 85.00% 75.76% 82.94% 66.69% 79.19%
+DiverseModel(Ma et al., 2023) 85.76% 76.75% 83.45% 67.85% 79.96%
+DiverseHead (DT) 84.66% 77.04% 80.78% 67.12% 78.87%
+DiverseHead (DF) 85.98% 79.15% 82.87% 69.63% 80.97%
+DFC2020 MT(Tarvainen and Valpola, 2017) 78.64% 59.59% 73.57% 50.44% 65.85%
+CCT(Ouali et al., 2020) 79.71% 59.87% 76.40% 51.04% 67.13%
+GCT(Ke et al., 2020) 80.84% 61.47% 71.43% 52.17% 66.07%
+CPS(Chen et al., 2021) 81.49% 61.74% 79.46% 53.20% 69.49%
+DiverseModel(Ma et al., 2023) 81.87% 62.20% 80.69% 53.71% 70.25%
+DiverseHead (DT) 82.02% 62.13% 80.61% 53.81% 70.18%
+DiverseHead (DF) 81.78% 61.81% 80.21% 53.46% 69.82%
+RoadNet MT(Tarvainen and Valpola, 2017) 94.53% 86.77% 87.94% 78.81% 87.35%
+CCT(Ouali et al., 2020) 95.45% 89.15% 89.91% 81.97% 89.53%
+GCT(Ke et al., 2020) 95.20% 86.27% 90.93% 80.36% 88.54%
+CPS(Chen et al., 2021) 95.66% 88.84% 90.95% 82.46% 89.88%
+DiverseModel(Ma et al., 2023) 96.84% 93.12% 92.58% 87.11% 92.84%
+DiverseHead (DT) 96.16% 91.70% 90.98% 84.70% 91.33%
+DiverseHead (DF) 96.68% 91.99% 92.73% 86.33% 92.36%
+Massachusetts MT(Tarvainen and Valpola, 2017) 90.16% 79.58% 85.97% 71.55% 82.65%
+CCT(Ouali et al., 2020) 90.59% 82.77% 85.23% 73.67% 83.98%
+GCT(Ke et al., 2020) 89.67% 78.75% 85.01% 70.42% 81.76%
+CPS(Chen et al., 2021) 89.95% 77.26% 87.33% 70.07% 81.98%
+DiverseModel(Ma et al., 2023) 90.62% 86.07% 84.01% 75.00% 85.03%
+DiverseHead (DT) 91.21% 82.73% 86.98% 74.67% 84.80%
+DiverseHead (DF) 91.43% 83.00% 87.45% 75.16% 85.17%
+In order to prove the efficiency of the multi-head framework, we considered a downgraded version of the current DiverseHead approach, called single-head supervision (SHS).
+The segmentation framework consists of only a single head.
+During each training iteration, the current model’s predictions for unlabeled data are used to generate pseudo labels through an $argmax$ operation.
 These generated pseudo labels are then used to calculate the unsupervised loss.
-The baseline is the segmentation network only trained by labelled data.
-The performance of the baseline and SHS, along with the proposed DiverseHead, are presented in Table IV in order to provide experimental evidence of the novelty of utilisation of multiple heads in DiverseHead architecture.
-While the single-head supervision benefits from the pseudo labels and exhibits better performance compared to the baseline model especially improving the metric of PA, both versions of DiverseHead approaches demonstrate a further significant improvement by taking advantage of multiple heads.
-Specifically, the mIoU of DiverseHead (DF) is 5.7% higher than that of Baseline and approximately 4% higher than that of SHS for Potsdam.
-Also, the mIoU value of DiverseHead (DT) is about 10% higher than that of SHS.
-Table IV: performance comparison with single-head supervision (SHS) and baseline model which is only supervised by labelled data. DT and DF present Dropout and dynamic freezing, respectively.
-OA UA PA mIoU $F_{1}$
-Potsdam Baseline 81.64% 73.86% 76.05% 62.72% 74.94%
+Also, we provide the performance of the segmentation network namely DeepLabv3+, called Base in Table 4, only trained by labelled data (part of each dataset) without using unlabelled data.
+The performance of the Base and SHS, along with the proposed DiverseHead, are presented in Table 4 to provide experimental evidence of the improvement and efficiency of the use of multiple heads in the DiverseHead architecture.
+While the single-head supervision benefits from the pseudo labels and exhibits better performance compared to the Base model, especially improving the metric of PA, both versions of DiverseHead approaches demonstrate a further significant improvement by taking advantage of multiple heads.
+Specifically, the mIoU of DiverseHead (DF) is 6.91% higher than that of Base and 5.12% higher than that of SHS for Potsdam.
+While the performance of DiverseHead (DT) is slightly lower than that of DiverseHead (DF), its performance is still much better than that of SHS and Base.
+Similarly, DiverseHead demonstrates superior performance on the RoadNet dataset compared to SHS and Base, particularly evident in the mIoU metric, where it exhibits a significant improvement of 6.22% over Base.
+Table 4: Performance comparison with single-head supervision (SHS) and baseline model which is only supervised by labelled data. DT and DF present dropout and dynamic freezing, respectively.
+Model OA UA PA mIoU $F_{1}$
+Potsdam Base 81.64% 73.86% 76.05% 62.72% 74.94%
 SHS 83.43% 74.57% 79.93% 64.51% 77.16%
-DiverseHead (DT) 84.60% 77.42% 80.49% 67.18% 78.92%
-DiverseHead (DF) 85.37% 78.43% 81.86% 68.49% 80.11%
-RoadNet Baseline 93.56% 82.85% 86.53% 75.06% 84.65%
-SHS 93.79% 81.19% 88.62% 74.87% 84.74%
-DiverseHead (DT) 96.28% 90.89% 91.95% 84.84% 91.42%
-DiverseHead (DF) 95.36% 87.99% 90.31% 81.35% 89.13%
-Table V: Ablation study about the number of heads for DiverseHead with dynamic freezing
+DiverseHead (DT) 84.66% 77.04% 80.78% 67.12% 78.87%
+DiverseHead (DF) 85.98% 79.15% 82.87% 69.63% 80.97%
+RoadNet Base 94.86% 88.22% 88.31% 80.11% 88.26%
+SHS 95.32% 87.08% 90.82% 80.95% 88.91%
+DiverseHead (DT) 96.16% 91.70% 90.98% 84.70% 91.33%
+DiverseHead (DF) 96.68% 91.99% 92.73% 86.33% 92.36%
+Table 5: Ablation study on the number of heads for DiverseHead with dynamic freezing
                  # of heads OA UA PA mIoU $F_{1}$
-Potsdam 8 85.70% 78.47% 82.09% 68.82% 80.24%
-10 85.37% 78.43% 81.86% 68.49% 80.11%
-12 85.51% 77.23% 83.03% 68.14% 80.02%
-RoadNet 8 94.28% 80.20% 92.23% 75.59% 85.79%
-10 95.36% 87.99% 90.31% 81.35% 89.13%
-12 95.35% 85.57% 92.22% 80.52% 88.77%
-Average 8 89.99% 79.33% 87.16% 72.21% 83.02%
-10 90.36% 83.21% 86.08% 74.92% 84.62%
-12 90.43% 81.40% 87.62% 74.33% 84.40%
+Potsdam                       8 85.92% 77.86% 83.86% 68.91% 80.75%
+                      10 85.98% 79.15% 82.87% 69.63% 80.97%
+                      12 85.70% 78.54% 82.55% 68.96% 80.50%
+RoadNet                       8 96.51% 91.36% 92.55% 85.67% 91.95%
+                      10 96.68% 91.99% 92.73% 86.33% 92.36%
+                      12 96.66% 92.12% 92.58% 86.31% 92.35%
+Average                       8 91.22% 84.61% 88.20% 77.29% 86.35%
+                      10 91.33% 85.57% 87.80% 77.98% 86.66%
+                      12 91.18% 85.33% 87.57% 77.64% 86.42%
 An ablation study was conducted on the Potsdam and RoadNet datasets to explore the impact of the number of heads for the proposed DiverseHead method with dynamic freezing perturbation.
-The results in Table V indicate that using 10 heads yields the optimal choice across most metrics in terms of average performance.
-The performance difference between the versions with different numbers of heads is very small on Potsdam, but the version of 8 heads shows a performance drop.
-Considering the stability, We chose the 10-head version of DiverseHead (DF) for all experiments in this paper.
-Table VI: Performance comparison of the DiverseModel with its constituent networks
-OA UA PA mIoU $F_{1}$
+The results in Table 5 indicate that using 10 heads yields the optimal choice across most metrics in terms of average performance though the performance difference between the versions with different numbers of heads is small on both datasets.
+Thus, we chose the 10-head version of DiverseHead for all experiments in this paper.
+Table 6: Performance comparison of the DiverseModel with its constituent networks on Potsdam dataset
+Model OA UA PA mIoU $F_{1}$
 UNet 83.28% 74.18% 78.88% 64.51% 76.46%
@@
 DiverseModel 85.76% 76.75% 83.45% 67.85% 79.96%
-Since the DiverseModel technique uses three different segmentation networks, we also evaluated the performance of each individual member network within the DiverseModel in Table VI.
-The results indicate that the cross-supervision of all networks shows a significant improvement in the segmentation task performance compared to each separate network.
-Figure 4: Visual results of each method on Potsdam Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
-Figure 5: Visual results of each method on RoadNet Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
-Figure 6: Visual results of each method on DFC2020 Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
-Figure 7: Visual results of each method on Massachusetts Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
-V-C Qualitative Results and Analysis
-The visible results are presented in Figures from 4 to 7.
-For each dataset, we chose one case and depicted the RGB image, ground truth, and predictions of all the methods.
-Compared to the ground truth, the proposed three methods (right-top 3 sub-figures) achieve better similarity to the ground truth compared to the reference models.
-We also calculated the IoU scores for all predictions with the ground truth for this case, which are shown in the sub-caption of each prediction.
-The IoU scores also demonstrate that the performance of the proposed DiverseNet methods is more competitive than that of the other approaches.
-VI Conclusion
-In this paper, we proposed a lightweight semi-supervised learning based on a multi-head structure called DiverseHead, which equipped a dynamic freezing and dropout perturbation approaches.
-The multi-head idea benefits from bagging and shows competitive performance for the segmentation of remote sensing imagery.
-Also, another high-efficiency multiple network based semi-supervised learning method called DiverseModel is investigated for further analysis.
-DiverModel shows the strongest performance on various datasets compared to the state of the arts but requires more computational resources, whilst the performance of DiverseHead is slightly lower than DiverseModel but dramatically lighter than all reference models.
+Since the DiverseModel technique uses three different segmentation networks, we also evaluated the performance of each member network within the DiverseModel on the Potsdam dataset in Table 6.
+The component models UNet, SegNet, and PSPNet use unlabeled data through the way of SHS.
+DiverseModel demonstrates a significant improvement across all performance metrics in the segmentation task compared to each network.
+In particular, the metric of PA experiences a notable improvement of 4.57% compared to the best-performing individual component model.
+This phenomenon can be attributed to the enhancement of pseudo-label diversity through cross-model supervision, resulting in a significant improvement in PA (recall) during the test phase.
+The findings suggest that the cross-supervision of different networks has the potential to achieve superior performance compared to the best-performing individual component.
+Potsdam DFC2020 RoadNet Massachusetts
+Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Ground Truth [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+MT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+IoU = 39.38% IoU = 70.08% IoU = 79.30% IoU = 73.37%
+CCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+IoU = 51.19% IoU = 78.21% IoU = 86.24% IoU = 75.97%
+GCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+IoU = 40.35% IoU = 76.66% IoU = 80.96 IoU = 74.20%
+Potsdam DFC2020 RoadNet Massachusetts
+CPS Refer to caption Refer to caption Refer to caption Refer to caption
+IoU = 50.21% IoU = 79.39% IoU = 84.92% IoU = 74.16%
+DiverseModel Refer to caption Refer to caption Refer to caption Refer to caption
+IoU = 58.76% IoU = 80.28% IoU = 94.67% IoU = 76.84%
+DiverseHead (Dropout) Refer to caption Refer to caption Refer to caption Refer to caption
+IoU = 57.46% IoU = 80.12% IoU = 93.01% IoU = 78.93%
+DiverseHead (Dynamic Freezing) Refer to caption Refer to caption Refer to caption Refer to caption
+IoU = 58.25% IoU = 80.31% IoU = 90.35% IoU = 78.17%
+Figure 7: Visual results of each method on four datasets. The IoU score of prediction for each method is shown in its sub-caption.
+6.3 Qualitative Results and Analysis
+The visible results are presented in Figures 7.
+For each dataset, we randomly selected one case of the RGB image, its ground truth, and predictions for all the methods.
+Also, the IoU scores are calculated for all predictions with the ground truth for each case, which are shown in the sub-caption of each prediction.
+Based on the IoU values of visual predictions, the highest score is obtained by either DiverseModel or DiverseHead (DT&DF).
+Especially for the RoadNet dataset, both the DiverseModel and DiverseHead families achieve an IoU of over 90% surpassing other methods by at least 3.5%.
+Visually, the segmentation maps of DiverseModel and DiverseHead display better continuity and show better similarity to the ground truth than other methods.
+7 Conclusion
+In this paper, we proposed a lightweight semi-supervised learning approach based on a multi-head structure called DiverseHead.
+Based on the multi-head structure we provide two perturbation methods, namely dynamic freezing and dropout.
+Taking inspiration from the theory of bagging, a voting mechanism is proposed to generate beneficial pseudo-labels in the training stage.
+This simple and lightweight semi-supervised learning framework shows competitive performance for the segmentation of remote sensing imagery.
+Furthermore, we conducted further analysis and additional evaluation on the previously proposed multi-network-based semi-supervised learning method known as DiverseModel.
+Based on the results obtained from the aforementioned four remote sensing datasets, DiverseHead and DiverseModel demonstrate comparable performance while significantly outperforming various widely used semi-supervised learning frameworks.
+It is important to highlight that the training structure of DiverseHead is significantly lighter than that of DiverseModel and other state-of-the-art methods, while still achieving competitive performance.
+This implies that DiverseHead is a better option for those who lack high-memory computation resources.

### v3

@@
 Abstract
-Semi-supervised learning aims to help reduce the cost of the manual labelling process by leveraging valuable features extracted from a substantial pool of unlabeled data alongside a limited set of labelled data during the training phase.
-Since pixel-level manual labelling in large-scale remote sensing imagery is expensive, semi-supervised learning becomes an appropriate solution to this.
-However, most of the existing consistency learning frameworks based on network perturbation are very bulky.
-There is still a lack of lightweight and efficient perturbation methods to promote the diversity of features and the precision of pseudo labels during training.
-In order to fill this gap, we propose DiverseNet which explores multi-head and multi-model semi-supervised learning algorithms by simultaneously enhancing precision and diversity during training.
-The two proposed methods in the DiverseNet family, namely DiverseHead and DiverseModel, both achieve the better semantic segmentation performance in four widely utilised remote sensing imagery data sets compared to state-of-the-art semi-supervised learning methods.
-Meanwhile, the proposed DiverseHead architecture is simple and relatively lightweight in terms of parameter space compared to the state-of-the-art methods whilst reaching high-performance results for all the tested data sets.
-keywords:
-Semi-supervised Learning, Consistency Learning, Semantic Segmentation
+Semi-supervised learning (SSL) aims to help reduce the cost of the manual labelling process by leveraging a substantial pool of unlabelled data alongside a limited set of labelled data during the training phase.
+Since pixel-level manual labelling in large-scale remote sensing imagery is expensive and time-consuming, semi-supervised learning has become a widely used solution to deal with this.
+However, the majority of existing SSL frameworks, especially various teacher-student frameworks, are too bulky to run efficiently on a GPU with limited memory.
+There is still a lack of lightweight SSL frameworks and efficient perturbation methods to promote the diversity of training samples and enhance the precision of pseudo labels during training.
+In order to fill this gap, we proposed a simple, lightweight, and efficient SSL architecture named DiverseHead, which promotes the utilisation of multiple decision heads instead of multiple whole networks.
+Another limitation of most existing SSL frameworks is the insufficient diversity of pseudo labels, as they rely on the same network architecture and fail to explore different structures for generating pseudo labels.
+To solve this issue, we propose DiverseModel to explore and analyse different networks in parallel for SSL to increase the diversity of pseudo labels.
+The two proposed methods, namely DiverseHead and DiverseModel, both achieve competitive semantic segmentation performance in four widely used remote sensing imagery datasets compared to state-of-the-art semi-supervised learning methods.
+Meanwhile, the proposed lightweight DiverseHead architecture can be easily applied to various state-of-the-art SSL methods while further improving their performance.
+The code is available at https://github.com/WANLIMA-CARDIFF/DiverseNet.
+Index Terms:
+Semi-supervised Learning, Semantic Segmentation, Land over classification, Building Detection, Roadnet Detection
+I Introduction
+Supervised deep learning has become the dominant technique in computer vision during the last decade.
+Building on its success in computer vision, many remote sensing applications, such as land cover classification, change detection and object detection have seen significant improvements as similar tasks [1, 2, 3, 4, 5].
+Nevertheless, supervised learning necessitates a substantial and meticulously labelled dataset.
+In the case of extensive remote sensing data, such as satellite imagery and drone-captured images in complex terrains, acquiring pixel-wise expert annotations is a time-consuming, labour-intensive, and costly process.
+While the field of computer vision provides numerous well-annotated datasets, transferring deep learning models trained on these datasets to the remote sensing domain is a formidable challenge.
+This is mainly due to the substantial differences between typical computer vision images and remote sensing data, such as hyperspectral and synthetic aperture radar (SAR) imagery, which often exhibit unconventional and non-intuitive characteristics.
+To address the issue of limited access to massive labelled datasets, SSL offers a viable solution by using a small amount of labelled data while leveraging the abundance of unlabelled data [6, 7, 8].
+This is because large volumes of unlabelled imagery generally can be easily and freely accessed from open-access remote sensing data sources.
 […]
-organization=School of Computer Science and Informatics, Cardiff University,city=Cardiff, postcode=CF24 4HJ, country=United Kingdom
-1 Introduction
-Supervised deep learning has gradually become the dominant technique in computer vision, especially during the last decade.
-Specifically, with its great success in semantic segmentation in computer vision, automated land cover classification approaches have been significantly improved  (Vali et al., 2020).
-Nevertheless, supervised learning necessitates a substantial and meticulously labelled dataset.
-In the case of extensive remote sensing data, such as satellite imagery and drone-captured images in complex terrains, acquiring pixel-wise expert annotations is a time-consuming, labour-intensive, and costly endeavour.
-While the field of computer vision provides numerous well-annotated datasets, transferring deep learning models trained on these datasets to the remote sensing domain is a formidable challenge.
-This is primarily due to the significant differences between images generally used in computer vision and remote sensing data, including hyperspectral and synthetic aperture radar (SAR) imagery, in which images often possess unconventional and non-intuitive characteristics.
-Fortunately, thanks to open-access remote sensing data sources, a large amount of unlabelled imagery can be accessed easily and freely.
-Thus, only using a limited amount of labelled data and exploiting the benefits of a large amount of unlabelled data could be a feasible solution for solving the problem of the lack of labelled remote sensing datasets (Wang et al., 2022a; Yang et al., 2023; Wang et al., 2023).
-In computer vision, semi-supervised learning has become a popular research direction to eliminate the labour-intensive and expensive annotation stages mentioned above.
-It leads to competitive performance compared to supervised learning in applications such as image classification (Zhang et al., 2021; Huang et al., 2021; Wang et al., 2021b) and segmentation (Hu et al., 2021; Wang et al., 2022b, 2021a), which exploit just a part of labelled data and a large amount of unlabelled data.
-Specifically, mainstream semi-supervised learning methods exploit “pseudo” labels that are fake and/or auto-generated label information obtained as a result of an unsupervised step.
-Since pseudo labels are involved in the model training stage, the accuracy of pseudo labels is regarded as a key factor affecting the performance of semi-supervised learning approaches.
-However, the accuracy of pseudo labels is not always satisfactory especially when labelled training data is not complete and sufficient enough.
-Meanwhile, increasing diversity of the pseudo-labels has become another key research direction within the semi-supervised learning domain, especially for consistency regularization (French et al., 2019).
-This is because a greater diversity of pseudo-labels enhances the effectiveness of training deep learning networks.
-In order to empower networks to harness the potentials within unlabeled data, a notable technique known as consistency regularization has emerged as a widely adopted method for semi-supervised learning.
-(French et al., 2019; Filipiak et al., 2021).
-Specifically, consistency regularization methods are built up on the theory of assumption of smoothness, which posits that if two points reside in a high-density region of feature space and are close to each other, their corresponding labels should be the same or consistent (Chen and Wang, 2010; Luo et al., 2018; French et al., 2019).
-In practice, these methods for consistency regularization ensure adherence to this assumption by compelling networks to generate predictions that remain consistent across modified iterations of unlabeled inputs.
+SSL has become a widely used technique in computer vision to reduce the labour-intensive and costly annotation process for various applications such as image classification [9, 10, 11] and segmentation [12, 13, 14], whilst leading to competitive performance.
+Specifically, taking advantage of “pseudo” labels generated by the prediction of unlabelled data has become a mainstream class of SSL methods.
+Thus, the quality of pseudo labels becomes a crucial factor in determining the training effectiveness of the models.
+With the success of SSL in computer vision, many SSL semantic segmentation approaches for remote sensing imagery have been explored, especially during the last decade [15, 16, 17].
+Specifically, based on consistency learning, U-MCL [18] proposes an uncertainty-aware technique integrating masks for SSL semantic segmentation of remote sensing imagery, and [19] promotes the importance of bias-correction .
+In the SSL literature, enhancing the accuracy and diversity of pseudo labels has become a major challenge and a central research focus.
+This is due to the fact that pseudo label accuracy is often inadequate, especially when the labelled training data is limited or incomplete.
+On the other hand, enhancing the diversity of pseudo labels becomes another key research focus in SSL to improve model robustness, particularly in the context of consistency regularisation.
+[20].
+In order to improve the quality of pseudo labels and empower networks to harness the potential of unlabelled data, a notable technique known as consistency regularisation has emerged as a widely adopted method for SSL.
+[20, 21].
+Specifically, consistency regularisation methods are built up on the theory of assumption of smoothness, which suggests that if two points lie in a high-density region of feature space and are close to each other, their corresponding labels should be the same or consistent [22, 23, 20].
+In practice, consistency regularisation SSL executes this assumption by forcing networks to produce consistent predictions for modified versions of unlabelled data or their features using various perturbation techniques.
 The strategies for the aforementioned perturbations can be categorised into three groups, namely input, feature, and network perturbations.
-Input perturbation is a method that involves making adjustments or alterations to the input data as part of the training process.
-In this way, the semi-supervised learning approach promotes the consistency of predictions made in response to these modified inputs.
-Widely used input perturbation methods employed in semi-supervised learning include Cutout (DeVries and Taylor, 2017), MixUp (Zhang et al., 2017), and CutMix (Yun et al., 2019).
-Among them, CutMix stands out by combining aspects of MixUp and CutOut techniques, demonstrating superior performance compared to them (French et al., 2019).
-However, input perturbation introduces artificial noise to the data, which might lead to incorrect or noisy labels when generating pseudo-labels for unlabeled examples.
-This might negatively impact the training process by providing incorrect guidance to the model when using pseudo labels.
-Cross-consistency training (CCT) (Ouali et al., 2020) presents a feature perturbation method that introduces noise to both low- and high-level features.
+Input perturbation involves modifying or altering input data, enabling SSL approaches to enforce consistency in predictions for these altered inputs.
+The widely used input perturbation method is adding artificial noise to input images.
+However, it might lead to incorrect or noisy pseudo labels for unlabelled examples and negatively impact the training efficiency by providing incorrect guidance to the model when training with these pseudo labels.
+Apart from input perturbation, feature perturbation methods introduce noise to both low- and high-level features.
 These perturbed features are then fed into multiple decoders to generate multiple outputs, followed by the enforcement of consistency among the outputs obtained from different decoders.
-However, feature perturbation alters the intrinsic characteristics and relationships between features, which also might lead to the creation of artificial feature representations that do not accurately reflect the underlying patterns in the original data.
-Consequently, the perturbed features could result in incorrect or noisy labels when generating pseudo-labels for unlabeled examples, which negatively impacts the learning process.
-Lastly, network perturbation uses different networks or the same network with different initialization weights to promote diversity of predictions, such as cross pseudo supervision (CPS) (Chen et al., 2021), confidence-guided semi-supervised learning (CGSSL) (Ma et al., 2023), mean teacher (MT), (Tarvainen and Valpola, 2017), CCT (Ouali et al., 2020), guided collaborative training (GCT) (Ke et al., 2020) and cross teaching between CNN and transformer (Luo et al., 2022).
-Compared to the previous two perturbation methods, although network perturbation methods introduce perturbations created by the model itself instead of artificial noise, they generally require rapidly increased computational resources due to the increased number of networks (whole or their inner stages) that are used.
-(a) DiverseHead
-(b) DiverseModel
-[…]
-1.1 Contributions
-The preceding overview underscores that, although previous perturbation techniques in semi-supervised learning have their advantages, they also bring inherent drawbacks which could be a key consideration affecting the performance in semi-supervised learning.
-In addressing the challenges posed by perturbation-based semi-supervised learning and its complexity, it becomes imperative to explore more sophisticated and efficient approaches.
-These approaches aim to enhance the capabilities of semi-supervised learning frameworks by integrating advanced perturbation methods.
-This paper proposes two perturbation-based semi-supervised network architectures, coined as DiverseNet, which consist of decision (DiverseHead) and novel feature (DiverseModel) diversifying network architectures for the purposes of semi-supervised semantic segmentation of remote sensing imagery.
-The proposed lightweight semi-supervised semantic segmentation network called DiverseHead is shown in Figure 1-(a) based on multiple decision heads to promote precision and diversity of pseudo labels whilst using two novel perturbation methods of dropout (DT) and dynamic freezing (DF).
-In addition, we also analysed a previously proposed high-performance multi-model structure called DiverseModel (Ma et al., 2023) as shown in Figure 1-(b) for scenarios equipped with high-memory computational resources.
+However, similar to input perturbation, feature perturbation potentially generates inaccurate representations due to introducing noise that fails to accurately capture the underlying patterns in the original data.
+This may result in incorrect or noisy pseudo-labels for unlabelled examples, ultimately hindering the learning process.
+Lastly, network perturbation uses multiple networks to promote diversity of predictions [24].
+Unlike the previous two perturbation methods, network perturbation techniques introduce perturbations in a more structured manner, generated by the model itself instead of artificial noise.
+However, they generally require significantly more computational resources due to the greater number of complete networks (or their internal stages) involved.
+To further justify our design, our network perturbation strategies (dynamic freezing and dropout) introduce structured, model-driven diversity without adding external noise.
+Unlike input and feature perturbations, which risk corrupting pseudo-labels or internal representations due to artificial noise injection, our approach preserves semantic integrity while enhancing prediction diversity.
+Furthermore, DiverseHead applies lightweight perturbations to the model head, achieving a balance between diversity and efficiency without the heavy computational cost of maintaining multiple full models.
+The previous overview emphasises that while earlier perturbation techniques in SSL provide certain benefits, they also have inherent limitations, including inefficiencies in generating high-quality pseudo labels and a high computational cost.
+Facing the challenges posed by perturbation-based SSL and its complexity, it becomes imperative to explore more efficient and lightweight approaches.
+This work proposes two perturbation-based semi-supervised network architectures, coined as DiverseNet, which consist of multiple head (DiverseHead) and multiple model (DiverseModel) based SSL frameworks for various semantic segmentation applications of remote sensing imagery.
+A brief demonstration of the proposed lightweight SSL framework called DiverseHead is shown in Figure 1-(a).
+In addition, we also further analysed a previously proposed cross-network based SSL structure called DiverseModel [25], as shown in Figure 1-(b) for scenarios equipped with high-memory computational resources.
 Specifically, the contributions of this work are as follows:
   1. 1.
-    We proposed a simple but efficient semi-supervised learning architecture, promoting the utilisation of multiple decision heads called DiverseHead.
-This structure is inspired by bagging (also called bootstrap aggregating) and provides diversity for parameters and features.
-Since each head only consists of a couple of convolutional layers instead of a whole network, the architecture is relatively lightweight and adaptable to various segmentation networks.
+    We introduce DiverseHead, a simple, lightweight, and efficient SSL framework which employs multiple decision heads within a single network.
+This structure is inspired by bagging (also called bootstrap aggregating), which helps enhance pseudo label quality by integrating perturbed parameters and features within the network architecture.
   2. 2.
-    In order to provide efficient pseudo labels in the training stage, we provide a comprehensive voting mechanism based on the proposed DiverseHead structure.
-The voting mechanism not only uses the joint output from multiple heads but also individually considers pseudo labels generated from these outputs, which are called Mean Voting and Max Voting in this paper, respectively.
+    To introduce perturbation for diversifying decisions, we incorporate two key techniques, dynamic freezing and dropout, into the DiverseHead architecture, aiming to diversify the network’s parameters and high-level features, respectively.
+The proposed perturbation strategies, incorporating multiple heads, are readily applicable to a variety of state-of-the-art SSL methods and can further enhance their performance.
   3. 3.
-    For the purposes of creating perturbation in diversifying decisions, we used two important techniques, namely ‘dynamic freezing’ and ‘dropout’.
-We make these techniques fit with the DiverseHead semi-supervised learning architecture to diversify the parameters of the network.
-The two perturbation methods differ from input/feature perturbation methods since the source of feature diversity is from parameter diversity instead of introducing artificial noise to data or features.
+    We propose a dual voting mechanism, Mean Voting and Max Voting, to aggregate multihead predictions and produce high-fidelity pseudo labels for DiverseHead.
+This mechanism leverages both collective consensus and individual confidence to further enhance pseudo-label robustness during training.
   4. 4.
-    Further analysis and more detailed evaluation has been carried out on a previously proposed architecture DiverseModel (Ma et al., 2023) which is based on consistency learning among multiple neural networks with distinct weights.
-We provide a detailed comparison study for various semantic segmentation datasets in this paper.
-Also, we use Grad-CAM (Selvaraju et al., 2020) to verify the observation that different networks exhibit varied attention to the same input, culminating in the combined pseudo label derived from the multiple model predictions showcasing superior quality.
-The rest of the paper is organised as follows: Section 2 discusses the related work on semi-supervised semantic segmentation in remote sensing and some basic knowledge on ensemble machine learning whilst in Section 3 and 4, the proposed algorithms DiverseHead and DiverseModel are presented.
-Section 5 describes the utilised segmentation dataset of remote sensing imagery.
-The experimental setting and the qualitative and quantitative analyses of the experimental results are performed in Section 6.
-Section 7 concludes the paper with a summary.
-2 Related Work
-Semantic segmentation is a well-studied and wide-spread research topic in remote sensing, incorporated within applications such as land cover classification/mapping (Dong et al., 2019), building change detection (Zheng et al., 2022), road extraction (Ghandorh et al., 2022), and marine debris detection (Kikaki et al., 2022; Booth et al., 2023).
-The volume of research on semantic segmentation in remote sensing in the literature is rapidly increasing with the success of deep learning in computer vision, due to the strong task-similarity between two areas.
-Fully convolutional networks (FCNs) (Long et al., 2015) is one group of the most widespread networks for segmentation tasks.
-They have made a considerable contribution to various segmentation tasks either in remote sensing or computer vision.
-Following the FCNs’ success, SegNet (Badrinarayanan et al., 2017) and UNet (Ronneberger et al., 2015) adopt a symmetrical encoder-decoder structure with skip connections, leveraging multi-stage features within the encoder.
-Alternatively, PSPNet (Zhao et al., 2017) introduces a pyramid pooling structure that helps provide a global contextual understanding for pixel-level scene parsing.
-DeepLab architecture (Chen et al., 2017) proposes a new type of convolution called atrous convolution and an atrous spatial pyramid pooling (ASPP) operation to enable the network to have the ability to control the spatial perception range of convolution kernels by setting different dilation rates.
-DeepLab has been extended to DeepLabv3+ (Chen et al., 2018) with an improved hybrid approach that combines an encoder-decoder structure and the ASPP.
-Within DeepLabv3+, the ASPP module is incorporated following a ResNet backbone, enabling the network to delve into deeper layers and extract high-level features, which is helpful for accurately segmenting pixels around boundaries (Xia et al., 2021; Zhang et al., 2019).
-DeepLabv3+ is one of the most used networks in the literature for semi-supervised learning segmentation in the computer vision area.
-Semi-supervised learning aims to alleviate the need for expensive annotation work and improve the models’ performances by making use of unlabelled data.
-Self-training (Lee et al., 2013) (also known as pseudo labelling) represents one of the primitive semi-supervised learning strategies for both classifications (Kim et al., 2020) and segmentation (Zhu et al., 2021), and involves the generation of pseudo labels through model predictions.
-These pseudo-labels are subsequently used to retrain the model.
-Various variations of self-training have been proposed focusing on determining the appropriate pseudo labels.
-Another widely developed semi-supervised learning approach called consistency regularization (French et al., 2019) is to force networks to give consistent predictions for unlabeled inputs that undergo diverse perturbations.
-CCT (Ouali et al., 2020) employs an encoder-decoder architecture with multiple auxiliary decoders.
+    We provide a more detailed comparison study for a previously proposed architecture DiverseModel [25] on various semantic segmentation datasets in this paper.
+Also, we use Grad-CAM [26] to verify the observation that different networks exhibit varied attention to the same input.
+The rest of the paper is organised as follows: Section II discusses the related work on semi-supervised semantic segmentation in remote sensing and some basic knowledge on ensemble machine learning whilst in Section III and IV, the proposed algorithms DiverseHead and DiverseModel are presented.
+Section V describes the utilised segmentation dataset of remote sensing imagery.
+The experimental setting along with both the qualitative and quantitative analyses of the results, are presented in Section VI.
+Section VII concludes the paper with a summary.
+II Related Work
+Semantic segmentation is rapidly developing in remote sensing with the success of deep learning in computer vision.
+Due to the strong task similarity, semantic segmentation techniques are used for various remote sensing applications, such as land cover classification/mapping [27], building change detection [28], road extraction [29], and marine debris detection [30, 31].
+Specifically, Fully Convolutional Networks (FCNs) [32] have made a considerable contribution to various segmentation tasks either in remote sensing or computer vision.
+Following the FCNs’ success, SegNet [33] and UNet [34] adopt a symmetrical encoder-decoder structure with skip connections, leveraging multi-stage features within the encoder.
+Alternatively, PSPNet [35] introduces a pyramid pooling structure that helps provide a global contextual understanding for pixel-level scene parsing.
+The DeepLab architecture [36] introduces atrous convolution and atrous spatial pyramid pooling (ASPP), allowing the network to adjust the spatial receptive field of convolution kernels by using different dilation rates.
+Then, DeepLab was extended to DeepLabv3+ [37] with an improved encoder-decoder structure, which is helpful to refine segmentation results, especially around object boundaries [38, 39].
+Recently, GLOTS [40] was proposed for semantic segmentation of remote sensing images, aiming to acquire consistent feature representations by leveraging transformers in both the encoder and decoder.
+DeepLabv3+ is one of the most widely used networks in the literature for semi-supervised learning segmentation in the computer vision area.
+Semi-supervised learning aims to alleviate the need for expensive annotation work by making use of both labelled and unlabelled data.
+Self-training [41] (also known as pseudo labelling) represents one of the primitive SSL strategies for both classifications [42] and segmentation [43].
+It generates pseudo-labels using model predictions for unlabelled data, which are then utilised to retrain the model.
+Another widely developed SSL approach called consistency regularisation [20] is to force networks to give consistent predictions for unlabelled inputs that undergo diverse perturbations.
+In the context of the remote sensing field, Lu et al.
+[44] propose a weak-to-strong consistency learning for semi-supervised semantic segmentation.
+Building on the weak-to-strong consistency learning, Lv et al.
+[45] further explore the efficient exploitation of labelled and unlabelled images.
+Additionally, MIMSeg [46] integrates weak-to-strong consistency learning with masked image modelling, and DWL [47] combines it with a decoupled weighting learning framework for semi-supervised semantic segmentation of remote sensing imagery.
+In computer vision, CCT [48] employs an encoder-decoder architecture with multiple auxiliary decoders.
 These decoders introduce diversity in the output by feature perturbations specific to each auxiliary decoder.
-They calculate MSE loss between predictions of the main decoder and each auxiliary decoder without creating pseudo labels.
+They calculate the MSE loss between the predictions of the main decoder and each auxiliary decoder without creating pseudo labels.
 It is worth noting that the unsupervised loss is not used to supervise the main decoder.
-Following CCT, subsequent methods like GCT (Ke et al., 2020) and CPS (Chen et al., 2021) have been proposed to introduce network perturbation for consistency regularization.
-Specifically, they use two same network structures but with different weight initialization.
-CPS differs from GCT by employing pseudo labels to enforce consistency, whereas GCT achieves consistency regularization by utilizing predictions of networks cooperating with a flow detector.
-Although the network perturbation was achieved through different weight initialization in GCT and CPS, the ability to generate diversity is still limited.
-ICNet (Wang et al., 2022a) was proposed to use teacher networks for increasing the model difference to enhance the diversity of pseudo labels based on an iterative contrast network.
-Nonetheless, these methods consistently rely on the same network architecture.
-To introduce more diversity, we advocate the use of different model architectures.
-By leveraging various networks, we can obtain distinct and complementary features from these models even with the same input data.
-Thus, the proposed DiverseModel differs from the aforementioned methods, exploring different networks in parallel to generate more diversity of pseudo labels to improve training effectiveness.
-Ensemble machine learning is a concept that employs multiple learners and combines their predictions (Sewell, 2008).
-Bagging, a form of ensemble learning, is a technique aimed at reducing prediction variance by creating multiple iterations of a predictor and then utilizing them to form an aggregated predictor (Breiman, 1996).
+Following CCT, subsequent methods like GCT [49] and CPS [24] have been proposed to introduce network perturbation for consistency regularisation.
+Both of them use the same network structures but with different weight initialisation.
+CPS differs from GCT by using pseudo-labels generated from two networks to enforce consistency, whereas GCT achieves consistency regularisation by minimising the loss between the probability predictions of networks working in conjunction with a flow detector.
+Although the network perturbation with different weight initialisation provides some diversity for pseudo labels in GCT and CPS, the ability to generate diversity is still limited.
+Another group of SSL methods is built upon the teacher-student architecture.
+As a classic teacher-student model, MT [50] uses an EMA of the student’s weights for the teacher, trains student models with augmented inputs, and applies a consistency cost to align their outputs.
+ICNet [6] was proposed to use teacher networks to improve the quality of pseudo labels and increase the model difference based on an iterative contrast network.
+Building on the teacher-student architecture, some of its variations have become SOTA.
+For example, iMAS [51] highlights the importance of instance differences and introduces instance-specific and model-adaptive supervision for semi-supervised semantic segmentation.
+AugSeg [52] employs an enhanced data augmentation technique for input data and adaptively injects labelled information into unlabelled samples based on the model’s estimated confidence in each sample.
+These methods consistently rely on the same network architecture with different weights.
+The previously proposed DiverseModel argues that advocating the use of different models can obtain distinct and complementary features from these models, even with the same input data.
+Specifically, DiverseModel explores different networks in parallel to generate more diversity of pseudo labels to improve training effectiveness.
+Although these discussed methods achieve competitive performance, they consistently rely on multiple networks, making them bulky and potentially unsuitable for users with limited computational resources.
+Ensemble machine learning is a concept that employs multiple learners and combines their predictions [53].
+Bagging, a form of ensemble learning, is a technique aimed at reducing prediction variance by creating multiple iterations of a predictor and then utilising them to form an aggregated predictor [54].
 Specifically, bagging creates sample subsets by randomly selecting from the training dataset and subsequently utilises these acquired subsets to train the foundational models for integration.
-When forecasting a numerical result, the aggregation involves taking an average of the versions, while for predicting a class it relies on a majority vote.
-Bagging is a commonly employed approach for enhancing the robustness and precision of machine learning algorithms for classification and regression (Ren et al., 2016).
-3 DiverseHead (Cross-head Supervision)
+When predicting a numerical outcome, aggregation is done by averaging the different versions, whereas for class prediction, it is based on a majority vote.
+Bagging is a commonly employed approach for enhancing the robustness and precision of machine learning algorithms for classification and regression [55].
+III DiverseHead (Cross-head Supervision)
 Figure 2: DiverseHead: an online semi-supervised learning approach.
-This figure applies the dynamic freezing strategy: slide windows of the freezer randomly select a certain number of heads to freeze the parameter of heads (not updated by backpropagation).
-Additionally, during every iteration, all heads undergo supervision through a supervised loss, yet only a single head is randomly chosen to be updated by unsupervised loss.
+This figure applies the dynamic freezing strategy: the freezers (Dynamic Selector in the figure) randomly select a certain number of heads to freeze the parameter of heads (not updated by backpropagation).
+Additionally, during every iteration, all heads undergo supervision through a supervised loss, yet each head is randomly chosen to be updated by an unsupervised loss.
 Figure 3: The Proposed Voting Module: a voting mechanism for the pseudo label creation.
@@
 The dashed arrow serves as an illustration of a pixel voting for its classification in a segmentation map.
-This section introduces the details of the proposed DiverseHead semi-supervised learning method which is proposed to create various predictions by multiple heads illustrated in Figure 2, specifically the number of convolutional layers for each head is 2.
-The source of perturbation is from two strategies, called dynamic freezing and dropout which are shown in Figure 4 and explained in detail in Section 3.1.
-The proposed DiverseHead does not distinguish between the main or auxiliary decoder, which differs from CCT.
-During every training iteration, all heads undergo supervision through supervised losses, yet only a single head is randomly chosen to be updated by an unsupervised loss.
-In addition, the proposed method combines self-training and consistency regularization, which uses their predictions to supervise itself and forces all perturbed head to give a consistent output.
-It makes an efficient voting module shown in Figure 3 to help calculate unsupervised loss by using optimized pseudo labels.
-To specifically explain the proposed semi-supervised framework, given both a labelled dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ containing $M$ images and an unlabelled data set $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$ with $N$ images, the network $Q$ is constructed with multiple heads denoted as $\left\{head^{i}\right\}_{i=1}^{L}$, where L is the number of heads.
-Each of these heads is initialized differently.
-The expectation of the proposed semi-supervised learning architecture is to gain the trained network $Q$ leveraging these labelled and unlabelled data.
-The output of each head refers to a probability-based prediction for all classifications corresponding to its input.
-When working with labelled data, the supervised loss $\mathcal{L}_{sup}^{j}$ between the $k^{\prime}th$ ground truth $y_{k}$ and its corresponding prediction $p_{k}^{j}$ for $j^{\prime}th$ head is defined by using the standard cross-entropy loss function $\ell_{ce}$:
-$\mathcal{L}_{sup}^{j}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{j},y_{k}\right),$ (1)
-where $W$ and $H$ refer to the weight and height of input images. The final supervised loss is determined by the mean of the losses across all heads
-$\mathcal{L}_{sup}=\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}$ (2)
+This section introduces the details of the proposed DiverseHead SSL method, which uses a single network with multiple heads, illustrated in Figure 2.
+Each head has 2 convolutional layers.
+Unlike CCT, the proposed DiverseHead method treats all heads equally, avoiding a distinction between the main and auxiliary heads.
+Meanwhile, all heads benefit from labelled data by applying supervised losses to each.
+This helps create better pseudo labels in the subsequent steps of the proposed method.
+On the other hand, rather than adding artificial noise to perturb head features as in CCT, DiverseHead introduces two perturbation strategies: Dynamic Freezing and Dropout.
+They are demonstrated in Figure 4 and explained in detail in Section III-A.
+During each training iteration, in addition to supervised losses, an unsupervised loss is computed between the pseudo-label and the prediction.
+The pseudo label is derived from an efficient voting module illustrated in Figure 3.
+Although non-differentiable operations (e.g.
+Voting and Argmax) could potentially limit optimisation during training, in our framework, parameter perturbations, feature perturbations, and independent initialisation across heads and models inject sufficient diversity and stochasticity into the training process.
+These factors enable effective gradient-based optimisation despite the presence of non-differentiable components.
+The proposed method can be seen as a combination of self-training and consistency regularisation, leveraging model predictions to supervise itself and forcing all perturbed heads to produce consistent outputs.
+To provide a detailed description of the proposed semi-supervised framework, given both a labelled data set $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ containing $M$ images and an unlabelled data set $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$ with $N$ images, the network $Q$ is constructed with multiple heads denoted as $\left\{head^{i}\right\}_{i=1}^{L}$, where L is the number of heads.
+Each of these heads is initialised differently.
+The proposed SSL approach expects to gain the trained network $Q$ leveraging the labelled and unlabelled data.
+The output of each head refers to a probability-based prediction for all classifications.
+When working with labelled data, the supervised loss $\mathcal{L}_{sup,s}^{j}$ between the $s^{\prime}th$ ground truth $y_{s}$ and its corresponding prediction $p_{s}^{j}$ for $j^{\prime}th$ head is defined by using the standard cross-entropy loss function $\ell_{ce}$:
+$\mathcal{L}_{sup,s}^{j}=\frac{1}{W\times H}\sum_{i=1}^{W\times H}\ell_{ce}\left(p_{i,s}^{j},y_{i,s}\right),$ (1)
+where $W$ and $H$ refer to the width and height of input images. The final supervised loss for the $s^{\prime}th$ labelled data is determined by the mean of the losses across all heads
+$\mathcal{L}_{sup,s}=\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup,s}^{j}$ (2)
 where L is the number of heads.
-In each iteration, we also get prediction $r_{k}^{j}$ corresponding to unlabelled data $u_{k}^{j}$ from the $j^{\prime}th$ head.
+In each iteration, we also get a prediction $r_{k}^{j}$ corresponding to the $k^{\prime}th$ unlabelled data $u_{k}^{j}$ from the $j^{\prime}th$ head.
 Inspired by ensemble learning, a voting module is proposed to get high-precision pseudo labels.
-There are two voting mechanisms in the proposed voting module, called mean voting and max voting, respectively.
+There are two voting mechanisms in the proposed voting module, called mean voting and max voting, respectively shown in Figure 3.
 The former aggregates the predictions from all heads to create a combined prediction $\hat{r}_{k}^{mean}$.
@@
 The latter regards all pseudo labels as voters, in which the mean pseudo label contributes $\varphi$ weight and each individual pseudo label contributes unit weight.
-$\varphi$ is a learnable parameter and its value changes depending on the dataset and training processing.
+$\varphi$ is a learnable parameter and its value changes depending on the dataset and training process.
 The max voting module returns the class number that gets the most votes for each pixel.
-After max voting, an optimal pseudo label $\hat{y}_{k}^{final}$ is created to calculate unsupervised loss by using the cross-entropy loss function:
-$\mathcal{L}_{unsup}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(r_{k}^{main},\hat{y}_{k}^{final}\right),$ (3)
-where $r_{k}^{main}$ is the prediction from the selected head. There is a random selector to pick only one head as the main branch to be supervised by minimizing the unsupervised loss.
-Finally, the whole loss is written as
-$\mathcal{L}=\mathcal{L}_{sup}+\lambda\mathcal{L}_{unsup},$ (4)
+After max voting, an optimal pseudo label $\hat{y}_{k}^{final}$ is created to calculate unsupervised loss $\mathcal{L}_{unsup,k}$ by using the cross-entropy loss function:
+$\mathcal{L}_{unsup,k}=\frac{1}{W\times H}\sum_{i=1}^{W\times H}\ell_{ce}\left(r_{i,k}^{main},\hat{y}_{i,k}^{final}\right),$ (3)
+where $r_{k}^{main}$ is the prediction from the randomly selected single head.
+Finally, the whole loss can be written as
+$\mathcal{L}=\mathcal{L}_{sup,s}+\lambda\mathcal{L}_{unsup,k},$ (4)
 where $\lambda$ is the trade-off weight between the supervised and unsupervised losses.
-(a) Dynamic Freezing
-(b) Dropout
-Figure 4: The proposed two perturbation methods: dropout and dynamic freezing.
-In (a), the dynamic selector is designed to amplify the parameter diversity of multiple heads.
-It randomly chooses a set number of heads not to be updated by backpropagation during each training iteration, thereby enhancing the variability of parameters.
-3.1 Perturbation Methods
-Algorithm 1 DiverseHead Semi-Supervised Learning with Dynamic Freezing Pseudocode
+We use different subscripts ($s$ and $k$) for the supervised and unsupervised losses, respectively, to account for the difference in the numbers of labelled and unlabelled data.
+Specifically, the number of labelled data samples is smaller than that of unlabelled data samples.
+The solution to this issue is that, in each epoch, the labelled data set is repeatedly used in cycles for multiple iterations until all unlabelled data have been processed once.
+Figure 4: The proposed two perturbation methods: (A) Dynamic Freezing and (B) Dropout.
+Dynamic Freezing was designed to enhance the parameter diversity across multiple heads.
+During each training iteration, a specific subset of heads is randomly selected (The Dynamic Selector in the figure is used for this purpose), and their parameters are frozen, meaning they are not updated by minimising the loss in that iteration.
+These parameters are unfrozen before the next iteration begins.
+Instead, for Dropout, each channel of features passed through each head is independently zeroed out with a dropout rate $p$ during each forward pass.
+III-A Perturbation Methods
+Algorithm 1 DiverseHead Semi-Supervised Learning with Dynamic Freezing Pseudocode.
+The Labelled Training Dataset is defined as $\mathcal{B}^{l}$.
+Since the number of labelled data is smaller than that of unlabelled data, the labelled data is used in cycles for one epoch.
+We define it as $\textit{cycle}\left(\mathcal{B}^{l}\right)$.
   Initialization:
+                 Randomly initialise model Q
                  Initialise backbone using ResNet-50
-              Initialise model $Q$
-              Randomly initialise parameters for each head
   INPUT: Labelled Training Dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$
@@
   $L=length(heads)$
-  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
+  for $\left\{\left(x_{s},y_{s}\right),u_{k}\right\}_{k=1,s=k\%M}^{N}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
         $\mathcal{R}=Randint(0,L,\frac{1}{2}L)$
@@
         for $j\in\left\{1,...,L\right\}$ do
-            $p_{k}^{j}=Q^{head_{j}}(x_{k}^{j})$
-            $\mathcal{L}_{sup}^{j}=loss(p_{k}^{j},y_{k})$ based on (1)
+             $p_{s}^{j}=Q^{head_{j}}(x_{s}^{j})$
+             $\mathcal{L}_{sup,s}^{j}=loss(p_{s}^{j},y_{s})$ based on (1)
              $r_{k}^{j}=Q^{head_{j}}(u_{k}^{j})$
@@
         $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{L}\right\}\right)$
-        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
-        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
-        Update model to minimize $\mathcal{L}$
-        $\mbox{\it Unfreeze}(\left\{head^{i}\right\}_{i=R})$
+        $\mathcal{L}_{unsup,k}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
+        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup,s}^{j}+\lambda\mathcal{L}_{unsup,k}$
+        Minimize $\mathcal{L}$ to update model $Q$
+        $\mbox{{Unfreeze}}(\left\{head^{i}\right\}_{i=R})$
   OUTPUT: Trained model $Q$
@@
 Before going to the next iteration, those frozen heads are unfrozen.
-Algorithm 2 DiverseHead Semi-Supervised Learning with Dropout Pseudocode
+Algorithm 2 DiverseHead Semi-Supervised Learning with Dropout Pseudocode.
+The Labelled Training Dataset is defined as $\mathcal{B}^{l}$.
+Since the number of labelled data is smaller than that of unlabelled data, the labelled data is used in cycles for one epoch.
+We define it as $\textit{cycle}\left(\mathcal{B}^{l}\right)$.
   Initialization:
+                 Randomly initialise model Q
                  Initialise backbone using ResNet-50
-          Randomly initialise parameters for each head
                  Add dropout layer before the last convolutional layer
+                 for each head
   INPUT: Labelled Training Dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$
@@
   $L=length(heads)$
-  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
+  for $\left\{\left(x_{s},y_{s}\right),u_{k}\right\}_{k=1,s=k\%M}^{N}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
         for $j\in\left\{1,...,L\right\}$ do
-            $p_{k}^{j}=Q^{head_{j}}(x_{k}^{j})$
-            $\mathcal{L}_{sup}^{j}=loss(p_{k}^{j},y_{k})$ based on (1)
+             $p_{s}^{j}=Q^{head_{j}}(x_{s}^{j})$
+             $\mathcal{L}_{sup,s}^{j}=loss(p_{s}^{j},y_{s})$ based on (1)
              $r_{k}^{j}=Q^{head_{j}}(u_{k}^{j})$
@@
         $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{L}\right\}\right)$
-        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
-        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
-        Update model to minimize $\mathcal{L}$
+        $\mathcal{L}_{unsup,k}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
+        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup,s}^{j}+\lambda\mathcal{L}_{unsup,k}$
+        Minimize $\mathcal{L}$ to update model Q
   OUTPUT: Trained Model Q
-Another form of perturbation involves the use of dropout layers to the proposed DiverseHead structure to increase the diversity of features during training.
-We introduce a dropout layer after each convolutional block in each segmentation head of the network as shown in Figure 4 (b).
+Another form of perturbation involves the use of dropout layers in the proposed DiverseHead structure to increase the diversity of features during training.
+We introduce a dropout layer after each convolutional block in each segmentation head of the network, as shown in Figure 4 (b).
 The pseudocode of this method is given in Algorithm 2.
-As both dynamic freezing and dropout perturbation utilize the same semi-supervised learning framework, DiverseHead, the network initialization and supervised loss calculation procedures stay the same.
+As both dynamic freezing and dropout perturbation utilise the same semi-supervised learning framework, DiverseHead, the network initialisation and supervised loss calculation procedures stay the same.
 Using dropout in DiverseHead, specific components of the weights in the heads of the network are randomly assigned a value of zero with a dropout rate (determining the probability), using samples derived from a Bernoulli distribution.
 This dropout operation is employed to enhance the variability of the output predictions.
-In the conducted experiments, the dropout rate is set as a hyperparameter with a value of 0.3.
+In the conducted experiments, the dropout rate is set as a hyperparameter with a value of 0.5.
 Although the dropout-based approach adheres to the same training pipeline with dynamic freezing outlined above, the distinction is that all heads remain unfrozen.
 To evaluate the efficacy of individual perturbation methods in conjunction with the proposed DiverseHead techniques, each method is applied independently within the proposed framework.
-The performance of each combination is assessed in Section 6.
-Considering various datasets may require differing levels of diversity in pseudo labels within the proposed DiverseHead framework, adjustments can be made by altering the number of frozen heads and dropout rates in the methods of dynamic freezing and dropout, respectively.
-4 Cross-model Supervision (DiverseModel)
+The performance of each combination is assessed in Section VI.
+Considering that various datasets may require differing levels of diversity in pseudo labels within the proposed DiverseHead framework, adjustments can be made by altering the number of frozen heads and dropout rates in the methods of dynamic freezing and dropout, respectively.
+IV Cross-model Supervision (DiverseModel)
 Figure 5: DiverseModel: an online semi-supervised learning approach.
@@
 As shown in Figure 5, the DiverseModel structure includes three distinct networks, which can be various semantic segmentation networks.
-In this paper, we choose three widely used segmentation networks for experiments, which are PSPNet (Zhao et al., 2017), UNet (Ronneberger et al., 2015), SegNet (Badrinarayanan et al., 2017).
+In this paper, we choose three widely used segmentation networks for experiments, which are PSPNet [35], UNet [34], SegNet [33].
 Algorithm 3 presents the pseudocode of the DiverseModel method.
-Since different networks pay different and complementary attention to the same input, this offers the basis that they are able to benefit from each other.
-In order to provide evidence for this claim, we executed the Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2020) technique for every network employed within the framework of the DiverseModel architecture.
-Grad-CAM visualizes the areas of an image that are important to the model predictions from each network.
+Since different networks pay different and complementary attention to the same input, this offers the basis for they are able to benefit from each other.
+In order to provide evidence for this claim, we executed the Gradient-weighted Class Activation Mapping (Grad-CAM) [26] technique for every network employed within the framework of the DiverseModel architecture.
+Grad-CAM visualises the areas of an image that are important to the model predictions from each network.
 Figure 6 depicts an example grad-CAM analysis for the BUILDING class in the Potsdam data set.
@@
 The lower section showcases both the ground truth and pseudo labels generated through predictions from each network.
-In the lower right corner, the pseudo label from DiverseModel is presented.
-The labelled data is used in a regular supervised learning manner to train these models by using the standard cross-entropy loss function $\ell_{ce}$. The supervised loss is expressed as:
-$\mathcal{L}_{sup}=\frac{1}{3}\sum_{n=1}^{3}\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{n},y_{k}\right),$ (5)
-where $p_{k}^{n}$ represents the prediction of the $k\leavevmode\nobreak\ th$ pixel from the $n\leavevmode\nobreak\ th$ network.
+In the lower right corner, the pseudo-label from DiverseModel is presented.
+The labelled data is used in a regular supervised learning manner to train these models by using the standard cross-entropy loss function $\ell_{ce}$.
+The supervised loss $\mathcal{L}_{sup,s}$ is expressed as:
+$\mathcal{L}_{sup,s}=\frac{1}{3}\sum_{n=1}^{3}\frac{1}{W\times H}\sum_{i=1}^{W\times H}\ell_{ce}\left(p_{i,s}^{n},y_{i,s}\right),$ (5)
+where $p_{i,s}^{n}$ represents the $i\leavevmode\nobreak\ th$ predicted pixel of the $s\leavevmode\nobreak\ th$ sample from the $n\leavevmode\nobreak\ th$ network.
 In addition, unlabelled data is used to generate pseudo labels, which are then exploited for cross-supervision to inform each network.
-Different from the version presented in (Ma et al., 2023), all loss calculations in this work solely focus on cross-entropy loss to avoid the variations in performance resulting from different types of loss functions.
+Different from the version presented in [25], all loss calculations in this work solely focus on cross-entropy loss to avoid the variations in performance resulting from different types of loss functions.
 The predictions obtained by each network are denoted as $\{p^{1},p^{2},p^{3}\}$, which are used for generating pseudo labels $\{\hat{r}^{1},\hat{r}^{2},\hat{r}^{3}\}$ through the $argmax$ operation.
 For instance, the cross pseudo supervision loss $\mathcal{L}_{unsup}^{12}$ between the prediction $p^{1}$ from the first network and the pseudo label $r^{2}$ generated by the second network is defined as:
-$\mathcal{L}_{unsup}^{12}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{1},\hat{r}_{k}^{2}\right).$ (6)
-Algorithm 3 DiverseModel Semi-supervised Learning Pseudocode
+$\mathcal{L}_{unsup,k}^{12}=\frac{1}{W\times H}\sum_{i=1}^{W\times H}\ell_{ce}\left(p_{i,k}^{1},\hat{r}_{i,k}^{2}\right).$ (6)
+where, $p_{i,k}^{1}$ represents the $i\leavevmode\nobreak\ th$ predicted pixel of the $k\leavevmode\nobreak\ th$ sample from the $1\leavevmode\nobreak\ st$ network.
+Algorithm 3 DiverseModel Semi-supervised Learning Pseudocode.
+The Labelled Training Dataset is defined as $\mathcal{B}^{l}$.
+Since the number of labelled data is smaller than that of unlabelled data, the labelled data is used in cycles for one epoch.
+We define it as $\textit{cycle}\left(\mathcal{B}^{l}\right)$.
   Initialization:
-               Randomly initialise 3 models, PSPNet $Q^{1}$, UNet $Q^{2}$, SegNet $Q^{3}$
+                Randomly initialise three models, PSPNet $Q^{1}$,
+                UNet $Q^{2}$, SegNet $Q^{3}$
   INPUT: Labelled Training Dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$
                  Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
-  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
-        $\mathcal{L}_{sup}=loss(Q^{1}(x_{k}),y_{k})$ + $loss(Q^{2}(x_{k}),y_{k})$
-                  + $loss(Q^{3}(x_{k}),y_{k})$ based on (1)
+  for $\left\{\left(x_{s},y_{s}\right),u_{k}\right\}_{k=1,s=k\%M}^{N}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
+        $\mathcal{L}_{sup,s}=loss(Q^{1}(x_{s}),y_{s})$ + $loss(Q^{2}(x_{s}),y_{s})$
+                    + $loss(Q^{3}(x_{s}),y_{s})$ based on (1)
         $\hat{r}_{k}^{1}\leftarrow argmax(Q^{3}(u_{k}))$
@@
         $\hat{r}_{k}^{3}\leftarrow argmax(Q^{3}(u_{k}))$
-        $\mathcal{L}_{unsup}=loss(Q^{1}(u_{k}),\hat{r}_{k}^{2})$ + $loss(Q^{1}(u_{k}),\hat{r}_{k}^{3})$
+        $\mathcal{L}_{unsup,k}=loss(Q^{1}(u_{k}),\hat{r}_{k}^{2})$ + $loss(Q^{1}(u_{k}),\hat{r}_{k}^{3})$
                     + $loss(Q^{2}(u_{k}),\hat{r}_{k}^{1})$ + $loss(Q^{2}(u_{k}),\hat{r}_{k}^{3})$
                     + $loss(Q^{3}(u_{k}),\hat{r}_{k}^{1})$ + $loss(Q^{3}(u_{k}),\hat{r}_{k}^{2})$
-        $\mathcal{L}\leftarrow\mathcal{L}_{sup}+\lambda\mathcal{L}_{unsup}$
-        Update model to minimize $\mathcal{L}$
+        $\mathcal{L}\leftarrow\mathcal{L}_{sup,s}+\lambda\mathcal{L}_{unsup,k}$
+        Minimize $\mathcal{L}$ to update model $Q^{1}$, $Q^{2}$, $Q^{3}$
   OUTPUT: Trained Model $Q^{1}$, $Q^{2}$, $Q^{3}$
-The cross-pseudo supervision among three networks creates six losses in the same way. The unsupervised loss $\mathcal{L}_{unsup}$ is the average of the six individual losses, as shown below
-$\begin{array}[]{r}\mathcal{L}_{unsup}=\frac{1}{6}\left(\mathcal{L}_{unsup}^{12}+\mathcal{L}_{unsup}^{13}+\mathcal{L}_{unsup}^{21}+\mathcal{L}_{unsup}^{23}+\mathcal{L}_{unsup}^{31}+\mathcal{L}_{unsup}^{32}\right).\end{array}$ (7)
-The total loss $\mathcal{L}$ is the linear addition of $\mathcal{L}_{sup}$ and $\mathcal{L}_{unsup}$, which is previously given in equation (4)
-5 Dataset Description
-We employed diverse remote sensing datasets to assess both the proposed techniques and state-of-the-art methods, specifically including (1) the ISPRS Potsdam dataset (Rottensteiner et al., 2012), (2) the DFC2020 dataset (Robinson et al., 2021), (3) the RoadNet dataset (Liu et al., 2018), and (4) the Massachusetts Buildings dataset (Mnih, 2013).
+The cross-pseudo supervision among the three networks creates six losses in the same way. The unsupervised loss $\mathcal{L}_{unsup}$ is the average of the six individual losses, as shown below
+$\mathcal{L}_{unsup,k}$ $=\frac{1}{6}\Big{(}\mathcal{L}_{unsup,k}^{12}+\mathcal{L}_{unsup,k}^{13}+\mathcal{L}_{unsup,k}^{21}$ (7)
+$\quad+\mathcal{L}_{unsup,k}^{23}+\mathcal{L}_{unsup,k}^{31}+\mathcal{L}_{unsup,k}^{32}\Big{)}.$
+The total loss $\mathcal{L}$ is the linear addition of $\mathcal{L}_{sup,s}$ and $\mathcal{L}_{unsup,k}$, which is previously given in equation (4)
+V Dataset Description
+We employed diverse remote sensing datasets to assess both the proposed techniques and state-of-the-art methods, specifically including (1) the ISPRS Potsdam dataset [56], (2) the DFC2020 dataset [57], (3) the RoadNet dataset [58], and (4) the Massachusetts Buildings dataset [59].
 In the sequel, we share the details of each dataset utilised in this paper.
-1) ISPRS Potsdam Semantic Labeling dataset is an open-access benchmark dataset provided by the International Society for Photogrammetry and Remote Sensing (ISPRS).
-The ground sampling distance of the two modalities of the true orthophoto (TOP) and the DSM is 5 cm.
-This dataset was annotated manually into six land cover classes, which are impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background.
-This dataset provides 38 patches (all of the size $6000\times 6000$ pixels), which contain infrared (IR), red, green and blue bands orthorectified optical images with corresponding digital surface models (DSM).
-For computation purposes, we partitioned all these data tiles into $512\times 512$ patches, resulting in 3456, 201, and 1815 samples for training, validation, and test sets, respectively.
-We randomly select a quarter of the training data as labelled data and use the remaining three quarters as unlabelled data for semi-supervised learning.
-2) DFC2020 is the 2022 IEEE GRSS Data Fusion Contest dataset which is based on the SEN12MS dataset (Schmitt et al., 2019).
-It provides Sentinel-1 SAR imagery, Sentinel-2 multispectral imagery, and corresponding land cover maps across seven areas namely Mexico City, Kippa-Ring, Khabarovsk, Black Forest, Mumbai, Cape Town, Bandar Anzali between 2016 and 2017.
+1) ISPRS Potsdam Semantic Labelling dataset is an open-access benchmark dataset provided by the International Society for Photogrammetry and Remote Sensing (ISPRS).
+The true orthophoto (TOP) and DSM modalities have a ground sampling distance of 5 cm.
+Six land cover classes were identified by hand annotation of this dataset: impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background.
+This dataset has 38 patches, each measuring $6000\times 6000$ pixels.
+The patches include orthorectified optical pictures of red, green, and blue bands, as well as infrared (IR) and matching digital surface models (DSM).
+All of these data tiles were divided into $512\times 512$ patches for computational reasons, yielding 3456, 201, and 1815 samples for the training, validation, and test sets, respectively.
+We randomly select a quarter of the training data as labelled data and use the remaining three quarters as unlabelled data for all SSL approaches.
+2) DFC2020 is the 2022 IEEE GRSS Data Fusion Contest dataset, which is based on the SEN12MS dataset [60].
+It provides Sentinel-1 SAR imagery, Sentinel-2 multispectral imagery, and corresponding land cover maps with ten coarser-grained classes.
 The size of all patches is $256\times 256$ pixels.
 There are 6112, 986 and 5127 images for training, validation and test sets, respectively.
-In this paper, we utilize only one-fifth of the labelled data for training, while the remaining four-fifths are employed as unlabeled data.
-The fine-grained International Geosphere-Biosphere Program (IGBP) classification scheme in SEN12MS was aggregated into ten coarser-grained classes.
-3) RoadNet is a benchmark dataset for road detection with 0.21-m spatial resolution per pixel covering 21 typical urban regionsThe annotation data includes road surfaces, road edges, and road centerlines for each image.
-We only used road surfaces for the segmentation task.
+In this paper, one-fifth of the labelled data is for training, while the remaining four-fifths of the data is employed as unlabelled data.
+3) RoadNet is a benchmark dataset for roadnet detection with 0.21-m spatial resolution.It includes RGB images and related road surface maps for the segmentation task.
 The number of samples for training, validation and testing is 410, 45 and 387, respectively.
-In the semi-supervised learning approaches, only a quarter of the annotated training data is used, while the remaining three-quarters of the training data are employed as unlabeled data for training purposes.
+A quarter of the annotated training data is used for the supervised part, while the remaining three-quarters of the training data are employed as unlabelled data for SSL training purposes.
 4) Massachusetts Buildings dataset predominantly encompasses urban and suburban regions, encompassing structures of varying scales.
-This includes diverse buildings such as individual houses and garages, all of which are included in the labels.
-The dataset consists of 151 aerial images capturing the Boston area covering roughly 340 square kilometres, with a resolution of $1\leavevmode\nobreak\ m^{2}$ per pixel.
-The dataset was divided into subsets as follows: 137 images for training, 10 images for testing, and 4 images for validation.
-Each subset consists of images sized $1500\times 1500$ pixels.
-A quarter of the annotated training images and the remaining three-quarters of the unlabelled training data are used in semi-supervised learning.
-6 Experimental analysis
-6.1 Implementation Details
-We implemented our methods by using the PyTorch framework.
-Following (Chen et al., 2018), we used a mini-batch SGD optimizer and adopted a poly learning rate policy, where the current learning rate equals the initial learning rate multiplied by $\left(1-\frac{\text{ iter }}{\text{ max-iter }}\right)^{\text{power }}$.
+It consists of 151 aerial RGB images with a resolution of $1\leavevmode\nobreak\ m^{2}$ per pixel and corresponding building masks with the size of $1500\times 1500$ pixels.
+The dataset was split into 137 training, 10 testing, and 4 validation images with labels.
+One-quarter of the annotated training images are used for supervision, while the remaining three-quarters are treated as unlabelled for SSL.
+VI Experimental analysis
+VI-A Implementation Details
+Our approaches are implemented using PyTorch.
+Following [37], we employed a polynomial learning rate policy with a mini-batch SGD optimizer, where the current learning rate is calculated as the initial learning rate multiplied by $\left(1-\frac{\text{ iter }}{\text{ max-iter }}\right)^{\text{power }}$.
 The initial learning rate and power are set to 0.01 and 0.9, respectively.
-All experiments were conducted on the GW4 Isambard with an NVIDIA A100-sxm GPU and an AMD EPYC 7543P CPU  (McIntosh-Smith, 2014).
-We implemented various classic semi-supervised learning methods (MT, CCT, GCT and CPS) on remote sensing datasets.
-All these methods use DeepLabv3+ with ResNet-50 backbone pre-trained on ImageNet as the semantic segmentation network.
-To be fair the proposed DiverseHead is also based on DeepLabv3+ with the pretrained ResNet-50 on ImageNet.
-We comprehensively analyzed all methods by quantifying performance via class-related measures, including overall accuracy (OA), user’s accuracy (UA), producer’s accuracy (PA), mean intersection over union (mIoU), and $F_{1}$-score.
+All experiments were conducted on the GW4 Isambard with an NVIDIA A100-sxm GPU and an AMD EPYC 7543P CPU  [61].
+For a fair comparison, both the state-of-the-art methods and the proposed DiverseHead use DeepLabv3+ with a ResNet-50 backbone pre-trained on ImageNet as the semantic segmentation network.
+We comprehensively analysed all methods by quantifying performance via class-related measures, including overall accuracy (OA), user’s accuracy (UA), producer’s accuracy (PA), mean intersection over union (mIoU), and $F_{1}$-score.
 Expressions of all five performance metrics are given as follows: $\mathrm{OA}=\frac{TP+TN}{TP+TN+FP+FN}$, $\mathrm{UA}=\frac{TP}{TP+FP}$, $\mathrm{PA}=\frac{TP}{TP+FN}$, $\mathrm{mIoU}=\frac{|\mathrm{TP}|}{|\mathrm{TP}+\mathrm{FN}+\mathrm{FP}|}$, $F_{1}=\frac{2\cdot PA\cdot UA}{PA+UA}$, where TP, TN, FP, and FN refer to the numbers of pixels that are true positives, true negatives, false positives, and false negatives for each class, respectively.
-6.2 Quantitative Results and Analysis
-We evaluated the performance of the proposed approaches and some existing semi-supervised frameworks using the five performance metrics above.
-Experiments are conducted on the four previously mentioned datasets, namely Potsdam, DFC2020, RoadNet, and Massachusetts Building.
-The average results are presented in Table 1.
-The reason we first share average results across datasets is to provide a global performance for all utilised techniques, and the detailed performance comparison for all methods on each dataset will be discussed later.
-From the average results in Table 1, the proposed DiverseHead (DF) achieves the best overall performance in which there are 4 of 5 metrics reach the best compared to other methods, which are marked as red.
-DiverseModel shows the second-best performance among all methods in Table 1.
-In particular, for metrics of UA, the proposed DiverseModel exhibits an improvement of over 3.6% compared to another cross-model-based approach, CPS.
-Although the performance of DiverseHead (DT) is marginally lower than that of DiverseHead (DF) and DiverseModel, the three methods exhibit similar performance and surpass other compared methods.
+VI-B Quantitative Results and Analysis
+We evaluate the proposed approaches, along with classic and state-of-the-art SSL methods, using the five performance metrics outlined above, across four remote sensing imagery segmentation datasets: Potsdam, DFC2020, RoadNet, and Massachusetts Building.
+The average results across four datasets are presented in Table I.
+It presents an overall performance of the proposed methods using a single network in comparison to traditional SSL methods that employ dual networks.
+Since some state-of-the-art SSL methods in computer vision, such as UniMatch, iMAS, and AugSeg, are designed for RGB image segmentation and rely on RGB-based datasets for pretraining, these methods are not included in the average performance across the four datasets, including multi-band remote sensing data.
+However, a detailed performance comparison of these state-of-the-art and classic SSL methods across various datasets, including both multi-band and RGB remote sensing imagery, will be presented and discussed later in this section.
+From the average results in Table I, the proposed DiverseHead (DF) delivers the best overall performance, achieving the highest score in 4 out of 5 metrics, which are highlighted in red.
+DiverseModel demonstrates the second-best performance across most of the metrics.
+In particular, for metrics of UA, the proposed DiverseModel exhibits an improvement of over 3.48% compared to another network-perturbation-based approach, CPS.
+Although the performance of DiverseHead (DT) is slightly lower than that of DiverseHead (DF) and DiverseModel, these three methods show similar average performance, but both of them outperform the other compared methods.
 With these used remote sensing datasets, MT demonstrates notably inferior average performance across various metrics.
 CCT, GCT, and CPS exhibit comparable performance, yet the performance superiority of CPS is notably evident in its PA metric.
-Table 1: Average performance comparison with the state-of-the-art methods on four datasets. DT and DF indicate dropout and dynamic freezing, respectively.
+It is important to highlight that DiverseHead is a very lightweight method and reaches a very competitive average performance compared to others.
+Table I: Average performance comparison with the classic SSL methods with multiple networks on four datasets. DT and DF indicate dropout and dynamic freezing, respectively.
 Models OA UA PA mIoU $F_{1}$
-MT (Tarvainen and Valpola, 2017) 86.33% 74.90% 81.47% 65.97% 77.95%
-CCT (Ouali et al., 2020) 87.10% 76.61% 82.29% 67.71% 79.19%
-GCT (Ke et al., 2020) 87.43% 75.53% 82.04% 67.19% 78.63%
-CPS (Chen et al., 2021) 88.03% 75.90% 85.17% 68.10% 80.13%
-DiverseModel (Ma et al., 2023) 88.77% 79.54% 85.18% 70.92% 82.02%
-DiverseHead (DT) 88.51% 78.40% 84.84% 70.07% 81.29%
-DiverseHead (DF) 88.97% 78.99% 85.82% 71.14% 82.08%
-It is important to highlight that the DiverseHead is a significantly lightweight semi-supervised learning architecture aimed at segmentation tasks within remote sensing imagery.
-Table 2 presents the required number of parameters during training for each semi-supervised learning architecture.
-Apart from DiverseHead, all other semi-supervised learning approaches typically involve parameters exceeding three hundred megabytes during training.
-Aside from DiverseModel, all other methods are based on the same segmentation network namely DeepLabv3+.
-DiverseModel is the largest architecture although it provides competitive performance compared to other cross-model-based methods namely MT, CCT, GCT, and CPS as shown in Table 1.
-It is evident that the architectures using multiple segmentation networks during training typically need higher memory demands.
-In contrast, the proposed DiverseHead is remarkably lightweight as it only employs a single model with multiple heads which consists of only a few small convolutional layers, eliminating the need for multiple models during training.
-Thus, the parameter size of the DiverseHead (DT&DF) is only 16% bigger than that of the single network (Base in Table 2), whereas the parameter size of other reference semi-supervised architectures is at least twice that of the single network.
-Although just a a few parameters are required, the DiverseHead proposed method outperforms this state-of-the-art with at least 1% accuracy and 3% mIoU whilst obtaining similar performance to the companion method of DiverseModel.
-Table 2: The required parameter size in each semi-supervised learning approach. Base means supervised segmentation network DeepLabv3+. DT and DF present dropout and dynamic freezing, respectively.
-CPS (Chen et al., 2021) MT (Tarvainen and Valpola, 2017) CCT (Ouali et al., 2020) GCT (Ke et al., 2020)
+MT [50] 86.35% 74.93% 81.53% 66.03% 77.99%
+CCT [48] 87.14% 76.49% 82.50% 67.76% 79.23%
+GCT [49] 87.45% 75.35% 82.31% 67.19% 78.65%
+CPS [24] 88.06% 75.97% 85.27% 68.23% 80.22%
+DiverseModel [25] 88.77% 79.54% 85.18% 70.92% 82.02%
+DiverseHead (DT) 88.69% 78.51% 85.42% 70.63% 81.64%
+DiverseHead (DF) 89.00% 79.14% 85.83% 71.28% 82.17%
+Table II presents the number of parameters of SSL frameworks during training.
+Apart from DiverseHead, all other classic SSL approaches typically involve parameters exceeding three hundred megabytes.
+Among these approaches, DiverseModel is the largest architecture; however, it outperforms other network-perturbation-based methods, including MT, CCT, GCT, and CPS, as demonstrated in Table I.
+In contrast, the proposed DiverseHead is highly lightweight, using just a single segmentation model with multiple heads, each consisting of only 2 convolutional layers, thus eliminating the need for multiple networks during training.
+Specifically, the parameter size of the DiverseHead (DT&DF) is only 16% bigger than that of the single network (Base in Table II), whereas the parameter size of other reference semi-supervised architectures is at least twice that of the single network.
+Despite requiring fewer parameters, the proposed DiverseHead method surpasses the classic SSL methods by at least 1% in accuracy and 3% in mIoU, while achieving performance comparable to the largest method, DiverseModel.
+Table II: The required parameter size of each semi-supervised learning approach. Base means segmentation network DeepLabv3+. DT and DF present dropout and dynamic freezing, respectively.
+CPS [24] MT [50] CCT [48] GCT [49]
 Size (MB) 303.378 303.378 337.655 335.049
-DiverseModel (Ma et al., 2023) DiverseHead(DT) DiverseHead(DF) Base
+DiverseModel [25] DiverseHead (DT) DiverseHead (DF) Base
 Size (MB) 936.262 175.806 175.806 151.689
-Specifically, we present the results of each semi-supervised learning method for the 4 utilised datasets in Table 3.
-The performance of the DiverseNet family, namely DiverseHead and DiverseModel, is noticeably superior to that of the other listed semi-supervised learning methods across all evaluated datasets, as indicated by the results.
-In particular, for the Potsdam and Massachusetts Building datasets, the proposed DiverseHead attains the highest performance across 4 out of 5 segmentation metrics whilst DiverseModel emerges as the second-best method based on its results.
-However, in the case of the RoadNet and DFC2020 datasets, DiverseModel reaches the best for most performance metrics, securing the top position whilst DiverseHead closely follows as the second-best method, despite their similar performance.
-The performance of DiverseHead (DT) consistently falls between that of DiverseModel and DiverseHead (DF) for most used datasets apart from DFC2020.
-This dataset provides very low-resolution labels, distinguishing it from other datasets.
-This could be one of the reasons why DiverseHead (DT) outperforms DiverseHead (DF) in DFC2020.
-Table 3: Performance comparison with the state of the arts on four datasets. DT and DF present Dropout and dynamic freezing, respectively.
+Specifically, we present the results of different semi-supervised learning methods applied to the two multi-band remote sensing imagery datasets, namely Potsdam and DFC2020, in the table III.
+The performance of the DiverseNet, namely DiverseHead and DiverseModel, is noticeably superior to that of the other listed semi-supervised learning methods across these evaluated datasets.
+In particular, for the Potsdam datasets, the proposed DiverseHead (DF) attains the highest performance across 4 out of 5 segmentation metrics, whilst DiverseModel emerges as the second-best method based on its results.
+However, in the case of the DFC2020 dataset, DiverseModel reaches the best for most performance metrics, securing the top position whilst DiverseHead closely follows as the second-best method, despite their similar performance.
+Images from Potsdam consist of 4 bands, while images from DFC2020 contain 13 bands.
+The DiverseModel exhibits greater proficiency in processing datasets with more bands, while Diverse is more effective in handling images with fewer bands.
+Table III: Performance comparison with the state-of-the-art on datasets containing images with more than three bands, namely Potsdam and DFC2020.
+DT and DF present Dropout and dynamic freezing, respectively.
 Model OA UA PA mIoU $F_{1}$
-Potsdam MT(Tarvainen and Valpola, 2017) 81.98% 73.66% 78.39% 63.07% 75.95%
-CCT(Ouali et al., 2020) 82.66% 74.64% 77.62% 64.16% 76.10%
-GCT(Ke et al., 2020) 83.99% 75.65% 80.81% 65.80% 78.14%
-CPS(Chen et al., 2021) 85.00% 75.76% 82.94% 66.69% 79.19%
-DiverseModel(Ma et al., 2023) 85.76% 76.75% 83.45% 67.85% 79.96%
+Potsdam MT[50] 81.98% 73.66% 78.39% 63.07% 75.95%
+CCT[48] 82.66% 74.64% 77.62% 64.16% 76.10%
+GCT[49] 83.99% 75.65% 80.81% 65.80% 78.14%
+CPS[24] 85.00% 75.76% 82.94% 66.69% 79.19%
+DiverseModel[25] 85.76% 76.75% 83.45% 67.85% 79.96%
 DiverseHead (DT) 84.66% 77.04% 80.78% 67.12% 78.87%
 DiverseHead (DF) 85.98% 79.15% 82.87% 69.63% 80.97%
-DFC2020 MT(Tarvainen and Valpola, 2017) 78.64% 59.59% 73.57% 50.44% 65.85%
-CCT(Ouali et al., 2020) 79.71% 59.87% 76.40% 51.04% 67.13%
-GCT(Ke et al., 2020) 80.84% 61.47% 71.43% 52.17% 66.07%
-CPS(Chen et al., 2021) 81.49% 61.74% 79.46% 53.20% 69.49%
-DiverseModel(Ma et al., 2023) 81.87% 62.20% 80.69% 53.71% 70.25%
+DFC2020 MT[50] 78.64% 59.59% 73.57% 50.44% 65.85%
+CCT[48] 79.71% 59.87% 76.40% 51.04% 67.13%
+GCT[49] 80.84% 61.47% 71.43% 52.17% 66.07%
+CPS[24] 81.49% 61.74% 79.46% 53.20% 69.49%
+DiverseModel[25] 81.87% 62.20% 80.69% 53.71% 70.25%
 DiverseHead (DT) 82.02% 62.13% 80.61% 53.81% 70.18%
 DiverseHead (DF) 81.78% 61.81% 80.21% 53.46% 69.82%
-RoadNet MT(Tarvainen and Valpola, 2017) 94.53% 86.77% 87.94% 78.81% 87.35%
-CCT(Ouali et al., 2020) 95.45% 89.15% 89.91% 81.97% 89.53%
-GCT(Ke et al., 2020) 95.20% 86.27% 90.93% 80.36% 88.54%
-CPS(Chen et al., 2021) 95.66% 88.84% 90.95% 82.46% 89.88%
-DiverseModel(Ma et al., 2023) 96.84% 93.12% 92.58% 87.11% 92.84%
-DiverseHead (DT) 96.16% 91.70% 90.98% 84.70% 91.33%
-DiverseHead (DF) 96.68% 91.99% 92.73% 86.33% 92.36%
-Massachusetts MT(Tarvainen and Valpola, 2017) 90.16% 79.58% 85.97% 71.55% 82.65%
-CCT(Ouali et al., 2020) 90.59% 82.77% 85.23% 73.67% 83.98%
-GCT(Ke et al., 2020) 89.67% 78.75% 85.01% 70.42% 81.76%
-CPS(Chen et al., 2021) 89.95% 77.26% 87.33% 70.07% 81.98%
-DiverseModel(Ma et al., 2023) 90.62% 86.07% 84.01% 75.00% 85.03%
-DiverseHead (DT) 91.21% 82.73% 86.98% 74.67% 84.80%
-DiverseHead (DF) 91.43% 83.00% 87.45% 75.16% 85.17%
-In order to prove the efficiency of the multi-head framework, we considered a downgraded version of the current DiverseHead approach, called single-head supervision (SHS).
-The segmentation framework consists of only a single head.
-During each training iteration, the current model’s predictions for unlabeled data are used to generate pseudo labels through an $argmax$ operation.
+Table IV: Performance comparison with the state-of-the-art on RGB-band datasets, namely Roadnet and Massachusetts. DT and DF present Dropout and dynamic freezing, respectively.
+Model OA UA PA mIoU $F_{1}$
+RoadNet MT[50] 94.61% 86.87% 88.18% 79.04% 87.52%
+CCT[48] 95.58% 88.67% 90.74% 82.18% 89.69%
+GCT[49] 95.29% 85.53% 91.99% 80.35% 88.64%
+CPS[24] 95.81% 89.10% 91.36% 82.96% 90.21%
+FixMatch [62] 96.02% 89.91% 91.63% 83.81% 90.76%
+UniMatch [63] 95.93% 90.58% 90.79% 83.71% 90.69%
+ICNet [62] 97.15% 93.48% 93.52% 88.19% 93.50%
+iMAS [51] 97.15% 94.15% 93.03% 88.32% 93.59%
+AugSeg [52] 97.36% 94.26% 93.79% 89.06% 94.03%
+DWL [47] 97.88% 94.81% 95.48% 90.96% 95.14%
+[…]
+DiverseHead(DT) w/ Sngl-model 96.85% 92.14% 93.31% 86.91% 92.72%
+DiverseHead(DF) w/ Sngl-model 96.81% 92.60% 92.80% 86.88% 92.70%
+DiverseHead(DT) w/ iMAS 97.47% 94.24% 94.20% 89.40% 94.22%
+DiverseHead(DT) w/ AugSeg 97.70% 95.85% 93.96% 90.50% 94.90%
+DiverseHead(DT) w/ DWL 97.96% 95.33% 95.36% 91.32% 95.35%
+Massachusetts MT[50] 90.16% 79.58% 85.97% 71.55% 82.65%
+CCT[48] 90.59% 82.77% 85.23% 73.67% 83.98%
+GCT[49] 89.67% 78.75% 85.01% 70.42% 81.76%
+CPS[24] 89.95% 77.26% 87.33% 70.07% 81.98%
+FixMatch [62] 90.00% 81.34% 84.41% 72.13% 82.85%
+UniMatch [63] 89.86% 79.95% 84.79% 71.28% 82.30%
+ICNet [62] 92.24% 84.81% 88.64% 77.34% 86.68%
+iMAS [51] 92.40% 83.89% 89.90% 77.26% 86.79%
+AugSeg [52] 91.27% 80.92% 88.68% 73.98% 84.62%
+DWL [47] 91.92% 88.29% 86.00% 77.98% 87.13%
+[…]
+DiverseHead(DT) w/ Sngl-model 91.21% 82.73% 86.98% 74.67% 84.80%
+DiverseHead(DF) w/ Sngl-model 91.43% 83.00% 87.45% 75.16% 85.17%
+DiverseHead(DT) w/ iMAS 92.92% 85.43% 90.27% 78.84% 87.78%
+DiverseHead(DT) w/ AugSeg 93.35% 87.60% 89.93% 80.47% 88.75%
+DiverseHead(DT) w/ DWL 92.57% 88.49% 87.37% 79.25% 87.93%
+The proposed DiverseHead with two perturbation methods is a foundational framework, requiring fewer parameters and making it highly suitable for users with limited computational resources.
+Also, this proposed foundational framework can be easily combined with various teacher-student SSL methods to improve their performance.
+Table IV presents the results of the proposed methods and various SSL methods, including state-of-the-art methods in computer vision, which incorporate various enhancement strategies, on RGB remote sensing image segmentation datasets, specifically RoadNet and Massachusetts Building.
+DiverseHead demonstrates superior performance compared to other classic frameworks, such as MT, CCT, GCT, CPS, FixMatch and UniMatch, while maintaining a lower parameter requirement and quicker training process.
+Although some currently proposed SSL methods, iMAS, AugSeg and DWL, outperform the proposed DiverseHead, integrating these state-of-the-art methods with the proposed DiverseHead idea leads to further performance improvement and reaches the best result in Table IV.
+MTCCTGCTCPSFixMatchUniMatchiMASAugSegDiverseHead005050505010010010010015015015015020020020020063.9663.9663.9663.96120.78120.78120.78120.78206.4206.4206.4206.493.893.893.893.887.8387.8387.8387.83104.08104.08104.08104.08152.18152.18152.18152.1881.7281.7281.7281.7240.6540.6540.6540.65Training Time (s)
+Figure 7: Training time (in seconds) for 100 iterations across different SSL methods. DiverseHead refers to the dropout version of our proposed method.
+To further support the computational efficiency of the proposed method, we measured the wall-clock training time required for 100 iterations across various state-of-the-art SSL frameworks, using identical hardware and implementation settings.
+As shown in Figure 7, DiverseHead (dropout) achieves the lowest training time (40.65 seconds) among all compared methods.
+This demonstrates that our approach introduces minimal computational overhead, providing strong empirical evidence of its lightweight nature during training.
+In order to further prove the efficiency of multi-head supervision, we considered a downgraded version of the current DiverseHead approach, called single-head supervision (SHS), which consists of only a single head.
+During each training iteration, the current model’s predictions for unlabelled data are used to generate pseudo labels through an $argmax$ operation.
 These generated pseudo labels are then used to calculate the unsupervised loss.
-Also, we provide the performance of the segmentation network namely DeepLabv3+, called Base in Table 4, only trained by labelled data (part of each dataset) without using unlabelled data.
-The performance of the Base and SHS, along with the proposed DiverseHead, are presented in Table 4 to provide experimental evidence of the improvement and efficiency of the use of multiple heads in the DiverseHead architecture.
+Also, we provide the performance of the segmentation network namely DeepLabv3+, called Base in Table V, only trained by labelled data (part of each dataset) without using unlabelled data.
+The performance of the Base and SHS, along with the proposed DiverseHead, are presented in Table V to provide experimental evidence of the improvement and efficiency of the use of multiple heads in the DiverseHead architecture.
 While the single-head supervision benefits from the pseudo labels and exhibits better performance compared to the Base model, especially improving the metric of PA, both versions of DiverseHead approaches demonstrate a further significant improvement by taking advantage of multiple heads.
@@
 While the performance of DiverseHead (DT) is slightly lower than that of DiverseHead (DF), its performance is still much better than that of SHS and Base.
-Similarly, DiverseHead demonstrates superior performance on the RoadNet dataset compared to SHS and Base, particularly evident in the mIoU metric, where it exhibits a significant improvement of 6.22% over Base.
-Table 4: Performance comparison with single-head supervision (SHS) and baseline model which is only supervised by labelled data. DT and DF present dropout and dynamic freezing, respectively.
+Similarly, Both DiverseHead versions demonstrate superior performance on the RoadNet dataset compared to SHS and Base, particularly evident in the mIoU metric, where DiverseHead (DT) exhibits a significant improvement of 6.8% over the Base.
+To evaluate the effectiveness of the proposed perturbation methods, DT and DF, in comparison to input perturbation, we conducted supplementary experiments involving data perturbation by introducing Gaussian noise to the input images.
+Consistency regularisation is achieved by enforcing the network to produce consistent predictions for the original and noisy input images, which is termed Input Perturbation in Table V.
+Both versions of DiverseHead exhibit better performance than Input Perturbation, whose performance is similar to SHS.
+To use the optimal Hyperparameter applied for input perturbation for fair comparison, hyperparameter tuning experiments are conducted on two datasets to confirm the standard deviation (SD) of Gaussian noise.
+Figure 8 illustrates the variation in mIoU as the SD changes.
+The mIoU value for both Potsdam and Roadnet peaked at an SD of 0.01.
+Thus, we set the SD of Gaussian noise for input perturbation 0.01 for the Potsdam and RoadNet datasets.
+Table V: Performance comparison with single-head supervision (SHS) and baseline model which is only supervised by labelled data. DT and DF present dropout and dynamic freezing, respectively.
 Model OA UA PA mIoU $F_{1}$
@@
 SHS 83.43% 74.57% 79.93% 64.51% 77.16%
+Input Perturbation 83.89% 75.73% 80.54% 65.69% 78.06%
 DiverseHead (DT) 84.66% 77.04% 80.78% 67.12% 78.87%
 DiverseHead (DF) 85.98% 79.15% 82.87% 69.63% 80.97%
-RoadNet Base 94.86% 88.22% 88.31% 80.11% 88.26%
-SHS 95.32% 87.08% 90.82% 80.95% 88.91%
-DiverseHead (DT) 96.16% 91.70% 90.98% 84.70% 91.33%
-DiverseHead (DF) 96.68% 91.99% 92.73% 86.33% 92.36%
-Table 5: Ablation study on the number of heads for DiverseHead with dynamic freezing
+RoadNet Base 95.17% 87.62% 89.81% 80.73% 88.71%
+SHS 95.44% 87.31% 91.17% 81.37% 89.20%
+Input Perturbation 95.31% 87.42% 90.52% 81.03% 88.94%
+DiverseHead (DT) 96.85% 92.14% 93.31% 86.91% 92.72%
+DiverseHead (DF) 96.81% 92.60% 92.80% 86.88% 92.70%
+Figure 8: Variation in mIoU as the standard deviation of the Gaussian noise added for input perturbation increases
+Table VI: Ablation study on the effect of head number in DiverseHead with dynamic freezing.
 # of heads OA UA PA mIoU $F_{1}$
@@
 12 85.70% 78.54% 82.55% 68.96% 80.50%
-RoadNet                       8 96.51% 91.36% 92.55% 85.67% 91.95%
-                      10 96.68% 91.99% 92.73% 86.33% 92.36%
-                      12 96.66% 92.12% 92.58% 86.31% 92.35%
-Average                       8 91.22% 84.61% 88.20% 77.29% 86.35%
-                      10 91.33% 85.57% 87.80% 77.98% 86.66%
-                      12 91.18% 85.33% 87.57% 77.64% 86.42%
-An ablation study was conducted on the Potsdam and RoadNet datasets to explore the impact of the number of heads for the proposed DiverseHead method with dynamic freezing perturbation.
-The results in Table 5 indicate that using 10 heads yields the optimal choice across most metrics in terms of average performance though the performance difference between the versions with different numbers of heads is small on both datasets.
-Thus, we chose the 10-head version of DiverseHead for all experiments in this paper.
-Table 6: Performance comparison of the DiverseModel with its constituent networks on Potsdam dataset
+RoadNet 8 96.63% 92.15% 92.42% 86.21% 92.29%
+10 96.81% 92.60% 92.80% 86.88% 92.70%
+12 96.80% 92.59% 92.77% 86.84% 92.68%
+Average 8 91.27% 85.01% 88.14% 77.56% 86.52%
+10 91.39% 85.88% 87.83% 78.25% 86.83%
+12 91.25% 85.57% 87.66% 77.90% 86.59%
+To investigate the influence of head count in the DiverseHead framework on performance, an ablation study was conducted on the Potsdam and RoadNet datasets using the proposed DiverseHead framework with dynamic freezing perturbation.
+The results in Table VI show that using 10 heads for DiverseHead achieves the best average performance across most metrics, despite the small performance differences among the variants.
+This supports the choice of using 10 heads as an effective design decision.
+To show the efficiency of DiverseModel, consisting of three different segmentation networks, we also evaluated the performance of each member network within the DiverseModel on the Potsdam dataset in Table VII.
+The component models UNet, SegNet, and PSPNet use unlabelled data through the way of SHS.
+DiverseModel demonstrates a significant improvement across all performance metrics in the segmentation task compared to each network.
+In particular, the metric of PA experiences a notable improvement of 4.57% compared to the best-performing individual component model.
+This phenomenon can be attributed to the enhancement of pseudo-label diversity through cross-model supervision, resulting in a significant improvement in PA (recall) during the test phase.
+The findings suggest that the cross-supervision of different networks has the potential to achieve superior performance compared to the best-performing individual component.
+Table VII: Performance comparison of the DiverseModel with its constituent networks on the Potsdam dataset
 Model OA UA PA mIoU $F_{1}$
@@
 DiverseModel 85.76% 76.75% 83.45% 67.85% 79.96%
-Since the DiverseModel technique uses three different segmentation networks, we also evaluated the performance of each member network within the DiverseModel on the Potsdam dataset in Table 6.
-The component models UNet, SegNet, and PSPNet use unlabeled data through the way of SHS.
-DiverseModel demonstrates a significant improvement across all performance metrics in the segmentation task compared to each network.
-In particular, the metric of PA experiences a notable improvement of 4.57% compared to the best-performing individual component model.
-This phenomenon can be attributed to the enhancement of pseudo-label diversity through cross-model supervision, resulting in a significant improvement in PA (recall) during the test phase.
-The findings suggest that the cross-supervision of different networks has the potential to achieve superior performance compared to the best-performing individual component.
-Potsdam DFC2020 RoadNet Massachusetts
-Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Ground Truth [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-MT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-IoU = 39.38% IoU = 70.08% IoU = 79.30% IoU = 73.37%
-CCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-IoU = 51.19% IoU = 78.21% IoU = 86.24% IoU = 75.97%
-GCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-IoU = 40.35% IoU = 76.66% IoU = 80.96 IoU = 74.20%
-Potsdam DFC2020 RoadNet Massachusetts
-CPS Refer to caption Refer to caption Refer to caption Refer to caption
-IoU = 50.21% IoU = 79.39% IoU = 84.92% IoU = 74.16%
-DiverseModel Refer to caption Refer to caption Refer to caption Refer to caption
-IoU = 58.76% IoU = 80.28% IoU = 94.67% IoU = 76.84%
-DiverseHead (Dropout) Refer to caption Refer to caption Refer to caption Refer to caption
-IoU = 57.46% IoU = 80.12% IoU = 93.01% IoU = 78.93%
-DiverseHead (Dynamic Freezing) Refer to caption Refer to caption Refer to caption Refer to caption
-IoU = 58.25% IoU = 80.31% IoU = 90.35% IoU = 78.17%
-Figure 7: Visual results of each method on four datasets. The IoU score of prediction for each method is shown in its sub-caption.
-6.3 Qualitative Results and Analysis
+VI-C Qualitative Results and Analysis
 The visible results are presented in Figures 9.
-For each dataset, we randomly selected one case of the RGB image, its ground truth, and predictions for all the methods.
-Also, the IoU scores are calculated for all predictions with the ground truth for each case, which are shown in the sub-caption of each prediction.
-Based on the IoU values of visual predictions, the highest score is obtained by either DiverseModel or DiverseHead (DT&DF).
-Especially for the RoadNet dataset, both the DiverseModel and DiverseHead families achieve an IoU of over 90% surpassing other methods by at least 3.5%.
-Visually, the segmentation maps of DiverseModel and DiverseHead display better continuity and show better similarity to the ground truth than other methods.
-7 Conclusion
-In this paper, we proposed a lightweight semi-supervised learning approach based on a multi-head structure called DiverseHead.
-Based on the multi-head structure we provide two perturbation methods, namely dynamic freezing and dropout.
+For each dataset, we randomly selected 2 cases of the original image, its ground truth, and predictions for the proposed methods and various classic methods.
+Also, the mIoU scores are calculated for all predictions with the ground truth for each case, which are shown in the sub-caption of each prediction.
+Based on the IoU values of visual predictions, the highest score is obtained by either DiverseModel or DiverseHead (DT&DF) for most cases.
+Especially for cases of the RoadNet dataset, both the DiverseModel and DiverseHead families achieve a mIoU of over 90% surpassing other methods by at least 6.35%.
+Visually, the segmentation maps of DiverseModel and DiverseHead display better overall similarity to the ground truth than other methods.
+Potsdam DFC2020 RoadNet Massa- chusetts Potsdam DFC2020 RoadNet Massa- chusetts
+(1) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(2) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(3) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+64.08% 80.99% 81.40% 76.12% 41.88% 70.08% 85.33% 75.90%
+(4) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+69.24% 85.86% 85.44% 78.03% 47.18% 78.21% 86.93% 77.44%
+(5) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+68.03% 82.38% 81.57% 76.56% 47.41% 76.66% 83.61% 72.72%
+(6) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+67.19% 87.44% 85.14% 77.48% 45.06% 79.39% 86.21% 77.09%
+(7) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+74.81% 88.06% 94.67% 79.10% 51.82% 80.28% 92.95% 78.71%
+(8) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+73.39% 88.32% 94.13% 81.63% 50.43% 80.12% 90.83% 78.89%
+(9) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+76.13% 88.30% 91.79% 80.34% 55.21% 80.31% 92.04 80.87%
+Figure 9: Visual results of each method on four datasets.
+The mIoU score of the prediction for each method is shown in its sub-caption.
+The numbers in the first column are – (1): Image (2): Ground Truth (3): MT (4): CCT (5): GCT (6): CPS (7): DiverseModel (8): DiverseHead(DT) (9): DiverseHead(DF)
+VII Conclusion
+In this paper, we proposed a lightweight and efficient semi-supervised learning approach based on a multi-head structure called DiverseHead.
+Based on the multi-head structure, we provide two perturbation methods, namely dynamic freezing and dropout.
 Taking inspiration from the theory of bagging, a voting mechanism is proposed to generate beneficial pseudo-labels in the training stage.
 This simple and lightweight semi-supervised learning framework shows competitive performance for the segmentation of remote sensing imagery.
+Also, It can be easily combined with state-of-the-art SSL methods and further improve their performance.
 Furthermore, we conducted further analysis and additional evaluation on the previously proposed multi-network-based semi-supervised learning method known as DiverseModel.
-Based on the results obtained from the aforementioned four remote sensing datasets, DiverseHead and DiverseModel demonstrate comparable performance while significantly outperforming various widely used semi-supervised learning frameworks.
-It is important to highlight that the training structure of DiverseHead is significantly lighter than that of DiverseModel and other state-of-the-art methods, while still achieving competitive performance.
-This implies that DiverseHead is a better option for those who lack high-memory computation resources.
+Based on the results obtained from the aforementioned four remote sensing datasets, DiverseHead and DiverseModel demonstrate comparable performance while significantly outperforming various classic semi-supervised learning frameworks.
+From the application perspective, the proposed DiverseNet could theoretically be utilised for a wide range of image-based tasks, including medical imaging, remote sensing, and natural image segmentation.
+Having evaluated the methods on datasets with multi-band images, our approach is likely to perform competitively in MRI and other multi-band image segmentation tasks.
+We will explore more applications using the proposed method in the future.
+For technical improvement, future work could explore investigating the integration of the two proposed and other perturbation strategies in SSL, and designing more sophisticated interaction mechanisms among component models and optimising the hyperparameters to enhance their synergy.